### v1

# Federated Instruction Tuning of LLMs with Domain Coverage Augmentation

Abstract
Federated Domain-specific Instruction Tuning (FedDIT) leverages a few cross-client private data and server-side public data for instruction augmentation, enhancing model performance in specific domains. While the factors affecting FedDIT remain unclear and existing instruction augmentation methods mainly focus on the centralized setting without considering the distributed environment. Firstly, our experiments show that cross-client domain coverage, rather than data heterogeneity, drives model performance in FedDIT. Thus, we propose FedDCA, which maximizes domain coverage through greedy client center selection and retrieval-based augmentation. To reduce client-side computation, FedDCA uses heterogeneous encoders with server-side feature alignment. Extensive experiments across four domains (code, medical, financial, and mathematical) validate the effectiveness of both methods. Additionally, we explore the privacy protection against memory extraction attacks with various amounts of public data and results show that there is no significant correlation between the amount of public data and the privacy-preserving capability. However, as the fine-tuning round increases, the risk of privacy leakage reduces or converges.
1 Introduction
Table 1: Performance(%) of different augmentation settings on each domain, which are conducted via FedAvg protocol with 10 clients. Specifically, the zero-shot directly inferences without fine-tuning, while base data utilizes only the client’s local data for FedDIT. Also, we compare FedDCA with other two augmentation strategies: random sampling and direct retrieval (described in Appendix A.3), respectively.
Metric Zero-shot Base Data Random Sampling Direct Retrieval FedDCA
HumanEval/Pass@1 29.88 39.03 32.93 34.14 36.58
MMLU-Med/Acc 70.60 68.40 71.30 72.20 74.50
FPB/Acc 55.94 58.25 64.19 66.31 67.24
FiQA/Acc 18.54 14.18 13.09 19.11 35.27
TFNS/Acc 59.21 66.62 65.53 67.62 73.32
GSM8K/Exact Match 23.27 47.46 47.38 50.87 52.46
Avg. 42.91 48.99 49.07 51.71 56.56
Recently, federated instruction tuning (FedIT) has gained attention as a novel approach that leverages the principles of federated learning (FL) to enable collaborative training of large language models (LLM) in a distributed environment without the need to share private data (McMahan et al. (2017); Ye et al. (2024b)). This approach involves exchanging model parameters among distributed data holders, striking a delicate balance between preserving privacy and optimizing model efficiency. Although many FedIT frameworks have been established(Ye et al. (2024b); Kuang et al. (2023); Zhang et al. (2024a)), they have not deeply explored the practical issues that FedDIT might encounter in real-world scenarios. For example, FedIT usually requires sufficient instruction data for fine-tuning, which is difficult to meet in domain-specific fine-tuning scenarios (Zhang et al. (2024d)).
In this study, we explore Federated Domain-specific Instruction Tuning (FedDIT), a novel approach within the FL paradigm that focuses on enhancing the performance of LLMs in specific domains. Unlike the general FedIT, which seeks to improve model performance across a variety of tasks without considering local data shortage, FedDIT faces the challenge of clients having only a few local domain-specific data. To overcome this, FedDIT leverages a server-hosted public dataset that spans multiple domains to enrich the local data through the specific instruction augmentation strategy. This strategic enrichment is essential for achieving effective instruction tuning and needs to be carefully designed to prevent performance degradation. Except for the code domain, which is more about following a certain paradigm. As shown in Table 1, when clients attempt to do FedDIT only based on their local data, although some high-quality in-domain local data may be available, the limited scale restricts effective instruction tuning. Consequently, this sometimes leads to performance degradation, as indicated by the underlined values. Overall, the goal of FedDIT is to develop a domain-specific LLM that utilizes collaborative training and instruction augmentation with public data while preserving client privacy, ensuring that the model is adept at handling tasks within its targeted domain.
Additionally, the factors affecting FedDIT are still unclear. What’s worse, augmenting instructions may make the results even more unpredictable and confused about how to improve. Shepherd(Zhang et al. (2024a)) approaches this problem from the perspective of heterogeneity, constructing heterogeneity based on the topic of general instruction datasets. The study demonstrates that, for general FedIT, unlike the consensus of traditional FL, heterogeneity has a positive effect. By aggregating diverse instructions from clients, the diversity increases, thereby enhancing the model’s adaptability to various tasks. However, it just scratches the surface and does not explore issues in FedDIT.
Going one step further, we conduct experiments to unveil a new key finding (Appendix A.2): there is no linear correlation between the degree of non-independent and identically distributed (non-iid) and the performance of LLM for FedDIT. Inspired by Explore-Instruct(Wan et al. (2023)), we first define the domain coverage metric and investigate the effect of domain coverage. Results show that the cross-client domain coverage impacts the model’s performance significantly.
To maximize the cross-client domain coverage without compromising client data privacy, based on INSTA(Lee et al. (2024)), we propose a novel FedDIT algorithm, called FedDCA (Federated Instruction Tuning of LLMs with Domain Coverage Augmentation), which performs greedy client center selection and instruction augmentation via dense retrieval on the server side. The core idea of FedDCA lies in greedy client center selection, which chooses client centers that enhance the diversity and coverage of the domain-specific characteristics within the public dataset. By concentrating on maximizing domain coverage at each step, the algorithm efficiently builds an augmented train set that boosts the model’s learning and generalization capabilities, resulting in superior performance on domain-specific tasks. Furthermore, to reduce the computation overhead on the client side, we propose FedDCA with heterogeneous encoders. For feature alignment, we train a projector on the server side using public data through contrastive learning.
We demonstrate the effectiveness of FedDCA by conducting extensive experiments across four domains: code, medical, financial, and mathematical, and comparing against multiple baselines, which can be categorized into two types: unaugmented and augmented methods. For the unaugmented setting, we compare our method with FedIT, which composes of four orthodox FL methods, including FedAvg(McMahan et al. (2017)), FedProx(Li et al. (2020)), SCAFFOLD(pmlr-v119-karimireddy20a), and FedAvgM(Hsu et al. (2019)). For the augmented setting, we compare our method with random sampling, direct retrieval, LESS(Xia et al. (2024)), and Self-Instruction(Wang et al. (2023c)). Additionally, we report FedDCA’s performance with different FL strategies. Furthermore, we compare the computational efficiency of FedDCA and FedDCA on the client side. To illustrate the impact of different retrieval volumes on client data privacy, we conduct experiments against memory extraction attacks. Results indicate that as the number of training rounds increases, the method only based on the client’s local data memorizes more privacy information while the privacy leakage risk gradually decreases or converges for the augmented setting.
The main contribution is as follows:
  • We unveil a key finding: For FedDIT, there is no linear correlation between the degree of heterogeneity and the model’s performance. Instead, cross-client domain coverage impacts the model’s performance significantly, which is presented in Appendix A.2.
  • We propose a novel FedDIT algorithm (Section 4) called FedDCA to maximize the cross-client domain coverage through greedy client center selection and then perform retrieval-based instruction augmentation on the server. To further reduce the computational overhead on the client side, we propose FedDCA with a heterogeneous encoder setting. For feature alignment, we train a projector on the server side.
  • With extensive experiments (Section 5), we show the effectiveness of FedDCA and FedDCA. Furthermore, we explore privacy preservation against memory extraction attacks and conduct experiments based on various public data amounts. Results indicate that the privacy-preserving capability is not significantly correlated with the amount of public data, but the risk of privacy leakage decreases or converges as the amount of fine-tuning rounds increases.
2 Preliminaries
Federated domain-specific instruction tuning (FedDIT) aims to leverage cross-client private domain-specific instruction data and utilize the multi-domain public data on the server to achieve instruction augmentation, collaboratively enhancing the model’s performance in specific domains. Suppose $N$ distributed clients with their local private data $D_{k}^{l}$. The server has a public dataset $D^{p}$ consisting of multiple domains and is responsible for data augmentation based on different strategies and aggregation model parameters from clients. The training process follows standard FL protocol(McMahan et al. (2017)). For computational efficiency, we adopt LoRA(Hu et al. (2021)) as the fine-tuning method, which involves tuning additional parameters $\Delta\phi$ while keeping the pre-trained LLM’s parameters $\phi$ frozen. In the initial training stage, the server dispatches the pre-trained LLM parameters $\phi$ and the initialized LoRA parameters $\Delta\phi$ to each client for training $\mathcal{R}$ rounds. In the $t$-th round, the server sends the aggregated LoRA parameters $\Delta\phi^{t}$ to clients. Clients use $\Delta\phi^{t}$ to update their local LoRA parameters $\Delta\phi_{k}^{t}$ and perform instruction tuning based on augmented instruction datasets, then send the updated LoRA parameters $\Delta\phi_{k}^{t+1}$ back to the server. The server receives the LoRA parameters $\{\Delta\phi_{k}^{t+1}\mid k=1,\dots,N\}$ from each client and aggregates them to obtain the global LoRA parameters $\Delta\phi^{t+1}$ for the next round.
3 Problem Formulation
For better understanding, we list the frequently used notation in Appendix A.11. The objective of FedDIT is to enhance the domain-specific performance of LLMs through FL without sharing private data(Ye et al. (2024b); Zhang et al. (2024d)). Under the FL framework, suppose we have $N$ clients, where each client $c_{k}$ has a local dataset $D_{k}^{l}$ with its size $N_{k}^{l}$, and an augmented dataset $D_{k}^{g}$ from the server public dataset $D^{p}$, respectively. Due to constraints such as memory, computational overhead, and maximum tolerated training time, client $c_{k}$ can accept at most $N_{k}^{p}$ public instructions. Denote the augmentation strategy as $\Lambda$, through which the server performs instruction augmentation on the public dataset. If $\Lambda$ is null, it denotes clients only perform FedDIT based only on their local private data, which may cause performance degradation. Otherwise, $\Lambda$ could be categorized into two types: 1) Focusing only on its own local data distribution. 2) Taking into account the cross-client data distribution. In conclusion, the global objective of FedDIT is defined as follows:
$\operatorname*{arg\,min}_{\Delta\phi}{\left\{F(\phi,\Delta\phi)\triangleq\sum_{k=1}^{N}p_{k}\left((1-\alpha_{k})F_{k}\left(\phi,\Delta\phi_{k};D_{k}^{l}\right)+\alpha_{k}F_{k}\left(\phi,\Delta\phi_{k};D_{k}^{g}\right)\right)\right\}},$ (1)
where the first term and second term denote the accumulated fine-tuning loss computed on the client $c_{k}$’s local instructions $D_{k}^{l}$ and the augmented public instructions $D_{k}^{g}$ from the server respectively. $p_{k}$ is the weight of the $k$-th client, which is determined by the ratio of $k$-th client’s data size to all clients’ total data size. $\alpha_{k}$ is the ratio of the public data amount of its augmented instruction dataset $D_{k}$, which is computed as $\frac{N_{k}^{p}}{N_{k}^{l}+N_{k}^{p}}$. Eq.1 minimizes the summed empirical loss across clients’ augmented instructions to pursue the in-domain utility of the obtained global model. Denote $P$ as the model parameters and $\mathcal{D}$ as a specific dataset, then the client $c_{k}$’s empirical loss $F_{k}(\phi,\Delta\phi_{k};\mathcal{D})$ base on $\mathcal{D}$ is calculated as:
$F_{k}(\phi,\Delta\phi_{k};\mathcal{D})\triangleq\frac{1}{|\mathcal{D}|}\sum_{j=1}^{|\mathcal{D}|}l(P_{\phi+\Delta\phi_{k}};x_{j}),$ (2)
where $x_{j}\in D,\forall j\in\{1,2,\dots,|D|\}$. The instruction tuning loss $l(\cdot;\cdot)$ on a sample $(x,y)$ is defined as $-\sum_{t=1}^{\left|y\right|}\log\left(w\left(y_{t}|x,y_{<t}\right)\right)$, where $x$ is the formated instruction with Alpaca instruction template [^1] https://crfm.stanford.edu/2023/03/13/alpaca.html and $y$ is the corressponding response.
In contrast to the problem formulations in many prior works on FedIT, the primary distinction in the formulation of FedDIT lies in acknowledging the lack of in-domain instructions across clients. FedDIT establishes a public multi-domain dataset on the server for domain-specific instruction augmentation by a specific sampling strategy $\Lambda$. Therefore, each client only needs to hold a few high-quality domain-specific private data to collaborate with other clients to obtain a strong domain-specific LLM.
4 Method
We propose Federated Instruction Tuning of LLMs with Domain Coverage Augmentation (FedDCA), which enhances domain coverage through greedy client center selection and performs retrieval-based instruction augmentation to obtain a LLM that performs well on domain-specific tasks. We first give the overview FedDCA (Section 4.1) and define the domain coverage metric (Section 4.2), then show details of the greedy client center selection algorithm (Section 4.3) and heterogeneous encoder method FedDCA (Section 4.4).
4.1 Overview
Figure 1: Overview of FedDCA, which consists of three stages: 1) The client $c_{k}$ performs local instructions clustering and sends cluster centers $\mathcal{C}_{k}$ to the server. 2) The server first does greedy client center selection to maximize domain coverage and performs client-center-based domain data retrieval, then sends the augmented instructions $D_{k}^{g}$ to the client $c_{k}$. 3) Clients fine-tune the LLM collaboratively and the LoRA parameters $\Delta\phi$ are exchanged between clients and the server.
As shown in Figure 1, FedDCA consists of three stages, in the first stage the client $c_{k}$ clusters its local instructions and sends the cluster centers $\mathcal{C}_{k}$ to the server. Then the server performs greedy client center selection to obtain client centers that maximize the domain coverage, then performs client-center-based retrieval for domain-specific instruction augmentation. Additionally, the third stage follows the standard FedIT procedure(Ye et al. (2024b)), in which the server sends the frozen pre-trained LLM parameters $\phi$ and the initialized LoRA parameters $\Delta\phi$ to clients and then $\Delta\phi_{k}$ is exchanged between the client and the server.
The novel design of FedDCA lies in the second stage before training. As shown in Algorithm 1, the core idea is to use a greedy client center selection algorithm to obtain client centers that maximize cross-client domain coverage, and then use these client centers to retrieve instructions on the public dataset. To further reduce the computational overhead of encoding on clients, we propose a heterogeneous encoder method called FedDCA. Specifically, the client uses a small encoder locally, while the server uses a larger encoder for accurate semantics. To align features with different dimensions, we introduce a projector $w_{p}$ on the server side, which is obtained through contrastive learning.
Algorithm 1 FedDCA: greedy client center selection & domain data retrieval
1:Parameters: Number of clusters $\xi$; Encoder model $w_{enc}$; Client local datasets $D=\{D_{1},D_{2},\ldots,D_{N}\}$; Public dataset $D^{p}$; Similarity score threshold $\alpha$; Number of public data samples retrieved per client $N_{k}^{p}$; Client centers $\mathcal{P}=\{p_{1},p_{2},\ldots,p_{N}\}$; Pre-encoded public instruction embeddings $\mathcal{E}$.
2:for $i\in\{1,2,\ldots,N\}$ do
3:     $\mathcal{E}^{\prime}\leftarrow\{w_{enc}(x)\mid x\in D_{i,instruction}^{l}\}$
4:     $\mathcal{C}_{i},S_{i}\leftarrow\text{k-means}(\xi).\text{fit}(\mathcal{E^{\prime}})$triangleright Cluster local instructions, return cluster centers $\mathcal{C}_{i}$ and sizes $S_{i}$
5:end for
6:Send $\{\mathcal{C}_{i},S_{i}\mid i\in\{1,2,\ldots,N\}\}$ to the server for the greedy client center selection
7:$\mathcal{P}\leftarrow\{\mathcal{C}_{0,k}\ \mid k=\arg\max(S_{0})\}$triangleright Initialize the client center set
8:for $i\in\{1,2,\ldots,N-1\}$ do $\triangleright$ Greedy client center selection
9:     $\mathcal{S}\leftarrow\sum(\mathcal{C}_{i}\cdot\mathcal{P}^{T},\text{dim}=-1)$triangleright Compute summed similarity score of each cluster center
10:     $\mathcal{I}\leftarrow(N-i)\text{-}\arg\text{sort}(\mathcal{S})$triangleright Filter $i$ cluster centers close to the client center set $\mathcal{P}$
11:     $\mathcal{I^{\prime}}\leftarrow\arg\text{sort}(-S_{i})$triangleright Sort cluster center by cluster size in descending order
12:     $j\leftarrow\text{first element of }\mathcal{I}\cap\mathcal{I}^{\prime}$triangleright Selected cluster center $\mathcal{C}_{i,j}$
13:     $\mathcal{P}\leftarrow\mathcal{P}\cup\mathcal{C}_{i,j}$triangleright Update the client center set
14:end for
15:for $i\in\{1,2,\ldots,N\}$ do $\triangleright$ Client center based domain data retrieval
16:     $\mathcal{S}\leftarrow\mathcal{P}_{i}\cdot\mathcal{E}^{T}$triangleright Compute similarity score between $\mathcal{P}_{i}$ and $\mathcal{E}$
17:     $\mathcal{S^{\prime}}\leftarrow\{s\mid s\in\mathcal{S},s<\alpha\}$triangleright Filter instructions with similarity score larger than $\alpha$
18:     indices of $\mathcal{I}\leftarrow\text{indices of }N_{k}^{p}\text{-top in }\mathcal{S^{\prime}}$
19:     $D_{i}^{g}\leftarrow\{D^{p}_{j}\mid j\in\mathcal{I}\}$
20:     $D_{i}\leftarrow D_{i}^{l}\cup D_{i}^{g}$triangleright Obtain the augmented instruction dataset $D_{i}$
21:end for
4.2 Domain Coverage
As domain coverage directly affects the in-domain performance of the LLM (Appendix A.2), FedDCA aims to maximize the domain coverage of the cross-client data distribution $\cup_{i=1}^{N}D_{i}$ respect to the in-domain data distribution $D^{d}$, which is defined in Eq. 5. However, as public data is unreachable for clients and clients can not send the local data to the server for privacy, directly finding a cross-client dataset that maximizes the domain coverage in Eq. 5 is unrealistic. To find a proper client center set $\mathcal{P}$ that attempts to maximize the domain coverage $d(D^{d},P)$ and use $\mathcal{P}$ to perform instruction retrieval on the public data is an approximation problem.
Suppose there exists a metric space $\mathcal{X}$, a set of cross-client local instruction embeddings $\mathcal{E}\in\mathcal{X}$, and a set of cluster centers $\mathcal{P}\in\mathcal{X}$. Each client $c_{k}$ has maximum $\xi$ clusters obtained by k-means algorithm(Vardakas & Likas (2023)). As in the federated setting, communication cost is always considered. Thus, we need to find a $\mathcal{P}$ that maximizes the domain coverage while considering the communication cost, we formulate the optimization problem as follows:
$\operatorname*{arg\,min}_{\mathcal{P}}\left\{\sum_{i=1}^{N}|\mathcal{P}_{i}|+\sum_{d\in D^{d}}\left(\text{min}_{p\in\mathcal{P}}sim(d,p)\right)\right\},$ (3)
[…]
4.3 Greedy Client Center Selection & Domain Data Retrieval
As described in Section 4.2, the goal of FedDCA is to find a proper client center set $\mathcal{P}$ that maximizes the domain coverage $d(D^{d},\mathcal{P})$, as defined in Eq. 3. For computational efficiency, we propose a greedy algorithm as shown in Algorithm 1 and Figure 2 to solve this problem in polynomial time and obtain a sub-optimal solution. Given the cluster centers $\{\mathcal{C}_{i},i=1,2,\dots,N\}$ received from each client, which are obtained by clustering local instructions. FedDCA on the server consists of two main steps: greedy client center selection and client-center-based domain data retrieval.
Greedy client center selection. We will consider this problem from two aspects: 1) Select a client center that can represent the distribution of the local data. 2) To maximize the cross-client domain coverage, we filter client centers that are close to the previously selected client centers. Firstly, we randomly choose a client and select the largest cluster center as its center and be the initial of the client center set $\mathcal{P}$. Then we iteratively select the client center $\mathcal{P}_{k}$ of the top cluster size, while avoiding selecting the cluster center close to the previously selected client center to maximize the domain coverage of $\mathcal{P}$. Specifically, for the $i$-th iteration, we filter $i$-top cluster centers that have the largest summed similarity with previously selected client centers and select the center of the rest largest cluster as the $i$-th client center. This procedure is repeated until have selected $N$ client centers.
Domain data retrieval. For each client center $\mathcal{P}_{k}$, the server performs dense retrieval (detailed in Appendix A.3) on public dataset $D^{p}$ to get the top-$N_{k}^{p}$ similar public instructions, then sends $\{D_{1}^{g},\dots,D_{N}^{g}\}$ to each clients. Specifically, to avoid the overlap between public data and local private data, we set a threshold $\alpha$ to filter the public instructions that have a similarity score larger than $\alpha$ with the client center.
In summary, FedDCA can be seen as a method for establishing a comprehensive training set that captures the essence of the domain and covers its distribution. Through the iterative selection of client centers that provide unique coverage, the algorithm ensures that the selected centers are representative of the domain and achieve great domain coverage. In addition, domain data retrieval exposes the model to a broader range of in-domain features, facilitating a more robust understanding of domain-specific tasks. This not only enriches the training data but also reduces the risks of overfitting to the limited local data available to each client, which ends up with performance degradation. By leveraging diverse examples across clients, the model can better generalize its learning, leading to enhanced performance on various tasks within the domain.
Figure 2: Greedy client center selection iteratively selects the client center in each step, which synthetically considers both the representativeness of the cluster center and its distance from the previously selected client center set $\mathcal{P}$ to maximize the cross-client domain coverage.
4.4 Heterogeneous Encoder with Feature Alignment
Due to the limited computational resources on the client side, we propose a heterogeneous encoder method FedDCA to reduce the client’s computational overhead by using a small encoder $\omega$ on the client side and a larger encoder $\omega^{\prime}$ on the server side. However, the output dimension of heterogeneous encoders may not be consistent. Therefore, we introduce a projector $w_{p}$ on the server for feature alignment, as shown in Appendix A.4, we use contrastive learning(Khosla et al. (2020)) and train $w_{p}$ on the public instructions. Specifically, $w_{p}$ is composed of two fully connected layers and a ReLU activation layer in between, projecting $\omega$’s dimension to the output dimension of $\omega^{\prime}$. For $\forall{x_{i}}\in D^{p}$, where $i\in\{1,2,\dots,\mathcal{N}\}$, $\omega$ outputs embedding $h_{i}$, $\omega^{\prime}$ outputs embedding $h_{i}^{\prime}$ and $w_{p}$ outputs embedding $\varphi_{i}$. For input $x_{i}$, the positive sample pair is $(h_{i}^{\prime},\varphi_{i})$ and negative sample pairs are $\{\varphi_{j},j\neq i\}$. Let the batch size be $B$, then the training objective is defined as follows:
$\mathcal{L}=-\sum_{i=1}^{B}\log\frac{\exp\left(\text{sim}(\varphi_{i},h^{\prime}_{i})/\tau\right)}{\sum_{j=1}^{B}\mathbb{I}_{[j\neq i]}\exp\left(\text{sim}(\varphi_{i},h^{\prime}_{j})/\tau\right)},$ (4)
where $\text{sim}(\cdot,\cdot)$ denotes the cosine similarity and $\tau$ denotes the temperature parameter.
Using a projector $w_{p}$ to achieve feature alignment is an effective strategy in FL for scenarios where clients and the server utilize encoders of different sizes or capacities. The projector maps the lower-dimensional feature representations $\mathcal{H}$ from the client-side encoder to the higher-dimensional space $\mathcal{H}^{\prime}$ of the server-side encoder. By presenting the projector with positive and negative pairs, $w_{p}$ learns to adjust the projections such that the feature representations of similar items are close to each other, while those of dissimilar items are distant. In conclusion, the heterogenous encoder setting allows the system to be more scalable and flexible based on different client capacities. As a result, the projector helps in creating a unified feature space where representations from diverse sources can be combined, and utilized collaboratively.
4.5 Discussions
Computation. Through greedy client center selection, FedDCA solves the optimization problem defined in Eq. 3 in the polynomial time and quite efficient. On the client side, the k-means algorithm(Vardakas & Likas (2023)) is $\mathcal{O}(|D_{k}^{l}|)$. On the server side, as mentioned in Section 4.2, the time complexity of greedy client center selection is $\mathcal{O}\left(N\left(N\xi+\xi\log\xi\right)\right)$. Specifically, for each client, the similarity computation between selected client center set $\mathcal{P}$ and client $c_{k}$’s cluster $\mathcal{C}_{k}$ is $\mathcal{O}\left(N\xi\right)$, and the sorting process is $\mathcal{O}\left(\xi\log\xi\right)$. For domain data retrieval, given that the sorting algorithm is $\mathcal{O}\left(\mathcal{N}\log\mathcal{N}\right)$, where $\mathcal{N}$ is the size of the public dataset, thus the time complexity is $\mathcal{O}\left(N(\mathcal{N}\log\mathcal{N})\right)$. In addition, the hetergenous encoder setting FedDCA further reduces the computation overhead on the client side.
Communication. In the domain instruction augmentation stage, each client sends $\xi$ cluster centers to the server. Next, the server performs greedy client center selection and domain data retrieval, then sends retrieved public data $D_{k}^{g}$ to the client $c_{k}$, which size is $N_{k}^{p}$. In addition, in the fine-tuning stage, FedDCA follows the standard FedIT procedure, which exchanges the LoRA parameters $\Delta\phi_{k}$ between clients and the server. Specifically, for Llama3-8B model, the number of trainable LoRA parameters is just 13.6 million. Compared with the number of the frozen pretrained LLM parameters, the communication for LoRA tuning is quite efficient.
Privacy. The proposed method FedDCA does not violate the client’s privacy. Compared with other FedIT methods(Zhang et al. (2024d); Ye et al. (2024b)), the difference lies in the greedy client center selection. In this stage, the client only uploads the cluster center to the server, which is the average of embeddings to its cluster. In addition, the potential privacy leakage can be further avoided through homomorphic encryption(Acar et al. (2018)), which allows the server directly computates on ciphertext for matrix multiplication for dense retrieval.
5 Experiments
To show the effectiveness of FedDCA, we conduct extensive experiments on various domains and compare FedDCA with several baselines. This section is organized as follows: First we give details of experimental setup in Section 5.1, then we report the performance of FedDCA and compare it with several baselines in Section 5.2, and next we evaluate the computational cost between the FedDCA and FedDCA in Section 5.3. Finally, we do privacy evaluations of FedDCA in Section 5.4. For more details and results please see in Appendix A.
5.1 Experimental Setup
Dataset and Evaluation Metrics. To evaluate the performance of FedDCA, we conduct experiments on the four domains: code, medical, financial and mathematcial. The detail of constructing the public dataset is described in Appendix A.5. We choose HumanEval (code, H-Eval for short)(Chen et al. (2021)), MMLU-Med (medical, abbreviated as M-Med)(Hendrycks et al. (2021)), FPB (financial), FiQA (financial), TFNS (financial)(Yang et al. (2023a)), GSM8K (mathematical)(Cobbe et al. (2021)) for evaluation. HumanEval is evaluated using Pass@1. For MMLU-Med, FPB, FiQA and TFNS, we use accuracy as evaluation metric and exact match (EM) is used for GSM8K. Please see Appendix A.5 for more dataset details.
Baselines. We compare FedDCA and FedDCA with the following baselines: 1) Unaugmented. Zero-shot inference and FedIT, which composes of four widely used FL methods, including FedAvg(McMahan et al. (2017)), FedProx(Li et al. (2020)), SCAFFOLD(pmlr-v119-karimireddy20a) and FedAvgM(Hsu et al. (2019)). 2) Augmented. Random sampling, direct retrieval, LESS(Xia et al. (2024)), Self-Instruct(Wang et al. (2023c)). Additionally, we report FedDCA’s performance with different FL strategies in Table 2. Specifically, zero-shot inference shows the performance of the pre-trained LLM without FedDIT, which gives the lower performance bound. Direct retrieval is described in Appendix A.3. Self-Instruct augments instruction from a generative way through GPT-3.5-turbo(Sun et al. (2023)). Base on the prompt provided by Self-Instruct, we define the prompts for generating instructions and responses as shown in Appendix A.7.
5.2 Performance Analysis
Performance. We evaluate the performance of FedDCA and compare it with several baselines on four domains (see training details in Appendix A.6). As shown in Table 2, FedDCA outperforms the other nine baselines in all domains. In particular, FedDCA+FedProx achieves the best performance in most domains, with a substantial improvement from at least 5.61% to the maximum of 14.63% over other baselines on the average performance. In addition, to reduce the computation overhead of clients, FedDCA attempts to utilize heterogeneous encoders. We see that although FedDCA has a performance drop of 1.67% than FedDCA, it still outperforms other baselines in the average performance. Furthermore, we report FedDCA’s performance with different FL strategies, we can see that no FL methods can keep the leading position in all domains. In specific, FedProx and SCAFFOLD FL strategies perform better in the average performance. Overall, the result shows the effectiveness of FedDCA and FedDCA.
Table 2: Performance(%) of FedDCA, FedDCA and other nine baselines on various domains. We report FedDCA with different FL strategies in the last four rows and FedDCA+FedProx achieves the best average performance across four domains.
Method H-Eval M-Med FPB FiQA TFNS GSM8K Avg.
Zero-shot 29.88 70.60 55.94 18.54 59.21 23.27 42.91
FedAvg 39.03 68.40 58.25 14.18 66.62 47.46 48.99
FedProx 37.20 69.10 56.51 14.90 66.45 47.15 48.55
SCAFFOLD 37.80 70.20 62.71 15.27 66.49 49.27 50.29
FedAvgM 32.32 64.70 68.14 29.27 70.32 46.85 51.93
Random Sampling 32.93 71.30 64.19 13.09 65.53 47.38 49.07
Direct Retrieval 34.14 72.20 66.31 19.11 67.62 50.87 51.71
LESS 28.04 71.00 60.56 16.00 61.14 43.13 46.65
Self-Instruct 32.92 71.90 59.73 20.67 66.54 50.79 50.43
FedDCA 34.75 73.30 67.10 30.54 71.01 51.55 54.71
FedDCA+FedAvg 36.58 74.50 67.24 35.27 73.32 52.46 56.56
FedDCA+FedProx 32.92 72.40 72.93 38.18 77.55 51.25 57.54
FedDCA+SCAFFOLD 39.87 73.20 72.68 33.09 75.50 50.26 57.29
FedDCA+FedAvgM 33.53 68.90 71.45 31.45 72.52 49.76 54.60
Furthermore, Table 2 can be divided into two parts: unaugmented methods and augmented methods. We can observe that FedIT methods perform well in the code domain, even better than most augmented methods. This could be attributed to two reasons: 1) The code domain is more about following a certain paradigm. 2) As the local data is few, so compared with the same epoch and batch size, the unaugmented methods learn the same data more times, which is kind of not a fair setting. However, FedDCA+SCAFFOLD still surpasses the best baseline in the code domain, which further demonstrates the effectiveness of FedDCA.
Efficiency. We see that direct retrieval is the best baseline in the average performance. However, it does not consider the cross-client domain coverage, resulting in an overlapping retrieval data distribution. Self-Instruct(Wang et al. (2023c)) represents the upper limit performance of the generation-based method, which is restricted to the high expense of cost, limited API call frequency, and heavy quality screening process. On the other hand, FewFedPIT(Zhang et al. (2024d)) attempts to augment local data by leveraging the pre-trained Llama2 model(Li & Flanigan (2023)) to perform self-instruct manner instruction generation during training. However, there exist two concerns: 1) The pre-trained LLM cannot provide effective, stable, and high-quality instruction generation. 2) Generating instructions locally is very costly in terms of both time and computational resources. Overall, generating instructions through a self-instruct manner is not a satisfactory method for the FedDIT scenario. LESS(Xia et al. (2024)) represents the augmentation methods through gradient feature retrieval, however, the performance of LESS is not satisfactory. What’s worse, it assumes success to the validation set which is not always available and requires a warmup process on the public dataset first and then computes the gradient of both the public data and the validation set, which is computationally expensive when the size of the public dataset is enormous. In conclusion, FedDCA achieves better performance while requiring lower computational resources and a more relaxed training condition, which shows the efficiency of FedDCA.
Table 3: Domain coverage of FedDCA, FedDCA and other baselines on four domains, where FedIT methods only use local data without domain augmentation.
Method Code Med. Fin. Math. Avg.
FedIT 0.8126 0.6990 0.8529 0.7871 0.7879
Random Sampling 0.8512 0.7940 0.9196 0.8651 0.8575
Direct Retrieval 0.9396 0.8830 0.9293 0.8967 0.9122
LESS 0.8509 0.7737 0.8917 0.8352 0.8379
Self-Instruct 0.8966 0.8586 0.9015 0.8811 0.8845
FedDCA 0.9532 0.8972 0.9538 0.9096 0.9285
FedDCA 0.9766 0.9348 0.9815 0.9320 0.9562
Domain coverage. The domain coverage of each method is shown in Table 3. Also, we show the visualization of the domain coverage across different instruction augmentation methods in Appendix A.10. For the augmented methods, it can be observed that the higher domain coverage in a particular domain correlates with the better performance of the method in that domain. Specifically, given that FedDCA aims to maximize the domain coverage through greedy client center selection, therefore it achieves the highest domain coverage in all four domains, which surpasses other baselines from 4.82% to 21.36% average relative improvement. In addition, as FedDCA utilizes a smaller encoder on the client side for computational efficiency, it sacrifices some semantic precision, resulting in a 2.89% averaged domain coverage drop in relative. However, FedDCA still surpasses the best baseline by 1.79% on average. In brief, the result proves the effectiveness of FedDCA and FedDCA in domain coverage augmentation.
5.3 Computation Analysis
Table 4: Computational cost comparison between FedDCA and FedDCA. We report the encoding time on the public dataset and the model size of each encoder.
Method Encoding Time Model Size
FedDCA 15 min 46 s 335 M
FedDCA 5 min 02 s 22.7 M
Heterogenous encoder setting FedDCA allows the system to be more scalable and flexible based on different client capacities. To show the computational efficiency of FedDCA compared with FedDCA, we compare these two methods from the following aspects: 1) Encoder model size. 2) Encoding time overhead. We first give the evaluation setup of computation analysis.
Evaluation setup. The encoders used for FedDCA and FedDCA on the client side are bge-large-en-v1.5 and all-MiniLM-L6-v2 respectively (detailed in Appendix A.6). Then we show the model size and the time overhead of encoding the public instructions in Table 4
We can observe that in comparison with FedDCA, there is a large computational and time overhead advantage with FedDCA, while achieves an acceptable performance (shown in Table 2), which is accomplished by implementing a learnable projector on the server side. This projector maps the lower-dimensional client-side feature to the higher-dimensional space of the server-side encoder output. Consequently, the projector facilitates the creation of a unified feature space where representations from diverse sources can be integrated and used collaboratively.
5.4 Privacy Analysis
Next, we evaluate the privacy-preserving capability of different ratios of public data against memory extraction attacks(Carlini et al. (2021); Zhang et al. (2024c)), which utilizes the autoregression nature of LLM to extract information from the memory(Xu et al. (2024)).
Table 5: Privacy-preserving capability against memory extraction attack with different amounts of augmented public data. We report the average ROUGE-L score of each setting in four domains.
Retrieval Num Code Med. Fin. Math.
Base Data 0.2338 0.1579 0.1527 0.1718
1000 0.2228 0.1505 0.1451 0.1635
2000 0.2226 0.1531 0.1445 0.1627
3000 0.2195 0.1513 0.1454 0.1555
4000 0.2162 0.1434 0.1438 0.1631
5000 0.2119 0.1501 0.1445 0.1647
Evaluation setup. We focus on a single client’s fine-tuning process, which has 100 local data and adopts FedDCA to do the instruction augmentation with different amounts of public data from 1000 to 5000. We set up 10 clients with full participation for 10 rounds. Specifically, we record every round’s ROUGE-L score(Lin (2004)) of the client $c_{0}$. We use the designed prompt to extract the instruction that LLM memorizes, which is shown in Appdix A.8. Specifically, for each setting, we repeat the memory extraction 100 times and report the average ROUGE-L score computed based on each generated instruction and all the local data’s instructions to evaluate the privacy-preserving capability of different amounts of public data.
Figure 3: ROUGE-L score of each pair of base-data-only and augmented fine-tuning in four domains per round, where different monochromatic represents each domain.
Results. We report the average ROUGE-L score of base-data-only fine-tuning and various amounts of public data fine-tuning in Table 5. The result shows that at the same training round, there is no significant correlation between the public data ratio and the privacy-preserving capability. In addition, after 10 rounds of FedDIT, the base-data-only fine-tuning’s average ROUGE-L score all exceeds the instruction-augmented settings in four domains. Specifically, base-data-only fine-tuning exceeds augmented settings by 6.69%, 5.55%, 5.55%, and 6.16% on average, relatively.
The trend of the ROUGE-L score of base-data-only and augmented fine-tuning of each domain per round is shown in Figure 3. In specific, the augmented fine-tuning uses 5000 public data. We can see that the ROUGE-L score of augmented settings in each domain firstly increases and then starts decreasing or converging to a certain range. The base-data-only setting keeps the rising trend, especially the code domain, which indicates that as the number of training rounds increases, the base-data-only fine-tuning will remember more privacy information, while the risk of privacy leakage of the augmented fine-tuning will gradually decrease or converge.
6 Conclusion
We reveal that for FedDIT, the model performance is not linearly correlated with data heterogeneity, but is significantly impacted by cross-client domain coverage. To address this, we propose FedDCA, a domain-specific instruction augmentation method that maximizes domain coverage through greedy client center selection and retrieval-based augmentation. Additionally, FedDCA uses a heterogeneous encoder to reduce client-side computation overhead. Experiments across four domains demonstrate the effectiveness of FedDCA and the efficiency of FedDCA. Further analysis shows that as fine-tuning progresses, the risk of privacy leakage from augmented data decreases or converges.
Appendix A Appendix
A.1 Related Work
Federated Instruction Tuning. Instruction tuning has been widely applied across various application areas of large language models (LLM), serving as a key technique to enhance the capabilities and controllability of LLM (Zhang et al. (2024b); Wei et al. (2022)). Recently, federated instruction tuning (FedIT) has emerged as an effective strategy for the distributed optimization of LLMs, leveraging federated learning (FL) protocols to improve the handling of privacy-sensitive tasks in real-world scenarios. So far, several FedIT frameworks(Ye et al. (2024b; a)) have been established to evaluate the effectiveness of FedIT across multiple datasets, tasks, and FL methods. While these platforms provide a foundation for research, they have not yet introduced more complex federated algorithms and deeply investigate the challenging problems and factors affecting FedIT, which are crucial for advancing this field.
PrivateLoRA(Wang et al. (2023b)) addresses privacy and efficiency issues by exploiting the low-rank properties of residual activations to reduce communication costs, significantly lowering the communication overhead through collaborative computation between the server and clients while effectively maintaining the privacy of local data. While FewFedPIT(Zhang et al. (2024d)) focuses on the few-shot learning setting in FedIT, using self-generated data by pre-trained LLMs locally to mitigate the paucity of data and first discussing memory extraction attacks within FedIT.
Domain Instruction Augmentation. In the real world, there is an urgent need for training LLMs with specific functionalities (e.g., reasoning capabilities) or domain-specific LLMs (e.g., code(Nijkamp et al. (2023); Luo et al. (2023b)), medical(2024-Zhang-AlpaCare-), financial(Yang et al. (2023b; a); Zhang et al. (2023a); Wu et al. (2023)), mathematical(Yue et al. (2024); Luo et al. (2023a))). Existing works tend to use open-source domain-specific instruction tuning datasets for training. However, the target domain may not always have corresponding ready-made domain-specific instruction datasets. Even if they exist, these datasets are often limited in scale.
Several studies investigate domain-specific instruction augmentation, which can be categorized into three aspects: 1) Reusing human-curated public datasets(Wang et al. (2023a); Zhang et al. (2023c); Jiao et al. (2023); Lee et al. (2024); Xia et al. (2024)). For instance, Parrot(Jiao et al. (2023)) enhances translation capabilities of LLMs in chat by converting bilingual sentences into instruction-following datasets. Furthermore, works like INSTA(Lee et al. (2024)) and LESS(Xia et al. (2024)) attempt efficient domain-specific instruction augmentation via dense retrieval. INSTA(Lee et al. (2024)) uses instructions without responses for effective retrieval. LESS(Xia et al. (2024)) assumes access to a validation set and uses the warmup LLM’s gradients of the train and validation sets for retrieval. 2) Using seed tasks for self-instruct(Luo et al. (2023b); Wan et al. (2023)): Explore-Instruct(Wan et al. (2023)), for example, employs activate exploration to tree-model domain tasks from both depth and breadth, increasing seed instructions’ coverage of the domain, and subsequently generating broader in-domain instructions through self-instruct. 3) Scaling instructions from the web: Recent works(Yue et al. (2024); Zhou et al. (2024)) highlight the immense potential of mining naturally occurring instructions from the internet. Compared to generated data, web-mined instructions exhibit less bias and greater diversity. MAmmoTH2(Yue et al. (2024)) firstly retrieves domain-relevant texts and employs a LLM to extract Q-A pairs, further refining them into instruction-response pairs. Jiuzhang3.0(Zhou et al. (2024)) distills GPT’s instruction generation capabilities into a smaller model and then uses it to generate instructions from the internet. In conclusion, models fine-tuned with augmented instructions have shown promising domain-specific capabilities.
Inspired by INSTA(Lee et al. (2024)), this paper uses a mixed dataset composed of code, medical, financial, mathematical, and general datasets to represent public data and does domain-specific instruction augmentation based on the client’s local instructions.
A.2 What Truly Counts in FedDIT
This section will first analyze the correlation between different non-iid levels and the model’s performance with separate experiments for both single and multiple domains in Section A.2.1. Further, to demonstrate the impact of non-iid on FedDIT with different test sets, we compare the performance of the global model trained on augmented data based on iid and non-iid cross-client data distribution. Additionally, we suggest that domain coverage is a key factor for FedDIT in Section A.2.2.
A.2.1 Data Heterogeneity Is Not Matter In FedDIT
Following the traditional approach of constructing different degrees of non-iid which are widely used in federated learning(Wang et al. (2020); Yurochkin et al. (2019)), we adopt Dirichlet distribution to construct various heterogeneity and use k-means with the cluster num $\xi=100$ to pseudo labeling instructions. Dirichlet distribution is affected by the hyperparameter $\alpha$, which enhances heterogeneity with a smaller $\alpha$ and decreases heterogeneity with a larger $\alpha$. We choose four widely used heterogeneity, which are $\alpha=[0.01,0.1,1,10]$(Ye et al. (2023); Zhang et al. (2023b); li2021model). Figure LABEL:fig:data_heterogeneity shows a visualization of the data distribution with different heterogeneity on the code dataset.
We perform instruction tuning both on single-domain and multi-domain with different heterogeneity. Specifically, for the single domain, we only use the in-domain data and then perform FedIT, and for multi-domain, we use the mixed dataset which consists of four domain-specific instruction datasets and a general instruction dataset, which are described in Appendix A.5. The training details are shown in Appendix A.6. Results for single domain and multi-domain are shown in Table 6. Both single-domain and multi-domain clearly show that the performance of LLM does not decrease due to the increase of data heterogeneity, but shows a nonlinear correlation, which indicates that the performance of LLM does not directly depend on data homogeneity, and exists other factors that play a key role.
Table 6: Performance(%) of different heterogeneity on single domain and multi-domain.
$\alpha$ Single Domain Multi-Domain
10 1 0.1 0.01 10 1 0.1 0.01
H-Eval 32.92 34.14 31.09 34.75 34.14 31.70 34.75 34.75
M-Med 70.6 69.9 71.0 71.3 71.7 71.3 69.9 71.3
FPB 67.90 72.11 73.76 68.89 60.06 61.63 60.97 62.04
FiQA 43.63 45.45 41.81 32.00 10.90 12.72 17.81 9.09
TFNS 74.37 76.00 75.71 72.69 65.36 65.24 64.69 64.61
GSM8K 53.75 51.40 51.93 52.76 46.5 43.5 41.5 45.5
Avg. 57.20 58.17 57.55 55.40 48.11 47.68 48.27 47.88
A.2.2 Domain Coverage: A Key Factor in FedDIT
Explore-Instruct(Wan et al. (2023)) enhances the coverage of domain-specific seed tasks through active exploration, then uses the self-instruct method for instruction data augmentation. This approach highlights the impact of domain coverage on domain-specific instruction tuning. Inspired by Explore-Instruct, we attempt to conduct more in-depth and extensive experiments to study the effect of domain coverage on FedDIT. Firstly we define the domain coverage of cross-client data in the FL setting. Assume the dataset of in-domain data $D^{d}$ represents the latent data distribution of this domain and the cross-client data is defined as $D^{c}=\cup_{k=1}^{N}\left(D_{k}^{l}\cup D_{k}^{g}\right)$. Inspired by the facility location function(Cornuejols et al. (1983)), we define the domain coverage of $D^{c}$ respect to $D^{d}$ as follows:
$d(D^{d},D^{c})=\frac{1}{\left|D^{c}\right|}\sum_{d\in D^{d}}\text{max}_{v\in(D^{c}\cap D^{d})}sim(d,v),$ (5)
where $sim(d,v)$ is the similarity between $d$ and $v$. Specifically, we use the cosine similarity function in FedDCA. Note that we only use the in-domain data in $D^{c}$ to calculate the domain coverage, because the out-of-domain data would mislead the domain coverage evaluation, as shown in Figure 7.
Figure 7: Misleadning of out-of-domain data on domain coverage calculation.
To better align with the real-world scenarios, we explore instruction augmentation based on both iid and non-iid cross-client data distribution and adopt direct data retrieval as described in Appendix A.3 for FedDIT. We set the number of clients to 10, while each client has 100 local instructions and obtains 5000 augmented public instructions from the server.
To construct iid data distribution, we randomly sample 1,000 from multi-domain datasets (code, medical, financial, mathematical, and general), which is detailed in Table 8 and divide them into 10 shards as each client’s local data. To construct non-iid data distribution, we perform k-means clustering with $\xi=100$. Each client randomly samples 100 instructions from different randomly selected clusters. Furthermore, for both iid and non-iid settings, direct data retrieval is performed based on the client’s local data.
Table 7: Performance(%) and domain coverage of iid and non-iid settings on different domains. The higher domain coverage correlates with better performance.
Test Set Performance (%) Domain Coverage
iid non-iid iid non-iid
H-Eval 35.36 33.53 0.8538 0.7994
M-Med 70.20 71.00 0.7800 0.8027
FPB 58.58 64.19 0.8523 0.9327
FiQA 17.09 19.27
TFNS 66.16 69.09
GSM8K 40.50 38.50 0.9137 0.8448
Table 7 presents the performance of FedDIT on different domains with iid and non-iid settings and shows the domain coverage in four domains. We can observe that both iid and non-iid settings outperform in some domains, but both collectively indicate that higher domain coverage correlates with better performance.
A.3 Direct domain data retrieval
In this section, we first show the detail of instruction-based dense retrieval in Appendix A.3.1, which is both used in direct retrieval and FedDCA. Then we explain the direct retrieval algorithm in Appendix A.3.2.
A.3.1 Instruction Based Dense Retrieval
[…]
$\textbf{Score}(\mathcal{E},I)=sim(\mathcal{E},w_{enc}(I)),$ (6)
where $sim(\cdot,\cdot)$ is the cosine similarity function. Based on the computed similarity between $\mathcal{E}$ and each $I$ in the public data, we then select the top-$N_{k}^{p}$ instructions as the retrieved public data for domain data augmentation.
A.3.2 Direct Retrieval
The direct domain data retrieval only utilizes instructions of the client’s local data without responses to perform the retrieval-based domain data augmentation. The detailed algorithm is described in Algorithm 2. In specific, for each client, we first encode the local instructions by the encoder model $w_{enc}$. Then we use the k-means algorithm to cluster the embeddings into $\xi$ clusters. Then the $\xi$ cluster centers are sent to the server and do the retrieval-based domain data augmentation. Then the retrieved public data is sent to the client for instruction tuning.
Algorithm 2 Direct domain data retrieval
1:Parameters:
2:Clients’ local datasets $D=\{D_{1},D_{2},\ldots,D_{N}\}$; Public dataset $D^{p}$; Local datasets $D^{l}=\{D^{l}_{1},D^{l}_{2},\ldots,D^{l}_{N}\}$; Number of clusters $\xi$; Encoder model $w_{enc}$; Pre-encoded public instruction embeddings $\mathcal{E}$.
3:$D^{g}\leftarrow\emptyset$
4:for $i\in\{1,2,\ldots,N\}$ do
5:     $\mathcal{E^{\prime}}\leftarrow\{w_{enc}(x)\mid x\in D_{i,instruction}^{l}\}$triangleright Encode the local instructions
6:     $\mathcal{C}\leftarrow\text{k-means}(\xi).fit(\mathcal{E^{\prime}})$triangleright Cluster local instructions, return cluster centers $\mathcal{C}$
7:     $\mathcal{S}\leftarrow\mathcal{C}\cdot\mathcal{E}^{T}$triangleright Compute the similarity score between $\mathcal{C}$ and $\mathcal{E}$
8:     $\mathcal{I}\leftarrow\emptyset$
9:     for $j\in\{1,2,\ldots,\xi\}$ do
10:         $\mathcal{S}^{\prime}\leftarrow\{s\mid s\in S_{j},\text{index}(s)\notin\mathcal{I}\}$triangleright Filter the selected indices
11:         $\mathcal{I^{\prime}}\leftarrow\text{indices of the top-}\frac{N_{k}^{p}}{\xi}\text{ elements in }\mathcal{S}^{\prime}$triangleright Retrieve the top $\frac{N_{k}^{p}}{\xi}$ indices
12:         $\mathcal{I}\leftarrow\mathcal{I}\cup\mathcal{I^{\prime}}$triangleright Update the selected indices
13:     end for
14:     $D^{g}_{i}\leftarrow\{D^{p}_{j}\mid j\in\mathcal{I}\}$triangleright Selected public instructions
15:     $D_{i}\leftarrow D^{g}_{i}\cup D_{i}^{l}$triangleright Obtain the augmented instruction dataset $D_{i}$
16:end for
A.4 Projector
To reduce the computational overhead of clients, we propose a heterogeneous encoder method called FedDCA. Inspired by LLaRA(Liao et al. (2023)), we utilize a projector to do the dimension alignment between the small encoder on the client side and the large decoder on the server side. As shown in Figure 8, where $d_{l}$ is the dimension of the client-side encoder, $d_{g}$ is the dimension of the server-side encoder.
Figure 8: The projector used on the server side for feature alignment, which consists of two fully connected layers and a ReLU activation layer in between.
A.5 Train and Test Dataset Information
Here we provide the train and test dataset details of code (CodeAlpaca [^2] https://huggingface.co/datasets/sahil2801/CodeAlpaca-20k), medical (MedAlpaca [^3] https://huggingface.co/datasets/medalpaca), financial (FinGPT(Yang et al. (2023a))) and mathematical (MathInstruct [^4] https://huggingface.co/datasets/TIGER-Lab/MathInstruct) domain, respectively.
Table 8: Train set information of each domain.
Dataset name Domain $N_{sample}$
CodeAlpaca Code 20,022
MedAlpaca Medical 33,955
FinGPT Financial 76,772
MathInstruct Mathematical 224,567
Alpaca General 52,002
As shown in Table 8, the public data consists of four domain-specific instruction datasets and a general instruction dataset, which are CodeAlpaca, MedAlpaca, FinGPT, MathInstruct, and Alpaca, respectively. For each domain’s FedDIT, we randomly select 1000 samples from the in-domain instruction and split them into 10 shards as each client’s local dataset. The rest instructions are used as the public dataset $D^{p}$.
Table 9: Test set information of each domain.
Dataset name Domain $N_{sample}$ Metric
HumanEval Code 164 Pass@1
MMLU-Med Medical 1,089 Acc
FPB Financial 152 Acc
FiQA Financial 35 Acc
TFNS Financial 299 Acc
GSM8K Mathematical 1,319 Exact Match
Table 9 shows the detail of the test set used for evaluation, HumanEval is used for the code domain, MMLU-Med is used for the medical domain, FPB, FiQA and TFNS are used for the financial domain, and GSM8K is used for the mathematical domain. Specifically, MMLU-Med uses subjects anatomy, clinical_knowledge, college_biology, college_medicine, medical_genetics and professional_medicine in MMLU.
A.6 Implementation Details
We consider FedDIT in the cross-device scenario, $N$ = 10 clients, $\mathcal{R}$ = 30 rounds, where we randomly sample 2 clients to be available for each round. Then each available client performs FedDIT for 10 steps with AdamW optimizer and the batch size $B=32$ in a round. The initial learning rate is $55e-$ 5 with a cosine learning rate scheduler. Our experiment utilizes the widely used LLM, Llama3-8B [^5] https://huggingface.co/meta-llama/Meta-Llama-3-8B as the base model with 2048 max sequence length and adopts LoRA tuning method. The rank of LoRA is 16 and the scalar alpha is 16. For k-means(Vardakas & Likas (2023)), we set cluster num $\xi$ = 10 and for FedDCA we set the similarity threshold $\alpha$ = 0.7. For FedDCA, we set the temperature parameter $\tau$ = 0.5 for contrastive learning(Khosla et al. (2020)). We utilize bge-large-en-v1.5666https://huggingface.co/BAAI/bge-large-en-v1.5 as both the client and server’s encoder as default, which outputs embeddings of 1024 dimensions. While we utilize all-MiniLM-L6-v2777https://huggingface.co/sentence-transformers/all-MiniLM-L6-v2 as the client’s small encoder which outputs embeddings of 384 dimensions and bge-large-en-v1.5 as the server’s encoder for FedDCA.
A.7 Prompts used in the Self-Instruct data generation
To generate the Self-Instruct data, we prompt GPT-3.5 to generate the instruction with the designed prompt in Figure 9. Specifically, we randomly sample two examples from the client’s local data to guide GPT-3.5 generating the in-domain instruction, and sample one example from the client’s local data for one-shot in-context learning to guide GPT-3.5 generating responses into the example’s format.
Figure 9: Prompts used in the Self-Instruct data generation. (a) Prompt for generating new instructions. Two examples are randomly sampled from the client’s local data for in-context demonstration. (b) Prompt for generating responses. We prompt GPT-3.5 to generate responses with a randomly selected example for one-shot in-context learning.
A.8 Prompt used for memory extraction attack
As we use Llama3-8B as our base model and format the instructions and responses into the Alpaca’s format, to utilize the auto-regression nature of LLM to extract the instruction, we prompt the model to generate the instruction using the prompt in Figure 10, which is exactly the prefix of the Alpaca’s template.
Figure 10: Prompt used for memory extraction attack.
A.9 Further Analysis
To further study the effect of different hyperparameters of FedDCA, we undertake a thorough analysis including various retrieval amounts, and different k-means cluster numbers $\xi$. In addition, we perform the ablation study on whether using the similarity threshold $\alpha$ in the greedy client center selection.
Effect of Retrieval Number. We report the performance and the corresponding domain coverage of FedDCA with different retrieval amounts on the four domains in Figure 11, respectively. We can see that the domain coverage of FedDCA is increasing along with the retrieval amount in different trends, as well as the performance. Specifically, the domain coverage of each domain increases by 6.36%, 18.69%, 5.04%, and 8.06% in relative, respectively. Along with domain coverage increasing, the performance of FedDCA is increasing by 3.05%, 2.20%, 4.08%, and 6.95% for each domain. Noted that although the code domain is more about following a certain paradigm, which could perform well with a few data and more fine-tuning rounds, it still could benefit from the instruction augmentation.
Figure 11: Effect of different retrieval amounts on the performance of FedDCA and its domain coverage. We show the results on four domains separately. Here we use the FPB test set to evaluate the performance in the financial domain.
Impact of Different Cluster Number. The hyperparameter $\xi$ is the number of clusters in the k-means algorithm. The experiment is conducted on $\xi=[N,2N,4N,8N]$, where $N$ is the number of clients. Following the setting in Appendix A.6, we set $N=10$. We report the domain coverage of the augmented dataset via FedDCA with different cluster numbers $\xi$ on the four domains in Figure 12, respectively. Results show that there is no best $\xi$ for all domains. Specifically, the best $\xi$ of code, medical, financial, and mathematical domains are 80, 80, 40, and 10, respectively.
Figure 12: Impact of different cluster number $\xi$ on the cross-client domain coverage. We report the domain coverage of the augmented dataset via FedDCA with different cluster numbers $\xi$ on the four domains.
Ablation Study. We conduct the experiment with FedDCA w/o similarity threshold $\alpha$ on the four domains based on the FedAvg FL strategy. The performance and the corresponding domain coverage are shown in Table 10, where FedDCA without the similarity threshold $\alpha$ is marked as FedDCA. The result shows that the performance of FedDCA with similarity threshold $\alpha$ is slightly better than FedDCA without using $\alpha$ in code, financial, and mathematical domains, as the similarity scores in the medical domain are relatively lower. We show the similarity score distribution of the four domains in Figure LABEL:fig:similarity_score. For each domain, we plot each similarity score’s distribution of 10 clients. The similarity score is computed between the the selected client center and the public data, then we show the similarity score distribution using the histogram plot.
Table 10: Ablation study on the performance and domain coverage of FedDCA w/o similarity threshold $\alpha$.
Metric Performance(%) Domain Coverage
FedDCA FedDCA FedDCA FedDCA
H-Eval 35.97 36.58 0.8972 0.9348
M-Med 73.40 73.40 0.9348 0.9348
FPB 66.25 67.24 0.9353 0.9815
FiQA 23.27 35.27
TFNS 69.34 73.32
GSM8K 51.78 52.46 0.9128 0.9320
A.10 Augmentation strategy visualization
To more intuitively compare the domain coverage of different instruction augmentation methods, we randomly sampled 5,000 instructions obtained through these methods and randomly sampled 10,000 in-domain instructions as the background, which represents the distribution of specific domains in the public dataset. We then visualized the results using t-SNE(van der Maaten & Hinton (2008)), as shown in Figure LABEL:fig:augmentation_strategy_visualization. The plot shows that FedDCA encompasses the most of in-domain data, which is consistent with FedDCA’s domain coverage of each domain shown in Table 3. Also, we can observe that the random sampling strategy selects a lot of out-of-domain data while does not have good coverage in specific domains.
A.11 Frequently Used Notation
Table 11: Frequently used notation.
NOTATIONS REMARK
$N,C$ Clients number, client set $C=\{c_{1},c_{2},\ldots,c_{N}\}$.
$D^{p},D^{d},D^{c},D^{l}$ The public datasets, the in-domain data, the cross-client dataset, client’s local private data.
$N_{k}^{l},N_{k}^{p}$ Number of local private data on the $k$-th client, number of the retrieved public data on the $k$-th client.
$D_{k}^{l},D_{k}^{g},D_{k}$ The local private data on the $k$-th client, the retrieved public data on the $k$-th client, the augmented dataset on the $k$-th client.
$\Lambda,\mathcal{P},\mathcal{C}$ A specific sampling strategy that performs instruction augmentation on the server side, selected client center set, the cluster centers obtained locally and sent to the server for the greedy client center selection.
$\phi,\Delta\phi$ LLM’s pre-trained parameters, additional LoRA parameters.
$w,w_{enc},w_{p}$ Merged model parameters from the frozen LLM’s parameters $\phi$ and the additional LoRA parameters $\Delta\phi$, encoder model, projector model.
$F_{k}(w;\mathcal{D}),l(w;x,y)$ Loss of model $w$ over a specific dataset $\mathcal{D}$, the instructon tuning loss of model $w$ over a data sample $\left(x,y\right)$.

### v3

@@
 Abstract
-Federated Domain-specific Instruction Tuning (FedDIT) leverages a few cross-client private data and server-side public data for instruction augmentation, enhancing model performance in specific domains.
-While the factors affecting FedDIT remain unclear and existing instruction augmentation methods mainly focus on the centralized setting without considering the distributed environment.
-Firstly, our experiments show that cross-client domain coverage, rather than data heterogeneity, drives model performance in FedDIT.
-Thus, we propose FedDCA, which maximizes domain coverage through greedy client center selection and retrieval-based augmentation.
-To reduce client-side computation, FedDCA uses heterogeneous encoders with server-side feature alignment.
-Extensive experiments across four domains (code, medical, financial, and mathematical) validate the effectiveness of both methods.
-Additionally, we explore the privacy protection against memory extraction attacks with various amounts of public data and results show that there is no significant correlation between the amount of public data and the privacy-preserving capability.
-However, as the fine-tuning round increases, the risk of privacy leakage reduces or converges.
+Federated Domain-specific Instruction Tuning (FedDIT) utilizes limited cross-client private data together with server-side public data for instruction augmentation, ultimately boosting model performance within specific domains.
+To date, the factors affecting FedDIT remain unclear, and existing instruction augmentation methods primarily focus on the centralized setting without considering distributed environments.
+Our experiments reveal that the cross-client domain coverage, rather than data heterogeneity, drives model performance in FedDIT.
+In response, we propose FedDCA, which optimizes domain coverage through greedy client center selection and retrieval-based augmentation.
+For client-side computational efficiency and system scalability, FedDCA, the variant of FedDCA, utilizes heterogeneous encoders with server-side feature alignment.
+Extensive experiments across four distinct domains (code, medical, financial, and mathematical) substantiate the effectiveness of both methods.
+Additionally, we investigate privacy preservation against memory extraction attacks utilizing various amounts of public data.
+Results show that there is no significant correlation between the volume of public data and the privacy-preserving capability.
+However, as the fine-tuning rounds increase, the risk of privacy leakage reduces or converges.
 1 Introduction
-Table 1: Performance(%) of different augmentation settings on each domain, which are conducted via FedAvg protocol with 10 clients.
-Specifically, the zero-shot directly inferences without fine-tuning, while base data utilizes only the client’s local data for FedDIT.
-Also, we compare FedDCA with other two augmentation strategies: random sampling and direct retrieval (described in Appendix A.3), respectively.
+Table 1: Performance(%) of different augmentation settings on each domain, conducted via FedAvg protocol with 10 clients.
+Specifically, the zero-shot directly inferences without any fine-tuning, while the base data method utilizes only the client’s local data for FedDIT.
+Additionally, we compare FedDCA with other two augmentation strategies: random sampling and direct retrieval (described in Appendix A.3), respectively.
 Metric Zero-shot Base Data Random Sampling Direct Retrieval FedDCA
@@
 Avg. 42.91 48.99 49.07 51.71 56.56
-Recently, federated instruction tuning (FedIT) has gained attention as a novel approach that leverages the principles of federated learning (FL) to enable collaborative training of large language models (LLM) in a distributed environment without the need to share private data (McMahan et al.
-(2017); Ye et al.
-(2024b)).
-This approach involves exchanging model parameters among distributed data holders, striking a delicate balance between preserving privacy and optimizing model efficiency.
-Although many FedIT frameworks have been established(Ye et al.
-(2024b); Kuang et al.
-(2023); Zhang et al.
-(2024a)), they have not deeply explored the practical issues that FedDIT might encounter in real-world scenarios.
-For example, FedIT usually requires sufficient instruction data for fine-tuning, which is difficult to meet in domain-specific fine-tuning scenarios (Zhang et al.
-(2024d)).
-In this study, we explore Federated Domain-specific Instruction Tuning (FedDIT), a novel approach within the FL paradigm that focuses on enhancing the performance of LLMs in specific domains.
-Unlike the general FedIT, which seeks to improve model performance across a variety of tasks without considering local data shortage, FedDIT faces the challenge of clients having only a few local domain-specific data.
-To overcome this, FedDIT leverages a server-hosted public dataset that spans multiple domains to enrich the local data through the specific instruction augmentation strategy.
-This strategic enrichment is essential for achieving effective instruction tuning and needs to be carefully designed to prevent performance degradation.
-Except for the code domain, which is more about following a certain paradigm.
-As shown in Table 1, when clients attempt to do FedDIT only based on their local data, although some high-quality in-domain local data may be available, the limited scale restricts effective instruction tuning.
-Consequently, this sometimes leads to performance degradation, as indicated by the underlined values.
-Overall, the goal of FedDIT is to develop a domain-specific LLM that utilizes collaborative training and instruction augmentation with public data while preserving client privacy, ensuring that the model is adept at handling tasks within its targeted domain.
+Recently, federated instruction tuning (FedIT) has gained attention as a novel approach that leverages the principles of federated learning (FL) to facilitate collaborative training of large language models (LLM) in distributed environments while maintaining the confidentiality of private data (McMahan et al., 2017; Ye et al., 2024b).
+This methodology allows for the exchange of model parameters among distributed data holders, thereby achieving a careful balance between privacy preservation and efficient model optimization.
+Despite the establishment of various FedIT frameworks (Ye et al., 2024b; Kuang et al., 2023; Zhang et al., 2024a), existing literature has not adequately addressed the practical challenges that Federated Domain-specific Instruction Tuning (FedDIT) may encounter in real-world applications.
+For instance, FedIT generally necessitates a sufficient amount of instruction data for fine-tuning, which is often a shortage in domain-specific fine-tuning contexts (Zhang et al., 2024d).
+In this study, we investigate FedDIT, a novel approach within the FL paradigm aimed at boosting the performance of LLMs in specific domains.
+Unlike general FedIT, which seeks to enhance model effectiveness across diverse tasks without accounting for local data shortage, FedDIT encounters the unique challenge of clients possessing only a limited quantity of local domain-specific data.
+To overcome this, FedDIT leverages a server-hosted public dataset that spans multiple domains to enrich the local data through a specific instruction augmentation strategy.
+This strategic enrichment is crucial for achieving effective instruction tuning and needs to be meticulously designed to avoid performance degradation.
+Except for the code domain, which primarily adheres to a standardized paradigm, our results reveal that when clients rely solely on their local data for FedDIT, even the presence of high-quality in-domain local data can be insufficient due to limited scale, leading to a decline in performance, as reflected in the underlined values in Table 1.
+In summary, the goal of FedDIT is to develop a domain-specific LLM that employs collaborative training and instruction augmentation with public data while safeguarding client privacy, thereby ensuring the model’s proficiency in executing tasks pertinent to its designated domain.
 Additionally, the factors affecting FedDIT are still unclear.
-What’s worse, augmenting instructions may make the results even more unpredictable and confused about how to improve.
-Shepherd(Zhang et al.
-(2024a)) approaches this problem from the perspective of heterogeneity, constructing heterogeneity based on the topic of general instruction datasets.
-The study demonstrates that, for general FedIT, unlike the consensus of traditional FL, heterogeneity has a positive effect.
+Compounding this uncertainty, introducing augmented instructions may further complicate results, making it difficult to ascertain effective improvement strategies.
+Shepherd (Zhang et al., 2024a) approaches this problem from the perspective of heterogeneity, constructing heterogeneity based on the topic of general instruction datasets.
+It demonstrates that, unlike the consensus of traditional FL, for general FedIT, heterogeneity has a positive effect.
 By aggregating diverse instructions from clients, the diversity increases, thereby enhancing the model’s adaptability to various tasks.
 However, it just scratches the surface and does not explore issues in FedDIT.
-Going one step further, we conduct experiments to unveil a new key finding (Appendix A.2): there is no linear correlation between the degree of non-independent and identically distributed (non-iid) and the performance of LLM for FedDIT.
-Inspired by Explore-Instruct(Wan et al.
-(2023)), we first define the domain coverage metric and investigate the effect of domain coverage.
-Results show that the cross-client domain coverage impacts the model’s performance significantly.
-To maximize the cross-client domain coverage without compromising client data privacy, based on INSTA(Lee et al.
-(2024)), we propose a novel FedDIT algorithm, called FedDCA (Federated Instruction Tuning of LLMs with Domain Coverage Augmentation), which performs greedy client center selection and instruction augmentation via dense retrieval on the server side.
-The core idea of FedDCA lies in greedy client center selection, which chooses client centers that enhance the diversity and coverage of the domain-specific characteristics within the public dataset.
-By concentrating on maximizing domain coverage at each step, the algorithm efficiently builds an augmented train set that boosts the model’s learning and generalization capabilities, resulting in superior performance on domain-specific tasks.
-Furthermore, to reduce the computation overhead on the client side, we propose FedDCA with heterogeneous encoders.
-For feature alignment, we train a projector on the server side using public data through contrastive learning.
-We demonstrate the effectiveness of FedDCA by conducting extensive experiments across four domains: code, medical, financial, and mathematical, and comparing against multiple baselines, which can be categorized into two types: unaugmented and augmented methods.
-For the unaugmented setting, we compare our method with FedIT, which composes of four orthodox FL methods, including FedAvg(McMahan et al.
-(2017)), FedProx(Li et al.
-(2020)), SCAFFOLD(pmlr-v119-karimireddy20a), and FedAvgM(Hsu et al.
-(2019)).
-For the augmented setting, we compare our method with random sampling, direct retrieval, LESS(Xia et al.
-(2024)), and Self-Instruction(Wang et al.
-(2023c)).
-Additionally, we report FedDCA’s performance with different FL strategies.
-Furthermore, we compare the computational efficiency of FedDCA and FedDCA on the client side.
-To illustrate the impact of different retrieval volumes on client data privacy, we conduct experiments against memory extraction attacks.
-Results indicate that as the number of training rounds increases, the method only based on the client’s local data memorizes more privacy information while the privacy leakage risk gradually decreases or converges for the augmented setting.
+Going one step further, we conduct experiments to unveil a significant finding (Appendix A.2): there is no linear correlation between the degree of non-independent and identically distributed (non-iid) and LLM’s performance in the context of FedDIT.
+Inspired by Explore-Instruct (Wan et al., 2023), which shows the potential of domain coverage in domain-specific instruction tuning.
+The cross-client domain coverage metric is initially defined, followed by an investigation into its impact on FedDIT.
+Results demonstrate that domain coverage significantly influences model performance in the corresponding domain.
+To maximize the cross-client domain coverage without compromising client data privacy, we propose a novel FedDIT algorithm, Federated Instruction Tuning of LLMs with Domain Coverage Augmentation, termed FedDCA.
+This algorithm employs a greedy client center selection process and implements instruction augmentation through dense retrieval on the server side.
+The fundamental idea of FedDCA is to select client centers to expand the diversity and coverage of augmented instruction datasets within a specific domain.
+By strategically optimizing domain coverage at each step, FedDCA efficiently constructs the augmented train set that enhances both the learning and generalization capabilities of the model, leading to superior performance on domain-specific tasks.
+Furthermore, to mitigate computational overhead on the client side and enhance the system scalability, we propose FedDCA, which employs heterogeneous encoders of different sizes and capacities.
+To achieve feature alignment, we train a projector on the server side using public data and employ contrastive learning techniques.
+We demonstrate the effectiveness of FedDCA through comprehensive experiments conducted across four domains: code, medical, financial, and mathematical.
+These are compared against a range of baselines, which can be categorized into unaugmented and augmented methods.
+In the unaugmented setting, our method is compared with FedIT, which includes four orthodox FL techniques: FedAvg (McMahan et al., 2017), FedProx (Li et al., 2020), SCAFFOLD (pmlr-v119-karimireddy20a), and FedAvgM (Hsu et al., 2019).
+In the augmented setting, we compare FedDCA against methods such as random sampling, direct retrieval, LESS (Xia et al., 2024), and Self-Instruction (Wang et al., 2023c).
+Additionally, we present the performance outcomes of FedDCA when applied under various FL strategies.
+We also compare the computational efficiency on the client side between FedDCA and its variant, FedDCA.
+For privacy analysis, experiments against memory extraction attacks are conducted to evaluate how different quantities of retrieved public data affect the privacy of client local data.
+Results indicate that while reliance on local data increases memorization of sensitive information, the risk of privacy leakage diminishes or converges in the augmented setting as the training rounds progress.
 The main contribution is as follows:
-  • We unveil a key finding: For FedDIT, there is no linear correlation between the degree of heterogeneity and the model’s performance.
-Instead, cross-client domain coverage impacts the model’s performance significantly, which is presented in Appendix A.2.
-  • We propose a novel FedDIT algorithm (Section 4) called FedDCA to maximize the cross-client domain coverage through greedy client center selection and then perform retrieval-based instruction augmentation on the server.
-To further reduce the computational overhead on the client side, we propose FedDCA with a heterogeneous encoder setting.
-For feature alignment, we train a projector on the server side.
-  • With extensive experiments (Section 5), we show the effectiveness of FedDCA and FedDCA.
-Furthermore, we explore privacy preservation against memory extraction attacks and conduct experiments based on various public data amounts.
-Results indicate that the privacy-preserving capability is not significantly correlated with the amount of public data, but the risk of privacy leakage decreases or converges as the amount of fine-tuning rounds increases.
+  • We reveal a critical finding: in the context of FedDIT, there exists no linear correlation between the degree of heterogeneity and model performance.
+Instead, cross-client domain coverage substantially impacts LLM’s effectiveness, as elaborated in Appendix A.2.
+  • We propose a novel FedDIT algorithm (Section 4), termed FedDCA, aimed at maximizing cross-client domain coverage through greedy client center selection followed by retrieval-based instruction augmentation executed on the server.
+Additionally, we introduce FedDCA to further lessen the client-side computational overhead while enhancing the system scalability.
+This variant utilizes a heterogeneous encoder structure, paired with a projector on the server side for feature alignment.
+  • Through extensive experiments (Section 5), we demonstrate the effectiveness of FedDCA and FedDCA.
+We also investigate privacy preservation against memory extraction attacks, conducting experiments based on various amounts of public data.
+Results suggest that the capacity for privacy preservation does not correlate significantly with the quantity of public data.
+In contrast, the risk of privacy leakage tends to decrease or converge as the fine-tuning rounds increase.
 2 Preliminaries
-Federated domain-specific instruction tuning (FedDIT) aims to leverage cross-client private domain-specific instruction data and utilize the multi-domain public data on the server to achieve instruction augmentation, collaboratively enhancing the model’s performance in specific domains.
-Suppose $N$ distributed clients with their local private data $D_{k}^{l}$.
-The server has a public dataset $D^{p}$ consisting of multiple domains and is responsible for data augmentation based on different strategies and aggregation model parameters from clients.
-The training process follows standard FL protocol(McMahan et al.
-(2017)).
-For computational efficiency, we adopt LoRA(Hu et al.
-(2021)) as the fine-tuning method, which involves tuning additional parameters $\Delta\phi$ while keeping the pre-trained LLM’s parameters $\phi$ frozen.
-In the initial training stage, the server dispatches the pre-trained LLM parameters $\phi$ and the initialized LoRA parameters $\Delta\phi$ to each client for training $\mathcal{R}$ rounds.
-In the $t$-th round, the server sends the aggregated LoRA parameters $\Delta\phi^{t}$ to clients.
-Clients use $\Delta\phi^{t}$ to update their local LoRA parameters $\Delta\phi_{k}^{t}$ and perform instruction tuning based on augmented instruction datasets, then send the updated LoRA parameters $\Delta\phi_{k}^{t+1}$ back to the server.
-The server receives the LoRA parameters $\{\Delta\phi_{k}^{t+1}\mid k=1,\dots,N\}$ from each client and aggregates them to obtain the global LoRA parameters $\Delta\phi^{t+1}$ for the next round.
+FedDIT aims to leverage cross-client private domain-specific instruction data and utilize the multi-domain public data on the server to achieve instruction augmentation, collaboratively enhancing the model’s performance in specific domains.
+Consider $N$ distributed clients, each with local private data $D_{k}^{l}$.
+The server maintains a public dataset $D^{p}$ that encompasses multiple domains and is responsible for implementing data augmentation strategies and aggregating model parameters received from clients.
+The training process follows the standard FL protocol (McMahan et al., 2017).
+For computational efficiency, we adopt Low-Rank Adaption (LoRA) (Hu et al., 2021) as the fine-tuning method, which involves tuning additional parameters $\Delta\phi$ while keeping the pre-trained LLM’s parameters $\phi$ frozen.
+In the initial training phase, the server dispatches $\phi$ and $\Delta\phi$ to each client for $\mathcal{R}$ training rounds.
+In the $t$-th round, the server sends the aggregated $\Delta\phi^{t}$ to clients.
+Clients use $\Delta\phi^{t}$ to update their local LoRA parameters $\Delta\phi_{k}^{t}$ and conduct instruction tuning based on augmented instruction datasets $D_{k}$.
+Subsequently, the clients return $\Delta\phi_{k}^{t+1}$ to the server.
+The server then aggregates $\{\Delta\phi_{k}^{t+1}\mid k=1,\dots,N\}$ to obtain $\Delta\phi^{t+1}$ for the next round.
 3 Problem Formulation
 For better understanding, we list the frequently used notation in Appendix A.11.
-The objective of FedDIT is to enhance the domain-specific performance of LLMs through FL without sharing private data(Ye et al.
-(2024b); Zhang et al.
-(2024d)).
+The objective of FedDIT is to enhance the domain-specific performance of LLMs through FL without sharing private data (Ye et al., 2024b; Zhang et al., 2024d).
 Under the FL framework, suppose we have $N$ clients, where each client $c_{k}$ has a local dataset $D_{k}^{l}$ with its size $N_{k}^{l}$, and an augmented dataset $D_{k}^{g}$ from the server public dataset $D^{p}$, respectively.
@@
 Denote the augmentation strategy as $\Lambda$, through which the server performs instruction augmentation on the public dataset.
-If $\Lambda$ is null, it denotes clients only perform FedDIT based only on their local private data, which may cause performance degradation.
-Otherwise, $\Lambda$ could be categorized into two types: 1) Focusing only on its own local data distribution.
-2) Taking into account the cross-client data distribution.
+If $\Lambda$ is null, it indicates that clients conduct FedDIT solely based on their local private data, which may lead to performance degradation.
+Conversely, $\Lambda$ may be classified into two categories: (1) focusing exclusively on the client’s own local data distribution or (2) considering the cross-client data distribution.
 In conclusion, the global objective of FedDIT is defined as follows:
 $\operatorname*{arg\,min}_{\Delta\phi}{\left\{F(\phi,\Delta\phi)\triangleq\sum_{k=1}^{N}p_{k}\left((1-\alpha_{k})F_{k}\left(\phi,\Delta\phi_{k};D_{k}^{l}\right)+\alpha_{k}F_{k}\left(\phi,\Delta\phi_{k};D_{k}^{g}\right)\right)\right\}},$ (1)
-where the first term and second term denote the accumulated fine-tuning loss computed on the client $c_{k}$’s local instructions $D_{k}^{l}$ and the augmented public instructions $D_{k}^{g}$ from the server respectively.
+where the first term and second term denote the accumulated fine-tuning loss computed on the client $c_{k}$’s local instructions $D_{k}^{l}$ and the augmented public instructions $D_{k}^{g}$ from the server, respectively.
 $p_{k}$ is the weight of the $k$-th client, which is determined by the ratio of $k$-th client’s data size to all clients’ total data size.
@@
 4 Method
-We propose Federated Instruction Tuning of LLMs with Domain Coverage Augmentation (FedDCA), which enhances domain coverage through greedy client center selection and performs retrieval-based instruction augmentation to obtain a LLM that performs well on domain-specific tasks.
-We first give the overview FedDCA (Section 4.1) and define the domain coverage metric (Section 4.2), then show details of the greedy client center selection algorithm (Section 4.3) and heterogeneous encoder method FedDCA (Section 4.4).
-4.1 Overview
+We propose Federated Instruction Tuning of LLMs with Domain Coverage Augmentation (FedDCA), which enhances domain coverage to obtain a LLM that performs well on domain-specific tasks.
+FedDCA follows the standard FedIT protocol(Ye et al., 2024b) to perform federated instruction tuning.
+The novel design of FedDCA lies in the phase before training, which consists of two key modules: greedy client center selection and domain data retrieval, which are elaborated in the following; also see details in Algorithm 1 and Figure 1.
+For computational efficiency and system scalability, we introduce a variant of FedDCA with heterogeneous encoder setting, named FedDCA.
+We first formulate the optimization problem to solve for FedDCA.
 Figure 1: Overview of FedDCA, which consists of three stages: 1) The client $c_{k}$ performs local instructions clustering and sends cluster centers $\mathcal{C}_{k}$ to the server.
 2) The server first does greedy client center selection to maximize domain coverage and performs client-center-based domain data retrieval, then sends the augmented instructions $D_{k}^{g}$ to the client $c_{k}$.
-3) Clients fine-tune the LLM collaboratively and the LoRA parameters $\Delta\phi$ are exchanged between clients and the server.
-As shown in Figure 1, FedDCA consists of three stages, in the first stage the client $c_{k}$ clusters its local instructions and sends the cluster centers $\mathcal{C}_{k}$ to the server.
-Then the server performs greedy client center selection to obtain client centers that maximize the domain coverage, then performs client-center-based retrieval for domain-specific instruction augmentation.
-Additionally, the third stage follows the standard FedIT procedure(Ye et al.
-(2024b)), in which the server sends the frozen pre-trained LLM parameters $\phi$ and the initialized LoRA parameters $\Delta\phi$ to clients and then $\Delta\phi_{k}$ is exchanged between the client and the server.
-The novel design of FedDCA lies in the second stage before training.
-As shown in Algorithm 1, the core idea is to use a greedy client center selection algorithm to obtain client centers that maximize cross-client domain coverage, and then use these client centers to retrieve instructions on the public dataset.
-To further reduce the computational overhead of encoding on clients, we propose a heterogeneous encoder method called FedDCA.
-Specifically, the client uses a small encoder locally, while the server uses a larger encoder for accurate semantics.
-To align features with different dimensions, we introduce a projector $w_{p}$ on the server side, which is obtained through contrastive learning.
+3) Clients fine-tune the LLM collaboratively, and the LoRA parameters $\Delta\phi$ are exchanged between clients and the server.
 Algorithm 1 FedDCA: greedy client center selection & domain data retrieval
@@
 21:end for
-4.2 Domain Coverage
-As domain coverage directly affects the in-domain performance of the LLM (Appendix A.2), FedDCA aims to maximize the domain coverage of the cross-client data distribution $\cup_{i=1}^{N}D_{i}$ respect to the in-domain data distribution $D^{d}$, which is defined in Eq.
+4.1 Optimization Problem
+As domain coverage directly affects the in-domain performance of the LLM (Appendix A.2), FedDCA aims to maximize the domain coverage of the cross-client augmented data $\cup_{i=1}^{N}D_{i}$ respect to the in-domain data distribution $D^{d}$, which is defined in Eq.
 5.
-However, as public data is unreachable for clients and clients can not send the local data to the server for privacy, directly finding a cross-client dataset that maximizes the domain coverage in Eq.
+However, as clients can not send the local data to the server for privacy, directly finding a cross-client dataset that maximizes the domain coverage in Eq.
 5 is unrealistic.
-To find a proper client center set $\mathcal{P}$ that attempts to maximize the domain coverage $d(D^{d},P)$ and use $\mathcal{P}$ to perform instruction retrieval on the public data is an approximation problem.
+To find a proper client center set $\mathcal{P}$ that maximizes domain coverage and uses it to perform instruction retrieval on the public data is an approximation problem.
 Suppose there exists a metric space $\mathcal{X}$, a set of cross-client local instruction embeddings $\mathcal{E}\in\mathcal{X}$, and a set of cluster centers $\mathcal{P}\in\mathcal{X}$.
-Each client $c_{k}$ has maximum $\xi$ clusters obtained by k-means algorithm(Vardakas & Likas (2023)).
-As in the federated setting, communication cost is always considered.
-Thus, we need to find a $\mathcal{P}$ that maximizes the domain coverage while considering the communication cost, we formulate the optimization problem as follows:
+Each client $c_{k}$ has maximum $\xi$ clusters obtained by k-means algorithm (Vardakas & Likas, 2023).
+In the federated setting, communication cost is always a critical factor.
+Consequently, we formulate the optimization problem to include communication costs as follows:
 $\operatorname*{arg\,min}_{\mathcal{P}}\left\{\sum_{i=1}^{N}|\mathcal{P}_{i}|+\sum_{d\in D^{d}}\left(\text{min}_{p\in\mathcal{P}}sim(d,p)\right)\right\},$ (3)
@@
 We will consider this problem from two aspects: 1) Select a client center that can represent the distribution of the local data.
-2) To maximize the cross-client domain coverage, we filter client centers that are close to the previously selected client centers.
-Firstly, we randomly choose a client and select the largest cluster center as its center and be the initial of the client center set $\mathcal{P}$.
-Then we iteratively select the client center $\mathcal{P}_{k}$ of the top cluster size, while avoiding selecting the cluster center close to the previously selected client center to maximize the domain coverage of $\mathcal{P}$.
+2) To optimize the cross-client domain coverage, we filter client centers that are close to the previously selected client centers.
+First, we randomly choose a client and select the largest cluster center as its center and be the initial of the client center set $\mathcal{P}$.
+Then, we iteratively select the client center $\mathcal{P}_{k}$ of the top cluster size while avoiding selecting the cluster center close to the previously selected client center to maximize the domain coverage of $\mathcal{P}$.
 Specifically, for the $i$-th iteration, we filter $i$-top cluster centers that have the largest summed similarity with previously selected client centers and select the center of the rest largest cluster as the $i$-th client center.
-This procedure is repeated until have selected $N$ client centers.
+This procedure is repeated until we have selected $N$ client centers.
 Domain data retrieval.
-For each client center $\mathcal{P}_{k}$, the server performs dense retrieval (detailed in Appendix A.3) on public dataset $D^{p}$ to get the top-$N_{k}^{p}$ similar public instructions, then sends $\{D_{1}^{g},\dots,D_{N}^{g}\}$ to each clients.
+For each client center $\mathcal{P}_{k}$, the server performs dense retrieval (detailed in Appendix A.3) on public dataset $D^{p}$ to get the top-$N_{k}^{p}$ similar public instructions, then sends retrieved public datasets $\{D_{1}^{g},\dots,D_{N}^{g}\}$ to each clients.
 Specifically, to avoid the overlap between public data and local private data, we set a threshold $\alpha$ to filter the public instructions that have a similarity score larger than $\alpha$ with the client center.
-In summary, FedDCA can be seen as a method for establishing a comprehensive training set that captures the essence of the domain and covers its distribution.
-Through the iterative selection of client centers that provide unique coverage, the algorithm ensures that the selected centers are representative of the domain and achieve great domain coverage.
-In addition, domain data retrieval exposes the model to a broader range of in-domain features, facilitating a more robust understanding of domain-specific tasks.
-This not only enriches the training data but also reduces the risks of overfitting to the limited local data available to each client, which ends up with performance degradation.
-By leveraging diverse examples across clients, the model can better generalize its learning, leading to enhanced performance on various tasks within the domain.
+In summary, FedDCA establishes a comprehensive training set that captures the essence and distribution of the domain by iteratively selecting client centers with unique coverage increments.
+This ensures that selected centers are representative and achieve broad domain coverage.
+Additionally, domain data retrieval exposes the model to a wider range of in-domain features, enhancing its understanding of domain-specific tasks.
+This not only enriches the training data but also reduces the risks of overfitting to the limited local data available to each client, ultimately improving performance across various tasks within the domain.
 Figure 2: Greedy client center selection iteratively selects the client center in each step, which synthetically considers both the representativeness of the cluster center and its distance from the previously selected client center set $\mathcal{P}$ to maximize the cross-client domain coverage.
 4.3 Heterogeneous Encoder with Feature Alignment
-Due to the limited computational resources on the client side, we propose a heterogeneous encoder method FedDCA to reduce the client’s computational overhead by using a small encoder $\omega$ on the client side and a larger encoder $\omega^{\prime}$ on the server side.
+For the client-side computational efficiency and system scalability, we propose a heterogeneous encoder method FedDCA to reduce the client’s computational overhead by using a small encoder $\omega$ on the client side and a larger encoder $\omega^{\prime}$ on the server side.
 However, the output dimension of heterogeneous encoders may not be consistent.
-Therefore, we introduce a projector $w_{p}$ on the server for feature alignment, as shown in Appendix A.4, we use contrastive learning(Khosla et al.
-(2020)) and train $w_{p}$ on the public instructions.
-Specifically, $w_{p}$ is composed of two fully connected layers and a ReLU activation layer in between, projecting $\omega$’s dimension to the output dimension of $\omega^{\prime}$.
+Therefore, we introduce a projector $w_{p}$ on the server for feature alignment, as shown in Appendix A.4.
+We use contrastive learning (Khosla et al., 2020) and train $w_{p}$ on the public instructions.
+Specifically, $w_{p}$ comprises two fully connected layers and a ReLU activation layer in between, projecting $\omega$’s dimension to the output dimension of $\omega^{\prime}$.
 For $\forall{x_{i}}\in D^{p}$, where $i\in\{1,2,\dots,\mathcal{N}\}$, $\omega$ outputs embedding $h_{i}$, $\omega^{\prime}$ outputs embedding $h_{i}^{\prime}$ and $w_{p}$ outputs embedding $\varphi_{i}$.
@@
 where $\text{sim}(\cdot,\cdot)$ denotes the cosine similarity and $\tau$ denotes the temperature parameter.
-Using a projector $w_{p}$ to achieve feature alignment is an effective strategy in FL for scenarios where clients and the server utilize encoders of different sizes or capacities.
-The projector maps the lower-dimensional feature representations $\mathcal{H}$ from the client-side encoder to the higher-dimensional space $\mathcal{H}^{\prime}$ of the server-side encoder.
-By presenting the projector with positive and negative pairs, $w_{p}$ learns to adjust the projections such that the feature representations of similar items are close to each other, while those of dissimilar items are distant.
-In conclusion, the heterogenous encoder setting allows the system to be more scalable and flexible based on different client capacities.
-As a result, the projector helps in creating a unified feature space where representations from diverse sources can be combined, and utilized collaboratively.
+The heterogeneous encoder setting, by employing a projector $w_{p}$ for feature alignment, proves effective in FL contexts where clients and the server use different encoder sizes or capacities.
+This approach facilitates mapping lower-dimensional client features $\mathcal{H}$ to the higher-dimensional server space $\mathcal{H}^{\prime}$ through a strategic training of $w_{p}$ with both positive and negative pairs.
+It adeptly aligns similar feature representations closely while positioning dissimilar ones further apart.
+As a result, the heterogeneous encoder setting not only enhances system scalability and flexibility but also facilitates the creation of a unified feature space that effectively integrates and utilizes diverse representations collaboratively.
 4.4 Discussions
@@
 Through greedy client center selection, FedDCA solves the optimization problem defined in Eq.
-3 in the polynomial time and quite efficient.
-On the client side, the k-means algorithm(Vardakas & Likas (2023)) is $\mathcal{O}(|D_{k}^{l}|)$.
+3 in the polynomial time and is quite efficient.
+On the client side, the k-means algorithm (Vardakas & Likas, 2023) is $\mathcal{O}(|D_{k}^{l}|)$.
 On the server side, as mentioned in Section 4.1, the time complexity of greedy client center selection is $\mathcal{O}\left(N\left(N\xi+\xi\log\xi\right)\right)$.
 Specifically, for each client, the similarity computation between selected client center set $\mathcal{P}$ and client $c_{k}$’s cluster $\mathcal{C}_{k}$ is $\mathcal{O}\left(N\xi\right)$, and the sorting process is $\mathcal{O}\left(\xi\log\xi\right)$.
-For domain data retrieval, given that the sorting algorithm is $\mathcal{O}\left(\mathcal{N}\log\mathcal{N}\right)$, where $\mathcal{N}$ is the size of the public dataset, thus the time complexity is $\mathcal{O}\left(N(\mathcal{N}\log\mathcal{N})\right)$.
-In addition, the hetergenous encoder setting FedDCA further reduces the computation overhead on the client side.
+For domain data retrieval, given that the sorting algorithm is $\mathcal{O}\left(\mathcal{N}\log\mathcal{N}\right)$, where $\mathcal{N}$ is the size of the public dataset.
+Thus the time complexity is $\mathcal{O}\left(N(\mathcal{N}\log\mathcal{N})\right)$.
+In addition, the heterogeneous encoder setting FedDCA further reduces the computation overhead on the client side.
 Communication.
 In the domain instruction augmentation stage, each client sends $\xi$ cluster centers to the server.
-Next, the server performs greedy client center selection and domain data retrieval, then sends retrieved public data $D_{k}^{g}$ to the client $c_{k}$, which size is $N_{k}^{p}$.
+Next, the server performs greedy client center selection and domain data retrieval, then sends retrieved public data $D_{k}^{g}$ to the client $c_{k}$, whose size is $N_{k}^{p}$.
 In addition, in the fine-tuning stage, FedDCA follows the standard FedIT procedure, which exchanges the LoRA parameters $\Delta\phi_{k}$ between clients and the server.
-Specifically, for Llama3-8B model, the number of trainable LoRA parameters is just 13.6 million.
-Compared with the number of the frozen pretrained LLM parameters, the communication for LoRA tuning is quite efficient.
+Specifically, for the Llama3-8B model, the number of trainable LoRA parameters is just 13.6 million.
+Compared with the number of frozen pre-trained LLM parameters, the communication for LoRA tuning is quite efficient.
 Privacy.
-The proposed method FedDCA does not violate the client’s privacy.
-Compared with other FedIT methods(Zhang et al.
-(2024d); Ye et al.
-(2024b)), the difference lies in the greedy client center selection.
+The proposed method, FedDCA, does not violate the client’s privacy.
+Compared with other FedIT methods (Zhang et al., 2024d; Ye et al., 2024b), the difference lies in the greedy client center selection.
 In this stage, the client only uploads the cluster center to the server, which is the average of embeddings to its cluster.
-In addition, the potential privacy leakage can be further avoided through homomorphic encryption(Acar et al.
-(2018)), which allows the server directly computates on ciphertext for matrix multiplication for dense retrieval.
+In addition, the potential privacy leakage can be further avoided through homomorphic encryption (Acar et al., 2018), which allows the server to directly compute on ciphertext for matrix multiplication for dense retrieval.
 5 Experiments
-To show the effectiveness of FedDCA, we conduct extensive experiments on various domains and compare FedDCA with several baselines.
-This section is organized as follows: First we give details of experimental setup in Section 5.1, then we report the performance of FedDCA and compare it with several baselines in Section 5.2, and next we evaluate the computational cost between the FedDCA and FedDCA in Section 5.3.
-Finally, we do privacy evaluations of FedDCA in Section 5.4.
-For more details and results please see in Appendix A.
+To demonstrate the effectiveness of FedDCA and its variant FedDCA, we conduct extensive experiments across various domains and with several baselines.
+For additional details and results, please refer to Appendix A.
 5.1 Experimental Setup
 Dataset and Evaluation Metrics.
-To evaluate the performance of FedDCA, we conduct experiments on the four domains: code, medical, financial and mathematcial.
+To evaluate the performance of FedDCA, we conduct experiments on four domains: code, medical, financial, and mathematical.
 The detail of constructing the public dataset is described in Appendix A.5.
-We choose HumanEval (code, H-Eval for short)(Chen et al.
-(2021)), MMLU-Med (medical, abbreviated as M-Med)(Hendrycks et al.
-(2021)), FPB (financial), FiQA (financial), TFNS (financial)(Yang et al.
-(2023a)), GSM8K (mathematical)(Cobbe et al.
-(2021)) for evaluation.
+For evaluation, we select a range of datasets, including HumanEval (H-Eval for short) for coding (Chen et al., 2021), MMLU-Med (abbreviated as M-Med) for medical (Hendrycks et al., 2021), and GSM8K for mathematics (Cobbe et al., 2021).
+Additionally, financial datasets such as FPB, FiQA, and TFNS (Yang et al., 2023a) are utilized.
 HumanEval is evaluated using Pass@1.
-For MMLU-Med, FPB, FiQA and TFNS, we use accuracy as evaluation metric and exact match (EM) is used for GSM8K.
+For MMLU-Med, FPB, FiQA, and TFNS, we use accuracy as the evaluation metric.
+While GSM8K is evaluated using exact match (EM).
 Please see Appendix A.5 for more dataset details.
 Baselines.
-We compare FedDCA and FedDCA with the following baselines: 1) Unaugmented.
-Zero-shot inference and FedIT, which composes of four widely used FL methods, including FedAvg(McMahan et al.
-(2017)), FedProx(Li et al.
-(2020)), SCAFFOLD(pmlr-v119-karimireddy20a) and FedAvgM(Hsu et al.
-(2019)).
-2) Augmented.
-Random sampling, direct retrieval, LESS(Xia et al.
-(2024)), Self-Instruct(Wang et al.
-(2023c)).
+We compare FedDCA and FedDCA with the following baselines: 1) Unaugmented methods, including zero-shot inference and FedIT, which are composed of four widely used FL methods, including FedAvg (McMahan et al., 2017), FedProx (Li et al., 2020), SCAFFOLD (pmlr-v119-karimireddy20a) and FedAvgM (Hsu et al., 2019).
+2) Augmented methods, which include random sampling, direct retrieval, LESS (Xia et al., 2024), and Self-Instruct (Wang et al., 2023c).
 Additionally, we report FedDCA’s performance with different FL strategies in Table 2.
@@
 Direct retrieval is described in Appendix A.3.
-Self-Instruct augments instruction from a generative way through GPT-3.5-turbo(Sun et al.
-(2023)).
-Base on the prompt provided by Self-Instruct, we define the prompts for generating instructions and responses as shown in Appendix A.7.
+Self-Instruct augments instruction in a generative way through GPT-3.5-turbo (Sun et al., 2023).
+Based on the prompt provided by Self-Instruct, we define the prompts for generating instructions and responses as shown in Appendix A.7.
 5.2 Performance Analysis
@@
 In addition, to reduce the computation overhead of clients, FedDCA attempts to utilize heterogeneous encoders.
-We see that although FedDCA has a performance drop of 1.67% than FedDCA, it still outperforms other baselines in the average performance.
-Furthermore, we report FedDCA’s performance with different FL strategies, we can see that no FL methods can keep the leading position in all domains.
+We see that although FedDCA has a performance drop of 1.85% than FedDCA, it still outperforms other baselines in the average performance.
+Furthermore, we report FedDCA’s performance with different FL strategies, and we can see that no FL method can keep the leading position in all domains.
 In specific, FedProx and SCAFFOLD FL strategies perform better in the average performance.
@@
 Table 2: Performance(%) of FedDCA, FedDCA and other nine baselines on various domains.
-We report FedDCA with different FL strategies in the last four rows and FedDCA+FedProx achieves the best average performance across four domains.
+We report FedDCA with different FL strategies in the last four rows, and FedDCA+FedProx achieves the best average performance across four domains.
 Method H-Eval M-Med FPB FiQA TFNS GSM8K Avg.
@@
 FedDCA+FedAvgM 33.53 68.90 71.45 31.45 72.52 49.76 54.60
-Furthermore, Table 2 can be divided into two parts: unaugmented methods and augmented methods.
+Additionally, Table 2 can be divided into two parts: unaugmented methods and augmented methods.
 We can observe that FedIT methods perform well in the code domain, even better than most augmented methods.
@@
 2) As the local data is few, so compared with the same epoch and batch size, the unaugmented methods learn the same data more times, which is kind of not a fair setting.
-However, FedDCA+SCAFFOLD still surpasses the best baseline in the code domain, which further demonstrates the effectiveness of FedDCA.
+However, FedDCA+SCAFFOLD still surpasses the best baseline in the code domain, further demonstrating the effectiveness of FedDCA.
 Efficiency.
 We see that direct retrieval is the best baseline in the average performance.
-However, it does not consider the cross-client domain coverage, resulting in an overlapping retrieval data distribution.
-Self-Instruct(Wang et al.
-(2023c)) represents the upper limit performance of the generation-based method, which is restricted to the high expense of cost, limited API call frequency, and heavy quality screening process.
-On the other hand, FewFedPIT(Zhang et al.
-(2024d)) attempts to augment local data by leveraging the pre-trained Llama2 model(Li & Flanigan (2023)) to perform self-instruct manner instruction generation during training.
-However, there exist two concerns: 1) The pre-trained LLM cannot provide effective, stable, and high-quality instruction generation.
-2) Generating instructions locally is very costly in terms of both time and computational resources.
-Overall, generating instructions through a self-instruct manner is not a satisfactory method for the FedDIT scenario.
-LESS(Xia et al.
-(2024)) represents the augmentation methods through gradient feature retrieval, however, the performance of LESS is not satisfactory.
-What’s worse, it assumes success to the validation set which is not always available and requires a warmup process on the public dataset first and then computes the gradient of both the public data and the validation set, which is computationally expensive when the size of the public dataset is enormous.
-In conclusion, FedDCA achieves better performance while requiring lower computational resources and a more relaxed training condition, which shows the efficiency of FedDCA.
+However, it does not consider cross-client domain coverage, resulting in an overlapping retrieved data distribution.
+Self-Instruct (Wang et al., 2023c) represents the upper limit performance of the generation-based method, which is restricted to the high expense of cost, limited API call frequency, and heavy quality screening process.
+On the other hand, FewFedPIT (Zhang et al., 2024d) attempts to augment local data by leveraging the pre-trained Llama2 model to perform self-instructed generation during training.
+However, two concerns exist: 1) The pre-trained LLM cannot provide effective, stable, and high-quality instruction generation.
+2) Generating instructions locally is very costly regarding both time and computational resources.
+Overall, generating instructions self-instructively is not a satisfactory method for the FedDIT scenario.
+LESS (Xia et al., 2024) represents the augmentation methods through gradient feature retrieval.
+However, its performance is underwhelming.
+Compounding this issue, LESS presupposes access to the validation set, which is not always available, and necessitates an initial warmup on the public dataset before calculating gradients for both the public and validation data.
+This process becomes computationally burdensome with large public datasets.
+In conclusion, FedDCA achieves better performance while requiring lower computational resources and a more relaxed training condition, which shows its efficiency.
 Table 3: Domain coverage of FedDCA, FedDCA and other baselines on four domains, where FedIT methods only use local data without domain augmentation.
@@
 Domain coverage.
-The domain coverage of each method is shown in Table 3.
-Also, we show the visualization of the domain coverage across different instruction augmentation methods in Appendix A.10.
-For the augmented methods, it can be observed that the higher domain coverage in a particular domain correlates with the better performance of the method in that domain.
-Specifically, given that FedDCA aims to maximize the domain coverage through greedy client center selection, therefore it achieves the highest domain coverage in all four domains, which surpasses other baselines from 4.82% to 21.36% average relative improvement.
-In addition, as FedDCA utilizes a smaller encoder on the client side for computational efficiency, it sacrifices some semantic precision, resulting in a 2.89% averaged domain coverage drop in relative.
-However, FedDCA still surpasses the best baseline by 1.79% on average.
+Each method’s domain coverage is shown in Table 3.
+Additionally, visualization of domain coverage across different strategies can be found in Appendix A.10.
+For augmented methods, a correlation is evident between higher domain coverage in a specific domain and improved performance of the method within that domain.
+Given that FedDCA aims to maximize the domain coverage through greedy client center selection, therefore it achieves the highest domain coverage in all domains, which surpasses other baselines from 4.82% to 21.36% average relative improvement.
+Furthermore, while FedDCA employs a smaller encoder on the client side to enhance computational efficiency, this leads to a slight compromise in semantic precision, observed as a relative drop of 2.89% in average domain coverage.
+Nevertheless, FedDCA still outperforms the best baseline by an average of 2.78%.
 In brief, the result proves the effectiveness of FedDCA and FedDCA in domain coverage augmentation.
@@
 The encoders used for FedDCA and FedDCA on the client side are bge-large-en-v1.5 and all-MiniLM-L6-v2 respectively (detailed in Appendix A.6).
-Then we show the model size and the time overhead of encoding the public instructions in Table 4
-We can observe that in comparison with FedDCA, there is a large computational and time overhead advantage with FedDCA, while achieves an acceptable performance (shown in Table 2), which is accomplished by implementing a learnable projector on the server side.
-This projector maps the lower-dimensional client-side feature to the higher-dimensional space of the server-side encoder output.
-Consequently, the projector facilitates the creation of a unified feature space where representations from diverse sources can be integrated and used collaboratively.
+Then, we show the model size and the time overhead of encoding the public instructions in Table 4
+Compared to FedDCA, FedDCA has significant computational and time advantages while maintaining acceptable performance (shown in Table 2).
+This is accomplished by employing heterogeneous encoders and configuring a projector on the server to achieve dimensional mapping.
 5.4 Privacy Analysis
-Next, we evaluate the privacy-preserving capability of different ratios of public data against memory extraction attacks(Carlini et al.
-(2021); Zhang et al.
-(2024c)), which utilizes the autoregression nature of LLM to extract information from the memory(Xu et al.
-(2024)).
+Next, we evaluate the privacy-preserving capability of different ratios of public data against memory extraction attacks (Carlini et al., 2021; Zhang et al., 2024c), which utilizes the autoregression nature of LLM to extract information from the memory (Xu et al., 2024).
 Table 5: Privacy-preserving capability against memory extraction attack with different amounts of augmented public data. We report the average ROUGE-L score of each setting in four domains.
@@
 Evaluation setup.
-We focus on a single client’s fine-tuning process, which has 100 local data and adopts FedDCA to do the instruction augmentation with different amounts of public data from 1000 to 5000.
+We focus on one client’s instruction tuning in FedDIT, using FedDCA for instruction augmentation with 1000 to 5000 public instructions.
 We set up 10 clients with full participation for 10 rounds.
-Specifically, we record every round’s ROUGE-L score(Lin (2004)) of the client $c_{0}$.
-We use the designed prompt to extract the instruction that LLM memorizes, which is shown in Appdix A.8.
-Specifically, for each setting, we repeat the memory extraction 100 times and report the average ROUGE-L score computed based on each generated instruction and all the local data’s instructions to evaluate the privacy-preserving capability of different amounts of public data.
-Figure 3: ROUGE-L score of each pair of base-data-only and augmented fine-tuning in four domains per round, where different monochromatic represents each domain.
+Specifically, we record the average ROUGE-L score (Lin, 2004) of client $c_{0}$ for each round.
+The designed prompt to extract instructions memorized by the LLM is detailed in Appendix A.8.
+For each setting, we repeat memory extraction 100 times and report the average ROUGE-L score based on each generated instruction and all local data instructions to evaluate the privacy-preserving capability of different public data amounts.
+Specifically, denoting the generated $\mathcal{N}$ instructions as $\mathcal{I}$ and the client’s local instructions $\mathcal{I}^{l}$.
+The calculation is defined as $\frac{1}{\mathcal{N}}\sum_{i=1}^{\mathcal{N}}\texttt{ROUGE-L}(\mathcal{I}_{i},\mathcal{I}^{l})$.
+Figure 3: Average ROUGE-L score of each pair of base-data-only and augmented fine-tuning in four domains per round, where different monochromatic represents each domain.
 Results.
-We report the average ROUGE-L score of base-data-only fine-tuning and various amounts of public data fine-tuning in Table 5.
-The result shows that at the same training round, there is no significant correlation between the public data ratio and the privacy-preserving capability.
-In addition, after 10 rounds of FedDIT, the base-data-only fine-tuning’s average ROUGE-L score all exceeds the instruction-augmented settings in four domains.
-Specifically, base-data-only fine-tuning exceeds augmented settings by 6.69%, 5.55%, 5.55%, and 6.16% on average, relatively.
-The trend of the ROUGE-L score of base-data-only and augmented fine-tuning of each domain per round is shown in Figure 3.
-In specific, the augmented fine-tuning uses 5000 public data.
-We can see that the ROUGE-L score of augmented settings in each domain firstly increases and then starts decreasing or converging to a certain range.
-The base-data-only setting keeps the rising trend, especially the code domain, which indicates that as the number of training rounds increases, the base-data-only fine-tuning will remember more privacy information, while the risk of privacy leakage of the augmented fine-tuning will gradually decrease or converge.
+We report the average ROUGE-L scores for base-data-only and various public data fine-tuning in Table 5.
+The results show no significant correlation between the public data ratio and privacy-preserving capability in the same training round.
+After 10 rounds of FedDIT, base-data-only fine-tuning outperforms the instruction-augmented settings in all four domains by 6.69%, 5.55%, 5.55%, and 6.16% on average.
+Figure 3 shows the average ROUGE-L score trends per round for each domain, where augmented fine-tuning uses 5000 public data.
+Initially, the ROUGE-L scores for augmented settings increase, then decrease or converge, while the base-data-only scores continue to rise, especially in the code domain.
+This indicates that with more training rounds, base-data-only fine-tuning captures more privacy information, while the privacy leakage risk in augmented fine-tuning decreases or converges.
 6 Conclusion
-We reveal that for FedDIT, the model performance is not linearly correlated with data heterogeneity, but is significantly impacted by cross-client domain coverage.
-To address this, we propose FedDCA, a domain-specific instruction augmentation method that maximizes domain coverage through greedy client center selection and retrieval-based augmentation.
-Additionally, FedDCA uses a heterogeneous encoder to reduce client-side computation overhead.
+We reveal that in the context of FedDIT, the model performance is not linearly correlated with data heterogeneity but is significantly impacted by the cross-client domain coverage.
+In response, we propose a novel FedDIT method called FedDCA, which optimizes the domain coverage through greedy client center selection and retrieval-based instruction augmentation.
+Additionally, FedDCA leverages heterogeneous encoders to reduce the client-side computation overhead and improve system scalability.
 Experiments across four domains demonstrate the effectiveness of FedDCA and the efficiency of FedDCA.
-Further analysis shows that as fine-tuning progresses, the risk of privacy leakage from augmented data decreases or converges.
+Further privacy analysis indicates that as fine-tuning advances, the risk of private data leakage diminishes or converges in FedDIT with instruction augmentation.
 Appendix A Appendix
@@
 Federated Instruction Tuning.
-Instruction tuning has been widely applied across various application areas of large language models (LLM), serving as a key technique to enhance the capabilities and controllability of LLM (Zhang et al.
-(2024b); Wei et al.
-(2022)).
+Instruction tuning has been widely applied across various application areas of large language models (LLM), serving as a key technique to enhance the capabilities and controllability of LLM (Zhang et al., 2024b; Wei et al., 2022).
 Recently, federated instruction tuning (FedIT) has emerged as an effective strategy for the distributed optimization of LLMs, leveraging federated learning (FL) protocols to improve the handling of privacy-sensitive tasks in real-world scenarios.
-So far, several FedIT frameworks(Ye et al.
-(2024b; a)) have been established to evaluate the effectiveness of FedIT across multiple datasets, tasks, and FL methods.
+So far, several FedIT frameworks (Ye et al., 2024b; a) have been established to evaluate the effectiveness of FedIT across multiple datasets, tasks, and FL methods.
 While these platforms provide a foundation for research, they have not yet introduced more complex federated algorithms and deeply investigate the challenging problems and factors affecting FedIT, which are crucial for advancing this field.
-PrivateLoRA(Wang et al.
-(2023b)) addresses privacy and efficiency issues by exploiting the low-rank properties of residual activations to reduce communication costs, significantly lowering the communication overhead through collaborative computation between the server and clients while effectively maintaining the privacy of local data.
-While FewFedPIT(Zhang et al.
-(2024d)) focuses on the few-shot learning setting in FedIT, using self-generated data by pre-trained LLMs locally to mitigate the paucity of data and first discussing memory extraction attacks within FedIT.
+PrivateLoRA (Wang et al., 2023b) addresses privacy and efficiency issues by exploiting the low-rank properties of residual activations to reduce communication costs, significantly lowering the communication overhead through collaborative computation between the server and clients while effectively maintaining the privacy of local data.
+While FewFedPIT (Zhang et al., 2024d) focuses on the few-shot learning setting in FedIT, using self-generated data by pre-trained LLMs locally to mitigate the paucity of data and first discussing memory extraction attacks within FedIT.
 Domain Instruction Augmentation.
-In the real world, there is an urgent need for training LLMs with specific functionalities (e.g., reasoning capabilities) or domain-specific LLMs (e.g., code(Nijkamp et al.
-(2023); Luo et al.
-(2023b)), medical(2024-Zhang-AlpaCare-), financial(Yang et al.
-(2023b; a); Zhang et al.
-(2023a); Wu et al.
-(2023)), mathematical(Yue et al.
-(2024); Luo et al.
-(2023a))).
+In the real world, there is an urgent need for training LLMs with specific functionalities (e.g., reasoning capabilities) or domain-specific LLMs (e.g., code (Nijkamp et al., 2023; Luo et al., 2023b), medical (2024-Zhang-AlpaCare-), financial (Yang et al., 2023b; a; Zhang et al., 2023a; Wu et al., 2023), mathematical (Yue et al., 2024; Luo et al., 2023a)).
 Existing works tend to use open-source domain-specific instruction tuning datasets for training.
@@
 Even if they exist, these datasets are often limited in scale.
-Several studies investigate domain-specific instruction augmentation, which can be categorized into three aspects: 1) Reusing human-curated public datasets(Wang et al.
-(2023a); Zhang et al.
-(2023c); Jiao et al.
-(2023); Lee et al.
-(2024); Xia et al.
-(2024)).
-For instance, Parrot(Jiao et al.
-(2023)) enhances translation capabilities of LLMs in chat by converting bilingual sentences into instruction-following datasets.
-Furthermore, works like INSTA(Lee et al.
-(2024)) and LESS(Xia et al.
-(2024)) attempt efficient domain-specific instruction augmentation via dense retrieval.
-INSTA(Lee et al.
-(2024)) uses instructions without responses for effective retrieval.
-LESS(Xia et al.
-(2024)) assumes access to a validation set and uses the warmup LLM’s gradients of the train and validation sets for retrieval.
-2) Using seed tasks for self-instruct(Luo et al.
-(2023b); Wan et al.
-(2023)): Explore-Instruct(Wan et al.
-(2023)), for example, employs activate exploration to tree-model domain tasks from both depth and breadth, increasing seed instructions’ coverage of the domain, and subsequently generating broader in-domain instructions through self-instruct.
-3) Scaling instructions from the web: Recent works(Yue et al.
-(2024); Zhou et al.
-(2024)) highlight the immense potential of mining naturally occurring instructions from the internet.
+Several studies investigate domain-specific instruction augmentation, which can be categorized into three aspects: 1) Reusing human-curated public datasets (Wang et al., 2023a; Zhang et al., 2023c; Jiao et al., 2023; Lee et al., 2024; Xia et al., 2024).
+For instance, Parrot (Jiao et al., 2023) enhances translation capabilities of LLMs in chat by converting bilingual sentences into instruction-following datasets.
+Furthermore, works like INSTA (Lee et al., 2024) and LESS (Xia et al., 2024) attempt efficient domain-specific instruction augmentation via dense retrieval.
+INSTA (Lee et al., 2024) uses instructions without responses for effective retrieval.
+LESS (Xia et al., 2024) assumes access to a validation set and uses the warmup LLM’s gradients of the train and validation sets for retrieval.
+2) Using seed tasks for self-instruct (Luo et al., 2023b; Wan et al., 2023): Explore-Instruct (Wan et al., 2023), for example, employs activate exploration to tree-model domain tasks from both depth and breadth, increasing seed instructions’ coverage of the domain, and subsequently generating broader in-domain instructions through self-instruct.
+3) Scaling instructions from the web: Recent works (Yue et al., 2024; Zhou et al., 2024) highlight the immense potential of mining naturally occurring instructions from the internet.
 Compared to generated data, web-mined instructions exhibit less bias and greater diversity.
-MAmmoTH2(Yue et al.
-(2024)) firstly retrieves domain-relevant texts and employs a LLM to extract Q-A pairs, further refining them into instruction-response pairs.
-Jiuzhang3.0(Zhou et al.
-(2024)) distills GPT’s instruction generation capabilities into a smaller model and then uses it to generate instructions from the internet.
+MAmmoTH2 (Yue et al., 2024) firstly retrieves domain-relevant texts and employs a LLM to extract Q-A pairs, further refining them into instruction-response pairs.
+Jiuzhang3.0 (Zhou et al., 2024) distills GPT’s instruction generation capabilities into a smaller model and then uses it to generate instructions from the internet.
 In conclusion, models fine-tuned with augmented instructions have shown promising domain-specific capabilities.
-Inspired by INSTA(Lee et al.
-(2024)), this paper uses a mixed dataset composed of code, medical, financial, mathematical, and general datasets to represent public data and does domain-specific instruction augmentation based on the client’s local instructions.
+To obtain a well-performing LLM in the specific domain within the distributed environment, we utilize a multi-domain dataset as public data on the server side and perform domain-specific instruction augmentation based on the client’s local instructions.
 A.2 What Truly Counts in FedDIT
 This section will first analyze the correlation between different non-iid levels and the model’s performance with separate experiments for both single and multiple domains in Section A.2.1.
-Further, to demonstrate the impact of non-iid on FedDIT with different test sets, we compare the performance of the global model trained on augmented data based on iid and non-iid cross-client data distribution.
+Further, to demonstrate the impact of non-iid on FedDIT, we compare the performance of the global model trained on augmented data based on iid and non-iid cross-client data distribution.
 Additionally, we suggest that domain coverage is a key factor for FedDIT in Section A.2.2.
 A.2.1 Data Heterogeneity Is Not Matter In FedDIT
-Following the traditional approach of constructing different degrees of non-iid which are widely used in federated learning(Wang et al.
-(2020); Yurochkin et al.
-(2019)), we adopt Dirichlet distribution to construct various heterogeneity and use k-means with the cluster num $\xi=100$ to pseudo labeling instructions.
+Following the traditional approach of constructing different degrees of non-iid, which are widely used in federated learning (Wang et al., 2020; Yurochkin et al., 2019), we adopt Dirichlet distribution to construct various heterogeneity and use k-means with the cluster num $\xi=100$ to pseudo labeling instructions.
 Dirichlet distribution is affected by the hyperparameter $\alpha$, which enhances heterogeneity with a smaller $\alpha$ and decreases heterogeneity with a larger $\alpha$.
-We choose four widely used heterogeneity, which are $\alpha=[0.01,0.1,1,10]$(Ye et al.
-(2023); Zhang et al.
-(2023b); li2021model).
+We choose four widely used heterogeneity, which are $\alpha=[0.01,0.1,1,10]$ (Ye et al., 2023; Zhang et al., 2023b; li2021model).
 Figure LABEL:fig:data_heterogeneity shows a visualization of the data distribution with different heterogeneity on the code dataset.
 We perform instruction tuning both on single-domain and multi-domain with different heterogeneity.
-Specifically, for the single domain, we only use the in-domain data and then perform FedIT, and for multi-domain, we use the mixed dataset which consists of four domain-specific instruction datasets and a general instruction dataset, which are described in Appendix A.5.
+Specifically, for the single domain, we only use the in-domain data and then perform FedIT.
+For multi-domain, we use the mixed dataset, which consists of four domain-specific instruction datasets and a general instruction dataset, which are described in Appendix A.5.
 The training details are shown in Appendix A.6.
 Results for single domain and multi-domain are shown in Table 6.
-Both single-domain and multi-domain clearly show that the performance of LLM does not decrease due to the increase of data heterogeneity, but shows a nonlinear correlation, which indicates that the performance of LLM does not directly depend on data homogeneity, and exists other factors that play a key role.
+Both single-domain and multi-domain clearly show that the performance of LLM does not decrease due to the increase of data heterogeneity but shows a nonlinear correlation, which indicates that the performance of LLM does not directly depend on data heterogeneity and other factors that play a key role.
 Table 6: Performance(%) of different heterogeneity on single domain and multi-domain.
@@
 A.2.2 Domain Coverage: A Key Factor in FedDIT
-Explore-Instruct(Wan et al.
-(2023)) enhances the coverage of domain-specific seed tasks through active exploration, then uses the self-instruct method for instruction data augmentation.
+Explore-Instruct (Wan et al., 2023) enhances the coverage of domain-specific seed tasks through active exploration, then uses the self-instruct method for instruction data augmentation.
 This approach highlights the impact of domain coverage on domain-specific instruction tuning.
 Inspired by Explore-Instruct, we attempt to conduct more in-depth and extensive experiments to study the effect of domain coverage on FedDIT.
-Firstly we define the domain coverage of cross-client data in the FL setting.
+Firstly, we define the domain coverage of cross-client data in the FL setting.
 Assume the dataset of in-domain data $D^{d}$ represents the latent data distribution of this domain and the cross-client data is defined as $D^{c}=\cup_{k=1}^{N}\left(D_{k}^{l}\cup D_{k}^{g}\right)$.
-Inspired by the facility location function(Cornuejols et al.
-(1983)), we define the domain coverage of $D^{c}$ respect to $D^{d}$ as follows:
+Inspired by the facility location function (Cornuejols et al., 1983), we define the domain coverage of $D^{c}$ respect to $D^{d}$ as follows:
 $d(D^{d},D^{c})=\frac{1}{\left|D^{c}\right|}\sum_{d\in D^{d}}\text{max}_{v\in(D^{c}\cap D^{d})}sim(d,v),$ (5)
@@
 Specifically, we use the cosine similarity function in FedDCA.
-Note that we only use the in-domain data in $D^{c}$ to calculate the domain coverage, because the out-of-domain data would mislead the domain coverage evaluation, as shown in Figure 7.
+Note that we only use the in-domain data in $D^{c}$ to calculate the domain coverage because the out-of-domain data would mislead the domain coverage evaluation, as shown in Figure 7.
 Figure 7: Misleadning of out-of-domain data on domain coverage calculation.
@@
 In this section, we first show the detail of instruction-based dense retrieval in Appendix A.3.1, which is both used in direct retrieval and FedDCA.
-Then we explain the direct retrieval algorithm in Appendix A.3.2.
+Then, we explain the direct retrieval algorithm in Appendix A.3.2.
 A.3.1 Instruction Based Dense Retrieval
@@
 The detailed algorithm is described in Algorithm 2.
-In specific, for each client, we first encode the local instructions by the encoder model $w_{enc}$.
-Then we use the k-means algorithm to cluster the embeddings into $\xi$ clusters.
-Then the $\xi$ cluster centers are sent to the server and do the retrieval-based domain data augmentation.
-Then the retrieved public data is sent to the client for instruction tuning.
+For each client, we start by encoding the local instructions using the encoder model $w_{enc}$.
+Next, we apply the k-means algorithm to cluster the embeddings into $\xi$ clusters.
+Then $\xi$ cluster centers are sent to the server for retrieval-based domain data augmentation.
+Subsequently, the retrieved public data is sent to the client for instruction tuning.
 Algorithm 2 Direct domain data retrieval
@@
 To reduce the computational overhead of clients, we propose a heterogeneous encoder method called FedDCA.
-Inspired by LLaRA(Liao et al.
-(2023)), we utilize a projector to do the dimension alignment between the small encoder on the client side and the large decoder on the server side.
+We utilize a projector to perform the dimension alignment between the small encoder on the client side and the large decoder on the server side.
 As shown in Figure 8, where $d_{l}$ is the dimension of the client-side encoder, $d_{g}$ is the dimension of the server-side encoder.
@@
 A.5 Train and Test Dataset Information
-Here we provide the train and test dataset details of code (CodeAlpaca [^2] https://huggingface.co/datasets/sahil2801/CodeAlpaca-20k), medical (MedAlpaca [^3] https://huggingface.co/datasets/medalpaca), financial (FinGPT(Yang et al.
-(2023a))) and mathematical (MathInstruct [^4] https://huggingface.co/datasets/TIGER-Lab/MathInstruct) domain, respectively.
+Here we provide the train and test dataset details of code (CodeAlpaca [^2] https://huggingface.co/datasets/sahil2801/CodeAlpaca-20k), medical (MedAlpaca [^3] https://huggingface.co/datasets/medalpaca), financial (FinGPT (Yang et al., 2023a)) and mathematical (MathInstruct [^4] https://huggingface.co/datasets/TIGER-Lab/MathInstruct) domain, respectively.
 Table 8: Train set information of each domain.
@@
 For each domain’s FedDIT, we randomly select 1000 samples from the in-domain instruction and split them into 10 shards as each client’s local dataset.
-The rest instructions are used as the public dataset $D^{p}$.
+The rest of the instructions are used as the public dataset $D^{p}$.
 Table 9: Test set information of each domain.
@@
 GSM8K Mathematical 1,319 Exact Match
-Table 9 shows the detail of the test set used for evaluation, HumanEval is used for the code domain, MMLU-Med is used for the medical domain, FPB, FiQA and TFNS are used for the financial domain, and GSM8K is used for the mathematical domain.
+Table 9 shows the detail of the test set used for evaluation: HumanEval is used for the code domain; MMLU-Med is used for the medical domain; FPB, FiQA and TFNS are used for the financial domain, and GSM8K is used for the mathematical domain.
 Specifically, MMLU-Med uses subjects anatomy, clinical_knowledge, college_biology, college_medicine, medical_genetics and professional_medicine in MMLU.
@@
 We consider FedDIT in the cross-device scenario, $N$ = 10 clients, $\mathcal{R}$ = 30 rounds, where we randomly sample 2 clients to be available for each round.
-Then each available client performs FedDIT for 10 steps with AdamW optimizer and the batch size $B=32$ in a round.
+Then, each available client performs FedDIT for 10 steps with AdamW optimizer, and the batch size is $B=32$ in a round.
 The initial learning rate is $55e-$ 5 with a cosine learning rate scheduler.
 Our experiment utilizes the widely used LLM, Llama3-8B [^5] https://huggingface.co/meta-llama/Meta-Llama-3-8B as the base model with 2048 max sequence length and adopts LoRA tuning method.
-The rank of LoRA is 16 and the scalar alpha is 16.
-For k-means(Vardakas & Likas (2023)), we set cluster num $\xi$ = 10 and for FedDCA we set the similarity threshold $\alpha$ = 0.7.
-For FedDCA, we set the temperature parameter $\tau$ = 0.5 for contrastive learning(Khosla et al.
-(2020)).
+The rank of LoRA is 16, and the scalar alpha is 16.
+For k-means (Vardakas & Likas, 2023), we set cluster num $\xi$ = 10 and for FedDCA we set the similarity threshold $\alpha$ = 0.7.
+For FedDCA, we set the temperature parameter $\tau$ = 0.5 for contrastive learning (Khosla et al., 2020).
 We utilize bge-large-en-v1.5666https://huggingface.co/BAAI/bge-large-en-v1.5 as both the client and server’s encoder as default, which outputs embeddings of 1024 dimensions.
@@
 To generate the Self-Instruct data, we prompt GPT-3.5 to generate the instruction with the designed prompt in Figure 9.
-Specifically, we randomly sample two examples from the client’s local data to guide GPT-3.5 generating the in-domain instruction, and sample one example from the client’s local data for one-shot in-context learning to guide GPT-3.5 generating responses into the example’s format.
+Specifically, we randomly sample two examples from the client’s local data to guide GPT-3.5 generating the in-domain instruction and one example from the client’s local data for one-shot in-context learning to guide GPT-3.5 generating responses into the example’s format.
 Figure 9: Prompts used in the Self-Instruct data generation.
@@
 A.9 Further Analysis
-To further study the effect of different hyperparameters of FedDCA, we undertake a thorough analysis including various retrieval amounts, and different k-means cluster numbers $\xi$.
+To further study the effect of different hyperparameters of FedDCA, we undertake a thorough analysis including various retrieval amounts and different k-means cluster numbers $\xi$.
 In addition, we perform the ablation study on whether using the similarity threshold $\alpha$ in the greedy client center selection.
@@
 We show the results on four domains separately.
-Here we use the FPB test set to evaluate the performance in the financial domain.
+Here, we use the FPB test set to evaluate the performance in the financial domain.
 Impact of Different Cluster Number.
@@
 For each domain, we plot each similarity score’s distribution of 10 clients.
-The similarity score is computed between the the selected client center and the public data, then we show the similarity score distribution using the histogram plot.
+The similarity score is computed between the selected client center and the public data.
+Then, we show the similarity score distribution using the histogram plot.
 Table 10: Ablation study on the performance and domain coverage of FedDCA w/o similarity threshold $\alpha$.
@@
 A.10 Augmentation strategy visualization
-To more intuitively compare the domain coverage of different instruction augmentation methods, we randomly sampled 5,000 instructions obtained through these methods and randomly sampled 10,000 in-domain instructions as the background, which represents the distribution of specific domains in the public dataset.
-We then visualized the results using t-SNE(van der Maaten & Hinton (2008)), as shown in Figure LABEL:fig:augmentation_strategy_visualization.
-The plot shows that FedDCA encompasses the most of in-domain data, which is consistent with FedDCA’s domain coverage of each domain shown in Table 3.
+To more intuitively compare the domain coverage of different instruction augmentation methods, we randomly sample 5,000 instructions obtained through these methods and 10,000 in-domain instructions as the background, representing the distribution of specific domains in the public dataset.
+We then visualized the results using t-SNE (van der Maaten & Hinton, 2008), as shown in Figure LABEL:fig:augmentation_strategy_visualization.
+The plot shows that FedDCA encompasses most of the in-domain data, which is consistent with FedDCA’s domain coverage of each domain shown in Table 3.
 Also, we can observe that the random sampling strategy selects a lot of out-of-domain data while does not have good coverage in specific domains.

### v4

@@
 This methodology allows for the exchange of model parameters among distributed data holders, thereby achieving a careful balance between privacy preservation and efficient model optimization.
-Despite the establishment of various FedIT frameworks (Ye et al., 2024b; Kuang et al., 2023; Zhang et al., 2024a), existing literature has not adequately addressed the practical challenges that Federated Domain-specific Instruction Tuning (FedDIT) may encounter in real-world applications.
-For instance, FedIT generally necessitates a sufficient amount of instruction data for fine-tuning, which is often a shortage in domain-specific fine-tuning contexts (Zhang et al., 2024d).
+Despite the establishment of various FedIT frameworks (Ye et al., 2024b; Kuang et al., 2023; Zhang et al., 2023c), existing literature has not adequately addressed the practical challenges that Federated Domain-specific Instruction Tuning (FedDIT) may encounter in real-world applications.
+For instance, FedIT generally necessitates a sufficient amount of instruction data for fine-tuning, which is often a shortage in domain-specific fine-tuning contexts (Zhang et al., 2024b).
 In this study, we investigate FedDIT, a novel approach within the FL paradigm aimed at boosting the performance of LLMs in specific domains.
@@
 Compounding this uncertainty, introducing augmented instructions may further complicate results, making it difficult to ascertain effective improvement strategies.
-Shepherd (Zhang et al., 2024a) approaches this problem from the perspective of heterogeneity, constructing heterogeneity based on the topic of general instruction datasets.
+Shepherd (Zhang et al., 2023c) approaches this problem from the perspective of heterogeneity, constructing heterogeneity based on the topic of general instruction datasets.
 It demonstrates that, unlike the consensus of traditional FL, for general FedIT, heterogeneity has a positive effect.
@@
 These are compared against a range of baselines, which can be categorized into unaugmented and augmented methods.
-In the unaugmented setting, our method is compared with FedIT, which includes four orthodox FL techniques: FedAvg (McMahan et al., 2017), FedProx (Li et al., 2020), SCAFFOLD (pmlr-v119-karimireddy20a), and FedAvgM (Hsu et al., 2019).
-In the augmented setting, we compare FedDCA against methods such as random sampling, direct retrieval, LESS (Xia et al., 2024), and Self-Instruction (Wang et al., 2023c).
+In the unaugmented setting, our method is compared with FedIT, which includes four orthodox FL techniques: FedAvg (McMahan et al., 2017), FedProx (Li et al., 2020), SCAFFOLD (Karimireddy et al., 2020), and FedAvgM (Hsu et al., 2019).
+In the augmented setting, we compare FedDCA against methods such as random sampling, direct retrieval, LESS (Xia et al., 2024), and Self-Instruction (Wang et al., 2022).
 Additionally, we present the performance outcomes of FedDCA when applied under various FL strategies.
@@
 The training process follows the standard FL protocol (McMahan et al., 2017).
-For computational efficiency, we adopt Low-Rank Adaption (LoRA) (Hu et al., 2021) as the fine-tuning method, which involves tuning additional parameters $\Delta\phi$ while keeping the pre-trained LLM’s parameters $\phi$ frozen.
+For computational efficiency, we adopt Low-Rank Adaption (LoRA) (Hu et al., 2022) as the fine-tuning method, which involves tuning additional parameters $\Delta\phi$ while keeping the pre-trained LLM’s parameters $\phi$ frozen.
 In the initial training phase, the server dispatches $\phi$ and $\Delta\phi$ to each client for $\mathcal{R}$ training rounds.
@@
 For better understanding, we list the frequently used notation in Appendix A.11.
-The objective of FedDIT is to enhance the domain-specific performance of LLMs through FL without sharing private data (Ye et al., 2024b; Zhang et al., 2024d).
+The objective of FedDIT is to enhance the domain-specific performance of LLMs through FL without sharing private data (Ye et al., 2024b; Zhang et al., 2024b).
 Under the FL framework, suppose we have $N$ clients, where each client $c_{k}$ has a local dataset $D_{k}^{l}$ with its size $N_{k}^{l}$, and an augmented dataset $D_{k}^{g}$ from the server public dataset $D^{p}$, respectively.
@@
 Suppose there exists a metric space $\mathcal{X}$, a set of cross-client local instruction embeddings $\mathcal{E}\in\mathcal{X}$, and a set of cluster centers $\mathcal{P}\in\mathcal{X}$.
-Each client $c_{k}$ has maximum $\xi$ clusters obtained by k-means algorithm (Vardakas & Likas, 2023).
+Each client $c_{k}$ has maximum $\xi$ clusters obtained by k-means algorithm (Wu, 2012).
 In the federated setting, communication cost is always a critical factor.
@@
 3 in the polynomial time and is quite efficient.
-On the client side, the k-means algorithm (Vardakas & Likas, 2023) is $\mathcal{O}(|D_{k}^{l}|)$.
+On the client side, the k-means algorithm (Wu, 2012) is $\mathcal{O}(|D_{k}^{l}|)$.
 On the server side, as mentioned in Section 4.1, the time complexity of greedy client center selection is $\mathcal{O}\left(N\left(N\xi+\xi\log\xi\right)\right)$.
@@
 The proposed method, FedDCA, does not violate the client’s privacy.
-Compared with other FedIT methods (Zhang et al., 2024d; Ye et al., 2024b), the difference lies in the greedy client center selection.
+Compared with other FedIT methods (Zhang et al., 2024b; Ye et al., 2024b), the difference lies in the greedy client center selection.
 In this stage, the client only uploads the cluster center to the server, which is the average of embeddings to its cluster.
@@
 Baselines.
-We compare FedDCA and FedDCA with the following baselines: 1) Unaugmented methods, including zero-shot inference and FedIT, which are composed of four widely used FL methods, including FedAvg (McMahan et al., 2017), FedProx (Li et al., 2020), SCAFFOLD (pmlr-v119-karimireddy20a) and FedAvgM (Hsu et al., 2019).
-2) Augmented methods, which include random sampling, direct retrieval, LESS (Xia et al., 2024), and Self-Instruct (Wang et al., 2023c).
+We compare FedDCA and FedDCA with the following baselines: 1) Unaugmented methods, including zero-shot inference and FedIT, which are composed of four widely used FL methods, including FedAvg (McMahan et al., 2017), FedProx (Li et al., 2020), SCAFFOLD (Karimireddy et al., 2020) and FedAvgM (Hsu et al., 2019).
+2) Augmented methods, which include random sampling, direct retrieval, LESS (Xia et al., 2024), and Self-Instruct (Wang et al., 2022).
 Additionally, we report FedDCA’s performance with different FL strategies in Table 2.
@@
 However, it does not consider cross-client domain coverage, resulting in an overlapping retrieved data distribution.
-Self-Instruct (Wang et al., 2023c) represents the upper limit performance of the generation-based method, which is restricted to the high expense of cost, limited API call frequency, and heavy quality screening process.
-On the other hand, FewFedPIT (Zhang et al., 2024d) attempts to augment local data by leveraging the pre-trained Llama2 model to perform self-instructed generation during training.
+Self-Instruct (Wang et al., 2022) represents the upper limit performance of the generation-based method, which is restricted to the high expense of cost, limited API call frequency, and heavy quality screening process.
+On the other hand, FewFedPIT (Zhang et al., 2024b) attempts to augment local data by leveraging the pre-trained Llama2 model to perform self-instructed generation during training.
 However, two concerns exist: 1) The pre-trained LLM cannot provide effective, stable, and high-quality instruction generation.
@@
 5.4 Privacy Analysis
-Next, we evaluate the privacy-preserving capability of different ratios of public data against memory extraction attacks (Carlini et al., 2021; Zhang et al., 2024c), which utilizes the autoregression nature of LLM to extract information from the memory (Xu et al., 2024).
+Next, we evaluate the privacy-preserving capability of different ratios of public data against memory extraction attacks (Carlini et al., 2021; Zhang et al., 2024a), which utilizes the autoregression nature of LLM to extract information from the memory (Xu et al., 2024).
 Table 5: Privacy-preserving capability against memory extraction attack with different amounts of augmented public data. We report the average ROUGE-L score of each setting in four domains.
-Retrieval Num Code Med. Fin. Math.
+Retrieval Num. Code Med. Fin. Math.
 Base Data 0.2338 0.1579 0.1527 0.1718
@@
 Federated Instruction Tuning.
-Instruction tuning has been widely applied across various application areas of large language models (LLM), serving as a key technique to enhance the capabilities and controllability of LLM (Zhang et al., 2024b; Wei et al., 2022).
+Instruction tuning has been widely applied across various application areas of large language models (LLM), serving as a key technique to enhance the capabilities and controllability of LLM (Zhang et al., 2023d; Wei et al., 2022).
 Recently, federated instruction tuning (FedIT) has emerged as an effective strategy for the distributed optimization of LLMs, leveraging federated learning (FL) protocols to improve the handling of privacy-sensitive tasks in real-world scenarios.
@@
 PrivateLoRA (Wang et al., 2023b) addresses privacy and efficiency issues by exploiting the low-rank properties of residual activations to reduce communication costs, significantly lowering the communication overhead through collaborative computation between the server and clients while effectively maintaining the privacy of local data.
-While FewFedPIT (Zhang et al., 2024d) focuses on the few-shot learning setting in FedIT, using self-generated data by pre-trained LLMs locally to mitigate the paucity of data and first discussing memory extraction attacks within FedIT.
+While FewFedPIT (Zhang et al., 2024b) focuses on the few-shot learning setting in FedIT, using self-generated data by pre-trained LLMs locally to mitigate the paucity of data and first discussing memory extraction attacks within FedIT.
 Domain Instruction Augmentation.
-In the real world, there is an urgent need for training LLMs with specific functionalities (e.g., reasoning capabilities) or domain-specific LLMs (e.g., code (Nijkamp et al., 2023; Luo et al., 2023b), medical (2024-Zhang-AlpaCare-), financial (Yang et al., 2023b; a; Zhang et al., 2023a; Wu et al., 2023), mathematical (Yue et al., 2024; Luo et al., 2023a)).
+In the real world, there is an urgent need for training LLMs with specific functionalities (e.g., reasoning capabilities) or domain-specific LLMs (e.g., code (Nijkamp et al., 2023; Luo et al., 2024), medical (Zhang et al., 2023e), financial (Yang et al., 2023b; a; Zhang et al., 2023a; Wu et al., 2023), mathematical (Yue et al., 2024; Luo et al., 2023)).
 Existing works tend to use open-source domain-specific instruction tuning datasets for training.
@@
 Even if they exist, these datasets are often limited in scale.
-Several studies investigate domain-specific instruction augmentation, which can be categorized into three aspects: 1) Reusing human-curated public datasets (Wang et al., 2023a; Zhang et al., 2023c; Jiao et al., 2023; Lee et al., 2024; Xia et al., 2024).
+Several studies investigate domain-specific instruction augmentation, which can be categorized into three aspects: 1) Reusing human-curated public datasets (Wang et al., 2023a; Zhang et al., 2023f; Jiao et al., 2023; Lee et al., 2024; Xia et al., 2024).
 For instance, Parrot (Jiao et al., 2023) enhances translation capabilities of LLMs in chat by converting bilingual sentences into instruction-following datasets.
@@
 LESS (Xia et al., 2024) assumes access to a validation set and uses the warmup LLM’s gradients of the train and validation sets for retrieval.
-2) Using seed tasks for self-instruct (Luo et al., 2023b; Wan et al., 2023): Explore-Instruct (Wan et al., 2023), for example, employs activate exploration to tree-model domain tasks from both depth and breadth, increasing seed instructions’ coverage of the domain, and subsequently generating broader in-domain instructions through self-instruct.
+2) Using seed tasks for self-instruct (Luo et al., 2024; Wan et al., 2023): Explore-Instruct (Wan et al., 2023), for example, employs activate exploration to tree-model domain tasks from both depth and breadth, increasing seed instructions’ coverage of the domain, and subsequently generating broader in-domain instructions through self-instruct.
 3) Scaling instructions from the web: Recent works (Yue et al., 2024; Zhou et al., 2024) highlight the immense potential of mining naturally occurring instructions from the internet.
@@
 Dirichlet distribution is affected by the hyperparameter $\alpha$, which enhances heterogeneity with a smaller $\alpha$ and decreases heterogeneity with a larger $\alpha$.
-We choose four widely used heterogeneity, which are $\alpha=[0.01,0.1,1,10]$ (Ye et al., 2023; Zhang et al., 2023b; li2021model).
+We choose four widely used heterogeneity, which are $\alpha=[0.01,0.1,1,10]$ (Ye et al., 2023; Zhang et al., 2023b; Li et al., 2021).
 Figure LABEL:fig:data_heterogeneity shows a visualization of the data distribution with different heterogeneity on the code dataset.
@@
 The rank of LoRA is 16, and the scalar alpha is 16.
-For k-means (Vardakas & Likas, 2023), we set cluster num $\xi$ = 10 and for FedDCA we set the similarity threshold $\alpha$ = 0.7.
+For k-means (Wu, 2012), we set cluster num $\xi$ = 10 and for FedDCA we set the similarity threshold $\alpha$ = 0.7.
 For FedDCA, we set the temperature parameter $\tau$ = 0.5 for contrastive learning (Khosla et al., 2020).

### v5

@@
 Abstract
-Federated Domain-specific Instruction Tuning (FedDIT) utilizes limited cross-client private data together with server-side public data for instruction augmentation, ultimately boosting model performance within specific domains.
+Federated Domain-specific Instruction Tuning (FedDIT) utilizes limited cross-client private data together with various strategies of instruction augmentation, ultimately boosting model performance within specific domains.
 To date, the factors affecting FedDIT remain unclear, and existing instruction augmentation methods primarily focus on the centralized setting without considering distributed environments.
@@
 In response, we propose FedDCA, which optimizes domain coverage through greedy client center selection and retrieval-based augmentation.
+At its core, the greedy selection procedure iteratively picks client centers that maximize the diversity and coverage of the instruction space while avoiding redundancy with previously selected centers.
+This ensures broad yet efficient coverage of the domain distribution across clients.
 For client-side computational efficiency and system scalability, FedDCA, the variant of FedDCA, utilizes heterogeneous encoders with server-side feature alignment.
-Extensive experiments across four distinct domains (code, medical, financial, and mathematical) substantiate the effectiveness of both methods.
-Additionally, we investigate privacy preservation against memory extraction attacks utilizing various amounts of public data.
-Results show that there is no significant correlation between the volume of public data and the privacy-preserving capability.
-However, as the fine-tuning rounds increase, the risk of privacy leakage reduces or converges.
+Extensive experiments across code, medical, financial, and mathematical domains substantiate the effectiveness of both methods, as well as plug-and-play capability.
+We further analyze privacy preservation against memory extraction attacks, showing that while privacy leakage risk is independent of augmented public data ratio, it decreases or converges as training progresses.
 1 Introduction
-Table 1: Performance(%) of different augmentation settings on each domain, conducted via FedAvg protocol with 10 clients.
-Specifically, the zero-shot directly inferences without any fine-tuning, while the base data method utilizes only the client’s local data for FedDIT.
-Additionally, we compare FedDCA with other two augmentation strategies: random sampling and direct retrieval (described in Appendix A.3), respectively.
-Metric Zero-shot Base Data Random Sampling Direct Retrieval FedDCA
-HumanEval/Pass@1 29.88 39.03 32.93 34.14 36.58
-MMLU-Med/Acc 70.60 68.40 71.30 72.20 74.50
-FPB/Acc 55.94 58.25 64.19 66.31 67.24
-FiQA/Acc 18.54 14.18 13.09 19.11 35.27
-TFNS/Acc 59.21 66.62 65.53 67.62 73.32
-GSM8K/Exact Match 23.27 47.46 47.38 50.87 52.46
-Avg. 42.91 48.99 49.07 51.71 56.56
-Recently, federated instruction tuning (FedIT) has gained attention as a novel approach that leverages the principles of federated learning (FL) to facilitate collaborative training of large language models (LLM) in distributed environments while maintaining the confidentiality of private data (McMahan et al., 2017; Ye et al., 2024b).
+Recently, federated instruction tuning (FedIT) has gained attention as a novel approach that leverages the principles of federated learning (FL) to facilitate collaborative training of large language models (LLM) in distributed environments while maintaining the confidentiality of private data McMahan et al.
+(2017); Ye et al.
+(2024b); Zhang et al.
+(2023b).
 This methodology allows for the exchange of model parameters among distributed data holders, thereby achieving a careful balance between privacy preservation and efficient model optimization.
-Despite the establishment of various FedIT frameworks (Ye et al., 2024b; Kuang et al., 2023; Zhang et al., 2023c), existing literature has not adequately addressed the practical challenges that Federated Domain-specific Instruction Tuning (FedDIT) may encounter in real-world applications.
-For instance, FedIT generally necessitates a sufficient amount of instruction data for fine-tuning, which is often a shortage in domain-specific fine-tuning contexts (Zhang et al., 2024b).
+Despite the establishment of various FedIT frameworks Ye et al.
+(2024b); Kuang et al.
+(2023); Zhang et al.
+(2023b), existing literature has not adequately addressed the practical challenges that Federated Domain-specific Instruction Tuning (FedDIT) may encounter in real-world applications.
+For instance, FedIT generally necessitates a sufficient amount of instruction data for fine-tuning, which is often a shortage in domain-specific fine-tuning contexts Zhang et al.
+(2024b).
+Figure 1: Visualization of different instruction augmentation strategies in the mathematical domain.
+The Random Sampling strategy randomly samples instructions from the public dataset.
+Direct Retrieval performs dense retrieval from the public dataset based on the client’s local data distribution without considering the cross-client domain coverage.
+FedDCA strategically selects client centers to maximize cross-client domain coverage and then performs domain data retrieval.
+The t-SNE visualization shows FedDCA has a better domain coverage than other sampling strategies.
 In this study, we investigate FedDIT, a novel approach within the FL paradigm aimed at boosting the performance of LLMs in specific domains.
-Unlike general FedIT, which seeks to enhance model effectiveness across diverse tasks without accounting for local data shortage, FedDIT encounters the unique challenge of clients possessing only a limited quantity of local domain-specific data.
-To overcome this, FedDIT leverages a server-hosted public dataset that spans multiple domains to enrich the local data through a specific instruction augmentation strategy.
-This strategic enrichment is crucial for achieving effective instruction tuning and needs to be meticulously designed to avoid performance degradation.
-Except for the code domain, which primarily adheres to a standardized paradigm, our results reveal that when clients rely solely on their local data for FedDIT, even the presence of high-quality in-domain local data can be insufficient due to limited scale, leading to a decline in performance, as reflected in the underlined values in Table 1.
-In summary, the goal of FedDIT is to develop a domain-specific LLM that employs collaborative training and instruction augmentation with public data while safeguarding client privacy, thereby ensuring the model’s proficiency in executing tasks pertinent to its designated domain.
+Unlike general FedIT, which enhances model effectiveness across diverse tasks without accounting for local data shortages, FedDIT addresses the unique challenge of clients only having limited domain-specific data.
+To overcome this, FedDIT enriches local data through a specific instruction augmentation strategy, crucial for effective instruction tuning and avoiding performance degradation.
+Our results reveal that, except for the code domain, where a standardized paradigm is primarily adhered to, relying solely on high-quality in-domain local data can be insufficient due to limited scale, leading to a decline in performance (see underlined values in Table 2).
+FedDIT aims to develop a domain-specific LLM through collaborative training and instruction augmentation while safeguarding client privacy, ensuring proficiency in tasks pertinent to its designated domain.
+For simplification and better instruction quality Zhang et al.
+(2024b); Toshniwal et al.
+(2024), we focus on a specific scenario of FedDIT involving a server-hosted public dataset, abstracting open-source instruction datasets on the web across multiple domains.
+This dataset is used for various instruction augmentation strategies, enhancing model performance in specific domains.
 Additionally, the factors affecting FedDIT are still unclear.
 Compounding this uncertainty, introducing augmented instructions may further complicate results, making it difficult to ascertain effective improvement strategies.
-Shepherd (Zhang et al., 2023c) approaches this problem from the perspective of heterogeneity, constructing heterogeneity based on the topic of general instruction datasets.
-It demonstrates that, unlike the consensus of traditional FL, for general FedIT, heterogeneity has a positive effect.
-By aggregating diverse instructions from clients, the diversity increases, thereby enhancing the model’s adaptability to various tasks.
-However, it just scratches the surface and does not explore issues in FedDIT.
-Going one step further, we conduct experiments to unveil a significant finding (Appendix A.2): there is no linear correlation between the degree of non-independent and identically distributed (non-iid) and LLM’s performance in the context of FedDIT.
-Inspired by Explore-Instruct (Wan et al., 2023), which shows the potential of domain coverage in domain-specific instruction tuning.
-The cross-client domain coverage metric is initially defined, followed by an investigation into its impact on FedDIT.
-Results demonstrate that domain coverage significantly influences model performance in the corresponding domain.
+We conduct experiments to unveil a significant finding: there is no monotonic correlation between the degree of non-independent and identically distributed (non-iid) and LLM’s performance in the context of FedDIT.
+Inspired by Explore-Instruct Wan et al.
+(2023), which highlights the potential of domain coverage in domain-specific instruction tuning.
+Unlike previous metric that measures domain coverage using the distribution of verb-noun pairs, we define a more general, representation-based cross-client domain coverage metric and investigate its impact on FedDIT.
+Our results show that domain coverage significantly influences model performance within the corresponding domain.
 To maximize the cross-client domain coverage without compromising client data privacy, we propose a novel FedDIT algorithm, Federated Instruction Tuning of LLMs with Domain Coverage Augmentation, termed FedDCA.
 This algorithm employs a greedy client center selection process and implements instruction augmentation through dense retrieval on the server side.
-The fundamental idea of FedDCA is to select client centers to expand the diversity and coverage of augmented instruction datasets within a specific domain.
+The fundamental idea of FedDCA is to select client centers to expand the diversity and coverage of augmented instruction datasets within a specific domain (as shown in Figure 1).
 By strategically optimizing domain coverage at each step, FedDCA efficiently constructs the augmented train set that enhances both the learning and generalization capabilities of the model, leading to superior performance on domain-specific tasks.
@@
 To achieve feature alignment, we train a projector on the server side using public data and employ contrastive learning techniques.
-We demonstrate the effectiveness of FedDCA through comprehensive experiments conducted across four domains: code, medical, financial, and mathematical.
+We demonstrate the effectiveness and plug-and-play capability of FedDCA through comprehensive experiments conducted across four domains: code, medical, financial, and mathematical.
 These are compared against a range of baselines, which can be categorized into unaugmented and augmented methods.
-In the unaugmented setting, our method is compared with FedIT, which includes four orthodox FL techniques: FedAvg (McMahan et al., 2017), FedProx (Li et al., 2020), SCAFFOLD (Karimireddy et al., 2020), and FedAvgM (Hsu et al., 2019).
-In the augmented setting, we compare FedDCA against methods such as random sampling, direct retrieval, LESS (Xia et al., 2024), and Self-Instruction (Wang et al., 2022).
-Additionally, we present the performance outcomes of FedDCA when applied under various FL strategies.
-We also compare the computational efficiency on the client side between FedDCA and its variant, FedDCA.
-For privacy analysis, experiments against memory extraction attacks are conducted to evaluate how different quantities of retrieved public data affect the privacy of client local data.
-Results indicate that while reliance on local data increases memorization of sensitive information, the risk of privacy leakage diminishes or converges in the augmented setting as the training rounds progress.
-The main contribution is as follows:
-  • We reveal a critical finding: in the context of FedDIT, there exists no linear correlation between the degree of heterogeneity and model performance.
-Instead, cross-client domain coverage substantially impacts LLM’s effectiveness, as elaborated in Appendix A.2.
-  • We propose a novel FedDIT algorithm (Section 4), termed FedDCA, aimed at maximizing cross-client domain coverage through greedy client center selection followed by retrieval-based instruction augmentation executed on the server.
-Additionally, we introduce FedDCA to further lessen the client-side computational overhead while enhancing the system scalability.
-This variant utilizes a heterogeneous encoder structure, paired with a projector on the server side for feature alignment.
-  • Through extensive experiments (Section 5), we demonstrate the effectiveness of FedDCA and FedDCA.
-We also investigate privacy preservation against memory extraction attacks, conducting experiments based on various amounts of public data.
-Results suggest that the capacity for privacy preservation does not correlate significantly with the quantity of public data.
-In contrast, the risk of privacy leakage tends to decrease or converge as the fine-tuning rounds increase.
-2 Preliminaries
-FedDIT aims to leverage cross-client private domain-specific instruction data and utilize the multi-domain public data on the server to achieve instruction augmentation, collaboratively enhancing the model’s performance in specific domains.
-Consider $N$ distributed clients, each with local private data $D_{k}^{l}$.
-The server maintains a public dataset $D^{p}$ that encompasses multiple domains and is responsible for implementing data augmentation strategies and aggregating model parameters received from clients.
-The training process follows the standard FL protocol (McMahan et al., 2017).
-For computational efficiency, we adopt Low-Rank Adaption (LoRA) (Hu et al., 2022) as the fine-tuning method, which involves tuning additional parameters $\Delta\phi$ while keeping the pre-trained LLM’s parameters $\phi$ frozen.
-In the initial training phase, the server dispatches $\phi$ and $\Delta\phi$ to each client for $\mathcal{R}$ training rounds.
-In the $t$-th round, the server sends the aggregated $\Delta\phi^{t}$ to clients.
-Clients use $\Delta\phi^{t}$ to update their local LoRA parameters $\Delta\phi_{k}^{t}$ and conduct instruction tuning based on augmented instruction datasets $D_{k}$.
-Subsequently, the clients return $\Delta\phi_{k}^{t+1}$ to the server.
-The server then aggregates $\{\Delta\phi_{k}^{t+1}\mid k=1,\dots,N\}$ to obtain $\Delta\phi^{t+1}$ for the next round.
+Results show that FedDCA outperforms other methods from at least 0.84% to the maximum of 29.19% in terms of model performance, with a relative domain coverage improvement from 4.82% to 21.36%.
+Extensive experiments across privacy analysis, computational efficiency and other further investigations are also conducted.
 2 Problem Formulation
-For better understanding, we list the frequently used notation in Appendix A.11.
-The objective of FedDIT is to enhance the domain-specific performance of LLMs through FL without sharing private data (Ye et al., 2024b; Zhang et al., 2024b).
-Under the FL framework, suppose we have $N$ clients, where each client $c_{k}$ has a local dataset $D_{k}^{l}$ with its size $N_{k}^{l}$, and an augmented dataset $D_{k}^{g}$ from the server public dataset $D^{p}$, respectively.
-Due to constraints such as memory, computational overhead, and maximum tolerated training time, client $c_{k}$ can accept at most $N_{k}^{p}$ public instructions.
-Denote the augmentation strategy as $\Lambda$, through which the server performs instruction augmentation on the public dataset.
-If $\Lambda$ is null, it indicates that clients conduct FedDIT solely based on their local private data, which may lead to performance degradation.
-Conversely, $\Lambda$ may be classified into two categories: (1) focusing exclusively on the client’s own local data distribution or (2) considering the cross-client data distribution.
-In conclusion, the global objective of FedDIT is defined as follows:
-$\operatorname*{arg\,min}_{\Delta\phi}{\left\{F(\phi,\Delta\phi)\triangleq\sum_{k=1}^{N}p_{k}\left((1-\alpha_{k})F_{k}\left(\phi,\Delta\phi_{k};D_{k}^{l}\right)+\alpha_{k}F_{k}\left(\phi,\Delta\phi_{k};D_{k}^{g}\right)\right)\right\}},$ (1)
-where the first term and second term denote the accumulated fine-tuning loss computed on the client $c_{k}$’s local instructions $D_{k}^{l}$ and the augmented public instructions $D_{k}^{g}$ from the server, respectively.
-$p_{k}$ is the weight of the $k$-th client, which is determined by the ratio of $k$-th client’s data size to all clients’ total data size.
-$\alpha_{k}$ is the ratio of the public data amount of its augmented instruction dataset $D_{k}$, which is computed as $\frac{N_{k}^{p}}{N_{k}^{l}+N_{k}^{p}}$.
+Federated Domain-specific Instruction Tuning (FedDIT) is a federated learning approach designed to improve the performance of LLMs in specific domains by utilizing limited cross-client private data in combination with domain-specific instruction augmentation strategies Zhang et al.
+(2024b); Xia et al.
+(2024); Wang et al.
+(2022).
+We abstract the public data as a server-hosted multi-domain dataset (the composition is detailed in Table 1).
+The presence of a public dataset on the server is only one possible scenario, and the proposed algorithm is independent of the presence of a public dataset on the server.
+Even if the server does not have a public dataset, clients can retrieve public instructions by themselves based on the received client centers from the server or perform self-instruct, thereby achieving data augmentation.
+Consider $N$ distributed clients, each with local private data $D_{k}^{l}$ of size $N_{k}^{l}$, and augmented data $D_{k}^{g}$ from the server’s public dataset $D^{p}$.
+Due to constraints like memory and computation, client $c_{k}$ can accept at most $N_{k}^{p}$ public instructions.
+The server maintains $D^{p}$ that spans multiple domains and is responsible for data augmentation strategies $\Lambda$ and parameter aggregation.
+For efficiency, we adopt Low-Rank Adaption (LoRA) Hu et al.
+(2022) as the fine-tuning method, tuning additional parameters $\Delta\phi$ while keeping pre-trained LLM parameters $\phi$ frozen.
+The objective of FedDIT is to enhance the domain-specific performance of LLMs through FL without sharing private data Ye et al. (2024b); Zhang et al. (2024b, 2023b), which is defined as:
+$\small\operatorname*{arg\,min}_{\Delta\phi}\left\{F(\phi,\Delta\phi)\triangleq\sum_{k=1}^{N}p_{k}F_{k}\left(\phi,\Delta\phi_{k};D_{k}^{l},D_{k}^{g};\alpha_{k}\right)\right\},$ (1)
+where $F_{k}(\cdot)$ is the accumulated instruction fine-tuning loss of model $P_{\phi+\Delta\phi_{k}}$ computed on the client $c_{k}$’s augmented dataset $D_{k}^{l}\cup D_{k}^{g}$.
+$p_{k}$ is client $k$’s weight based on data ratio, and $\alpha_{k}=\frac{N_{k}^{p}}{N_{k}^{l}+N_{k}^{p}}$ is the ratio of public data.
 Eq.1 minimizes the summed empirical loss across clients’ augmented instructions to pursue the in-domain utility of the obtained global model.
-Denote $P$ as the model parameters and $\mathcal{D}$ as a specific dataset, then the client $c_{k}$’s empirical loss $F_{k}(\phi,\Delta\phi_{k};\mathcal{D})$ base on $\mathcal{D}$ is calculated as:
+The client $c_{k}$’s empirical loss $F_{k}(\phi,\Delta\phi_{k};\mathcal{D})$ is calculated as:
 $F_{k}(\phi,\Delta\phi_{k};\mathcal{D})\triangleq\frac{1}{|\mathcal{D}|}\sum_{j=1}^{|\mathcal{D}|}l(P_{\phi+\Delta\phi_{k}};x_{j}),$ (2)
 where $x_{j}\in D,\forall j\in\{1,2,\dots,|D|\}$.
-The instruction tuning loss $l(\cdot;\cdot)$ on a sample $(x,y)$ is defined as $-\sum_{t=1}^{\left|y\right|}\log\left(w\left(y_{t}|x,y_{<t}\right)\right)$, where $x$ is the formated instruction with Alpaca instruction template [^1] https://crfm.stanford.edu/2023/03/13/alpaca.html and $y$ is the corressponding response.
-In contrast to the problem formulations in many prior works on FedIT, the primary distinction in the formulation of FedDIT lies in acknowledging the lack of in-domain instructions across clients.
-FedDIT establishes a public multi-domain dataset on the server for domain-specific instruction augmentation by a specific sampling strategy $\Lambda$.
-Therefore, each client only needs to hold a few high-quality domain-specific private data to collaborate with other clients to obtain a strong domain-specific LLM.
+The instruction tuning loss $l(\cdot;\cdot)$ on a sample $(x,y)$ is defined as $-\sum_{t=1}^{\left|y\right|}\log\left(w\left(y_{t}|x,y_{<t}\right)\right)$, where $x$ is the formated instruction and $y$ is the corressponding response.
+In summary, we formalize FedDIT as a novel FL framework that focuses specifically on domain-specific instruction tuning.
+In contrast, FedIT typically demands a substantial quantity of instructions for fine-tuning.
+Domain-specific instruction augmentation can be divided into two categories: 1) Generative methods, which generate instructions through the pre-trained LLM locally or external APIs.
+2) Retrieval-based methods, which retrieve instructions from the web.
+The former methods either compromise privacy or have high computational overhead.
+Thus, in this work, we focus on retrieval-based methods to utilize diverse and high-quality public data.
+3 What Truly Counts in FedDIT
+(a) Exp1. Code
+(b) Exp1. Med
+(c) Exp1. Fin
+(d) Exp1. Math
+(e) Exp2. Performance (%)
+(f) Exp2. Domain Coverage
+Figure 2: (a)-(d) Performance (%) of different heterogeneity in each domain with 10 and 100 clients.
+(e)-(f) Performance and domain coverage of iid and non-iid settings on different domains.
+Experiments show that cross-client domain coverage significantly impacts model effectiveness, while data heterogeneity shows no monotonic correlation with performance.
+Data Heterogeneity Is Not Matter in FedDIT.
+We adopt the Dirichlet distribution Wang et al.
+(2020); Yurochkin et al.
+(2019) to construct various heterogeneity ($\alpha=[0.01,0.1,1,10]$) and use k-means with $\xi=100$ to pseudo labeling instructions.
+We perform instruction tuning on both 10 and 100 clients with 2 randomly selected clients participating in each round.
+For each domain, we only use the in-domain data and then perform FedIT.
+We perform the experiments 3 times with different random seeds (42, 43 and 44) and report the average performance and the standard deviation in each domain.
+As shown in Figures 2(a), 2(b), 2(c) and 2(d), the performance of LLM does not decrease due to the increase of data heterogeneity but shows a non-monotonic correlation, which indicates that the performance of LLM does not directly depend on data heterogeneity and other factors that play a key role.
+Domain Coverage: A Key Factor in FedDIT.
+Different from Explore-Instruct, which defines domain coverage through the distribution of verb-noun pairs, we attempt to conduct more in-depth and extensive experiments to study the effect of domain coverage on FedDIT.
+Firstly, we define the domain coverage in the FL setting, considering the cross-client data distribution.
+Assume the dataset of in-domain data $D^{d}$ represents the latent data distribution of this domain and the cross-client data is defined as $D^{c}=\cup_{k=1}^{N}\left(D_{k}^{l}\cup D_{k}^{g}\right)$.
+Inspired by the facility location function Cornuejols et al. (1983), we define the domain coverage of $D^{c}$ respect to $D^{d}$ as follows:
+$\small d(D^{d},D^{c})=\frac{1}{\left|D^{d}\right|}\sum_{d\in D^{d}}\text{max}_{v\in(D^{c}\cap D^{d})}sim(d,v),$ (3)
+where $sim(d,v)$ is the cosine similarity between $d$ and $v$. Note that we only use the in-domain data in $D^{c}$ for the domain coverage calculation to avoid misleading.
+To better align with the real-world scenarios, we explore instruction augmentation based on both iid and non-iid cross-client data distribution and then perform dense retrieval for FedDIT.
+We set the number of clients to 10, while each client has 100 local instructions and obtains 1000 augmented public instructions from the server.
+For the iid setting, we randomly sample 1k from the public dataset and divide them into 10 shards as each client’s local data.
+For the non-iid distribution, we perform k-means clustering with $\xi=100$.
+Each client randomly samples 100 instructions from randomly selected distinct clusters.
+Figures 2(e) and 2(f) presents the performance and the according domain coverage of FedDIT in different domains with iid and non-iid settings.
+We can observe that both iid and non-iid settings outperform in some domains, but both collectively indicate that higher domain coverage correlates with better performance.
 4 Method
-We propose Federated Instruction Tuning of LLMs with Domain Coverage Augmentation (FedDCA), which enhances domain coverage to obtain a LLM that performs well on domain-specific tasks.
-FedDCA follows the standard FedIT protocol(Ye et al., 2024b) to perform federated instruction tuning.
-The novel design of FedDCA lies in the phase before training, which consists of two key modules: greedy client center selection and domain data retrieval, which are elaborated in the following; also see details in Algorithm 1 and Figure 1.
-For computational efficiency and system scalability, we introduce a variant of FedDCA with heterogeneous encoder setting, named FedDCA.
-We first formulate the optimization problem to solve for FedDCA.
 Figure 3: Overview of FedDCA, which consists of three stages: 1) The client $c_{k}$ performs local instructions clustering and sends cluster centers $\mathcal{C}_{k}$ to the server.
-2) The server first does greedy client center selection to maximize domain coverage and performs client-center-based domain data retrieval, then sends the augmented instructions $D_{k}^{g}$ to the client $c_{k}$.
+2) The server first does greedy client center selection to maximize domain coverage (obtains $\mathcal{P}$) and performs client-center-based domain data retrieval, then sends the augmented instructions $D_{k}^{g}$ to the client $c_{k}$.
 3) Clients fine-tune the LLM collaboratively, and the LoRA parameters $\Delta\phi$ are exchanged between clients and the server.
-Algorithm 1 FedDCA: greedy client center selection & domain data retrieval
-1:Parameters: Number of clusters $\xi$; Encoder model $w_{enc}$; Client local datasets $D=\{D_{1},D_{2},\ldots,D_{N}\}$; Public dataset $D^{p}$; Similarity score threshold $\alpha$; Number of public data samples retrieved per client $N_{k}^{p}$; Client centers $\mathcal{P}=\{p_{1},p_{2},\ldots,p_{N}\}$; Pre-encoded public instruction embeddings $\mathcal{E}$.
-2:for $i\in\{1,2,\ldots,N\}$ do
-3:     $\mathcal{E}^{\prime}\leftarrow\{w_{enc}(x)\mid x\in D_{i,instruction}^{l}\}$
-4:     $\mathcal{C}_{i},S_{i}\leftarrow\text{k-means}(\xi).\text{fit}(\mathcal{E^{\prime}})$triangleright Cluster local instructions, return cluster centers $\mathcal{C}_{i}$ and sizes $S_{i}$
-5:end for
-6:Send $\{\mathcal{C}_{i},S_{i}\mid i\in\{1,2,\ldots,N\}\}$ to the server for the greedy client center selection
-7:$\mathcal{P}\leftarrow\{\mathcal{C}_{0,k}\ \mid k=\arg\max(S_{0})\}$triangleright Initialize the client center set
-8:for $i\in\{1,2,\ldots,N-1\}$ do $\triangleright$ Greedy client center selection
-9:     $\mathcal{S}\leftarrow\sum(\mathcal{C}_{i}\cdot\mathcal{P}^{T},\text{dim}=-1)$triangleright Compute summed similarity score of each cluster center
-10:     $\mathcal{I}\leftarrow(N-i)\text{-}\arg\text{sort}(\mathcal{S})$triangleright Filter $i$ cluster centers close to the client center set $\mathcal{P}$
-11:     $\mathcal{I^{\prime}}\leftarrow\arg\text{sort}(-S_{i})$triangleright Sort cluster center by cluster size in descending order
-12:     $j\leftarrow\text{first element of }\mathcal{I}\cap\mathcal{I}^{\prime}$triangleright Selected cluster center $\mathcal{C}_{i,j}$
-13:     $\mathcal{P}\leftarrow\mathcal{P}\cup\mathcal{C}_{i,j}$triangleright Update the client center set
-14:end for
-15:for $i\in\{1,2,\ldots,N\}$ do $\triangleright$ Client center based domain data retrieval
-16:     $\mathcal{S}\leftarrow\mathcal{P}_{i}\cdot\mathcal{E}^{T}$triangleright Compute similarity score between $\mathcal{P}_{i}$ and $\mathcal{E}$
-17:     $\mathcal{S^{\prime}}\leftarrow\{s\mid s\in\mathcal{S},s<\alpha\}$triangleright Filter instructions with similarity score larger than $\alpha$
-18:     indices of $\mathcal{I}\leftarrow\text{indices of }N_{k}^{p}\text{-top in }\mathcal{S^{\prime}}$
-19:     $D_{i}^{g}\leftarrow\{D^{p}_{j}\mid j\in\mathcal{I}\}$
-20:     $D_{i}\leftarrow D_{i}^{l}\cup D_{i}^{g}$triangleright Obtain the augmented instruction dataset $D_{i}$
-21:end for
+Based on the above empirical observations, we propose FedDCA, which enhances domain coverage to obtain a LLM that performs well on domain-specific tasks (shown in Figure 3).
+We first formulate and analyze the cross-client domain coverage optimization problem.
+Then, we introduce the FedDCA algorithm, which consists of two main stages: greedy client center selection and client-center-based domain data retrieval.
+For computational efficiency and system scalability, we introduce a variant of FedDCA with heterogeneous encoder settings, named FedDCA.
 4.1 Optimization Problem
-As domain coverage directly affects the in-domain performance of the LLM (Appendix A.2), FedDCA aims to maximize the domain coverage of the cross-client augmented data $\cup_{i=1}^{N}D_{i}$ respect to the in-domain data distribution $D^{d}$, which is defined in Eq.
-5.
-However, as clients can not send the local data to the server for privacy, directly finding a cross-client dataset that maximizes the domain coverage in Eq.
-5 is unrealistic.
-To find a proper client center set $\mathcal{P}$ that maximizes domain coverage and uses it to perform instruction retrieval on the public data is an approximation problem.
+As domain coverage directly affects the in-domain performance of the LLM, FedDCA aims to maximize the domain coverage of the cross-client augmented data $D^{c}$ with respect to the in-domain data distribution $D^{d}$.
+However, as clients can not send the local data to the server for privacy, directly finding a cross-client dataset that maximizes the domain coverage is unrealistic.
+Instead, finding a proper client center set $\mathcal{P}$ with optimized domain coverage and using it to perform instruction retrieval on the public data is an approximation problem.
 Suppose there exists a metric space $\mathcal{X}$, a set of cross-client local instruction embeddings $\mathcal{E}\in\mathcal{X}$, and a set of cluster centers $\mathcal{P}\in\mathcal{X}$.
-Each client $c_{k}$ has maximum $\xi$ clusters obtained by k-means algorithm (Wu, 2012).
-In the federated setting, communication cost is always a critical factor.
-Consequently, we formulate the optimization problem to include communication costs as follows:
-$\operatorname*{arg\,min}_{\mathcal{P}}\left\{\sum_{i=1}^{N}|\mathcal{P}_{i}|+\sum_{d\in D^{d}}\left(\text{min}_{p\in\mathcal{P}}sim(d,p)\right)\right\},$ (3)
-[…]
-4.2 Greedy Client Center Selection & Domain Data Retrieval
-As described in Section 4.1, the goal of FedDCA is to find a proper client center set $\mathcal{P}$ that maximizes the domain coverage $d(D^{d},\mathcal{P})$, as defined in Eq.
-3.
-For computational efficiency, we propose a greedy algorithm as shown in Algorithm 1 and Figure 2 to solve this problem in polynomial time and obtain a sub-optimal solution.
-Given the cluster centers $\{\mathcal{C}_{i},i=1,2,\dots,N\}$ received from each client, which are obtained by clustering local instructions.
+Each client $c_{k}$ has maximum $\xi$ clusters obtained by k-means algorithm Wu (2012).
+Additionally, in the federated setting, communication cost is always a critical factor.
+Thus, we formulate the optimization problem as follows:
+$\footnotesize\operatorname*{arg\,min}_{\mathcal{P}}$$\left\{\sum_{i=1}^{N}|\mathcal{P}_{i}|-\frac{1}{|D^{d}|}\sum_{d\in D^{d}}\left(\max_{p\in\mathcal{P}}\operatorname{sim}(d,p)\right)\right\},$ (4)
+s.t. $\quad|\mathcal{P}_{i}|\leq\xi,\quad\forall i\in\{1,2,\dots,N\}$ (5)
+The second term is the domain coverage of the selected client center set $\mathcal{P}$, and $sim(\cdot,\cdot)$ is the cosine similarity function. However, this optimization problem is NP-hard.
+For simplicity, we let $\xi$ be the constant $N$ (further discussed in Section 6.3).
+To solve this problem, the time complexity is $\mathcal{O}\left(C_{\xi N}^{N}\left(\mathcal{N}\left(N+N\log N\right)\right)\right)$, where $\mathcal{N}$ is the size of the public data $D^{p}$.
+As it is a factorial equation, the computational cost increases exponentially with $N$.
+Furthermore, the enormous amount of public data results in a large $\mathcal{N}$.
+To enhance computational efficiency, we propose a greedy algorithm as described below.
+4.2 FedDCA
+As the goal of FedDCA is to find a proper client center set $\mathcal{P}$ to optimize the objective defined in Eq.
+4, we propose a greedy algorithm as shown in Algorithm 1 to solve this problem in polynomial time and obtain a sub-optimal solution.
+Given the cluster centers $\{\mathcal{C}_{i}\}_{i=1}^{N}$ received from each client, which are obtained by clustering local instructions.
 FedDCA on the server consists of two main steps: greedy client center selection and client-center-based domain data retrieval.
-Greedy client center selection.
-We will consider this problem from two aspects: 1) Select a client center that can represent the distribution of the local data.
-2) To optimize the cross-client domain coverage, we filter client centers that are close to the previously selected client centers.
-First, we randomly choose a client and select the largest cluster center as its center and be the initial of the client center set $\mathcal{P}$.
-Then, we iteratively select the client center $\mathcal{P}_{k}$ of the top cluster size while avoiding selecting the cluster center close to the previously selected client center to maximize the domain coverage of $\mathcal{P}$.
-Specifically, for the $i$-th iteration, we filter $i$-top cluster centers that have the largest summed similarity with previously selected client centers and select the center of the rest largest cluster as the $i$-th client center.
-This procedure is repeated until we have selected $N$ client centers.
+Naive Greedy Client Center Selection.
+In this paper’s FedDIT setting, a natural attempt to achieve greedy domain coverage optimization is using the public dataset $D^{p}$ on the server and maximizing $d(D^{d},\mathcal{P})$.
+Here, each client provides a set of candidate embeddings, and the server greedily picks (or updates) the cluster center to maximize domain coverage on $D^{p}$.
+While this approach is intuitive, it has three main limitations: 1) High computation complexity when $D^{p}$ is large.
+Each step of finding the best center $q^{*}$ (to add or update) must re-evaluate coverage w.r.t.
+$|D^{d}|$ data points, often incurring complexity on the order of $\mathcal{O}\bigl{(}N\xi|D^{d}|(N+N\log N)\bigr{)}$ per selection step.
+This quickly becomes prohibitively expensive for large $|D^{d}|$.
+2) Difficulty in domain coverage computation.
+The server may have limited knowledge about which domain each public instruction belongs to, or only know it partially.
+Additionally, the definition of domain is flexible without unified annotation standards - clients’ perceptions of domains may differ from the server’s understanding.
+This makes it challenging for the server to accurately compute domain coverage.
+3) No guaranteed public data.
+In many realistic FedDIT deployments, the server may not actually have an in-domain public dataset.
+Clients might rely on external retrieval or self-instruct generation for data augmentation, leaving no $D^{p}$ accessible on the server.
+The coverage-based greedy approach is then infeasible.
+Cluster-center-based Client Center Selection.
+To address these issues, we propose an alternative method that does not rely on public dataset.
+We denote the full set of these uploaded cluster centers $\{\mathcal{C}_{i}\}_{i=1}^{N}$ as $\mathcal{C}_{\text{all}}$, which is the proxy for domain coverage space.
+Since each client must pick one cluster center from $\mathcal{C}_{\text{all}}$.
+We then define the coverage score for client center selection:
+$\quad C(\mathcal{P})$ $=\sum_{z\,\in\,\mathcal{C}_{\text{all}}}\max_{p\in\mathcal{P}}\mathrm{sim}(z,p),$ (6)
+$where\quad\mathcal{P}$ =$=\{\,p_{1},\dots,p_{N}\,\},\quad p_{i}\in\mathcal{C}_{\text{all}},$ (7)
+and seek to optimize it. This avoids relying on $D^{d}$, greatly reducing coverage-computation overhead and allowing operation without a public dataset.
+Specifically, each client $c_{i}$ locally runs $k$-means to produce $\xi$ candidate centers $\mathcal{C}_{i}$, then uploads these centers to the server.
+The server initializes a selection $\mathcal{P}$ by picking one cluster center from each client (thus $|\mathcal{P}|=N$) and computes the overall coverage solely based on these uploaded centers.
+In each iteration, the algorithm employs a coordinate-ascent style update: it removes the currently chosen center from one client’s portion of $\mathcal{P}$, tries replacing it with other candidate cluster centers, and keeps the replacement if it improves the coverage; this process is repeated until no further improvement is found (as shown in Algorithm 1).
+Algorithm 1 FedDCA: Greedy Client Center Selection
+1:Initialize:
+2: For each client $c_{i}$, pick any initial center $p_{i}\in\mathcal{C}_{i}$.
+3: Let $\mathcal{P}=\{p_{1},\dots,p_{N}\}$ and compute its coverage $C(\mathcal{P})$.
+4:$\Omega^{*}\leftarrow C(\mathcal{P})$
+5:for $i=1,\dots,N$ do
+6:    $\mathcal{P}_{-i}\leftarrow\mathcal{P}\setminus\{p_{i}\}$triangleright Remove $p_{i}$ from the current center set
+7:    $\nu^{*}\leftarrow p_{i}$
+8:    for $q\in(\mathcal{C}_{\text{all}}-\mathcal{P}_{-i})$ do
+9:         $C_{i}=C(\mathcal{P}_{-i}\cup\{q\})$
+10:         if $C_{i}>\Omega^{*}$ then
+11:             $\Omega^{*}\leftarrow C_{i}$; $\nu^{*}\leftarrow q$
+12:         end if
+13:    end for
+14:    $p_{i}\leftarrow\nu^{*}$; $\mathcal{P}\leftarrow\mathcal{P}_{-i}\cup\{p_{i}\}$; $C(\mathcal{P})\leftarrow\Omega^{*}$
+15:end for
+16:Output: The final center set $\mathcal{P}$ and its coverage $C(\mathcal{P})$.
+Since coverage is computed over just $N^{2}$ local centers (i.e., $\mathcal{C}_{\text{all}}$) rather than a large $D^{p}$, the computational overhead is much more manageable.
+Moreover, by not relying on a public dataset, this approach is applicable for various scenarios, including both data augmentation and data selection.
 Domain Data Retrieval.
-For each client center $\mathcal{P}_{k}$, the server performs dense retrieval (detailed in Appendix A.3) on public dataset $D^{p}$ to get the top-$N_{k}^{p}$ similar public instructions, then sends retrieved public datasets $\{D_{1}^{g},\dots,D_{N}^{g}\}$ to each clients.
+For each client center $p_{k}$, the server performs dense retrieval (detailed in Appendix A.2) on public dataset $D^{p}$ to get the top-$N_{k}^{p}$ similar public instructions, then sends retrieved public datasets $\{D_{1}^{g},\dots,D_{N}^{g}\}$ to each clients.
 Specifically, to avoid the overlap between public data and local private data, we set a threshold $\alpha$ to filter the public instructions that have a similarity score larger than $\alpha$ with the client center.
-In summary, FedDCA establishes a comprehensive training set that captures the essence and distribution of the domain by iteratively selecting client centers with unique coverage increments.
-This ensures that selected centers are representative and achieve broad domain coverage.
-Additionally, domain data retrieval exposes the model to a wider range of in-domain features, enhancing its understanding of domain-specific tasks.
-This not only enriches the training data but also reduces the risks of overfitting to the limited local data available to each client, ultimately improving performance across various tasks within the domain.
-Figure 2: Greedy client center selection iteratively selects the client center in each step, which synthetically considers both the representativeness of the cluster center and its distance from the previously selected client center set $\mathcal{P}$ to maximize the cross-client domain coverage.
-4.3 Heterogeneous Encoder with Feature Alignment
-For the client-side computational efficiency and system scalability, we propose a heterogeneous encoder method FedDCA to reduce the client’s computational overhead by using a small encoder $\omega$ on the client side and a larger encoder $\omega^{\prime}$ on the server side.
+FedDCA: Heterogeneous Encoders.
+To enhance client-side computational efficiency and system scalability, we introduce FedDCA.
+This method employs a small encoder $\omega$ on the client side to reduce computational overhead and a larger encoder $\omega^{\prime}$ on the server side.
 However, the output dimension of heterogeneous encoders may not be consistent.
-Therefore, we introduce a projector $w_{p}$ on the server for feature alignment, as shown in Appendix A.4.
-We use contrastive learning (Khosla et al., 2020) and train $w_{p}$ on the public instructions.
-Specifically, $w_{p}$ comprises two fully connected layers and a ReLU activation layer in between, projecting $\omega$’s dimension to the output dimension of $\omega^{\prime}$.
-For $\forall{x_{i}}\in D^{p}$, where $i\in\{1,2,\dots,\mathcal{N}\}$, $\omega$ outputs embedding $h_{i}$, $\omega^{\prime}$ outputs embedding $h_{i}^{\prime}$ and $w_{p}$ outputs embedding $\varphi_{i}$.
+Therefore, we introduce a projector $w_{p}$ on the server for feature alignment, which consists of two fully connected layers and a ReLU activation layer in between.
+We use contrastive learning Khosla et al.
+(2020) and train $w_{p}$ on the public instructions.
+For $\forall{x_{i}}\in D^{p}$, $\omega$ outputs embedding $h_{i}$, $\omega^{\prime}$ outputs embedding $h_{i}^{\prime}$ and $w_{p}$ outputs embedding $\varphi_{i}$.
 For input $x_{i}$, the positive sample pair is $(h_{i}^{\prime},\varphi_{i})$ and negative sample pairs are $\{\varphi_{j},j\neq i\}$.
-Let the batch size be $B$, then the training objective is defined as follows:
-$\mathcal{L}=-\sum_{i=1}^{B}\log\frac{\exp\left(\text{sim}(\varphi_{i},h^{\prime}_{i})/\tau\right)}{\sum_{j=1}^{B}\mathbb{I}_{[j\neq i]}\exp\left(\text{sim}(\varphi_{i},h^{\prime}_{j})/\tau\right)},$ (4)
-where $\text{sim}(\cdot,\cdot)$ denotes the cosine similarity and $\tau$ denotes the temperature parameter.
-The heterogeneous encoder setting, by employing a projector $w_{p}$ for feature alignment, proves effective in FL contexts where clients and the server use different encoder sizes or capacities.
-This approach facilitates mapping lower-dimensional client features $\mathcal{H}$ to the higher-dimensional server space $\mathcal{H}^{\prime}$ through a strategic training of $w_{p}$ with both positive and negative pairs.
-It adeptly aligns similar feature representations closely while positioning dissimilar ones further apart.
-As a result, the heterogeneous encoder setting not only enhances system scalability and flexibility but also facilitates the creation of a unified feature space that effectively integrates and utilizes diverse representations collaboratively.
-4.4 Discussions
-Computation.
+Let the batch size be $B$ and $\tau$ be the temperature parameter, then the training objective is defined as follows:
+$\footnotesize\mathcal{L}=-\sum_{i=1}^{B}\log\frac{\exp\left(\text{sim}(\varphi_{i},h^{\prime}_{i})/\tau\right)}{\sum_{j=1}^{B}\mathbb{I}_{[j\neq i]}\exp\left(\text{sim}(\varphi_{i},h^{\prime}_{j})/\tau\right)},$ (8)
+In summary, FedDCA leverages greedy client center selection and domain data retrieval to maximize cross-client domain coverage, leading to better model performance on domain-specific tasks.
+Additionally, FedDCA employs heterogeneous encoders with server-side feature alignment to reduce client-side computational overhead while maintaining effectiveness.
+5 Discussions
+Computation & Communication.
 Through greedy client center selection, FedDCA solves the optimization problem defined in Eq.
-3 in the polynomial time and is quite efficient.
-On the client side, the k-means algorithm (Wu, 2012) is $\mathcal{O}(|D_{k}^{l}|)$.
-On the server side, as mentioned in Section 4.1, the time complexity of greedy client center selection is $\mathcal{O}\left(N\left(N\xi+\xi\log\xi\right)\right)$.
-Specifically, for each client, the similarity computation between selected client center set $\mathcal{P}$ and client $c_{k}$’s cluster $\mathcal{C}_{k}$ is $\mathcal{O}\left(N\xi\right)$, and the sorting process is $\mathcal{O}\left(\xi\log\xi\right)$.
-For domain data retrieval, given that the sorting algorithm is $\mathcal{O}\left(\mathcal{N}\log\mathcal{N}\right)$, where $\mathcal{N}$ is the size of the public dataset.
-Thus the time complexity is $\mathcal{O}\left(N(\mathcal{N}\log\mathcal{N})\right)$.
+4 in the polynomial time.
+As mentioned in Section 4.1, the time complexity of greedy client center selection is $\mathcal{O}\left(N^{4}\xi^{2}\left(1+\log N\right)\right)$.
+For domain data retrieval, its time complexity is $\mathcal{O}\left(N(\mathcal{N}\log\mathcal{N})\right)$.
 In addition, the heterogeneous encoder setting FedDCA further reduces the computation overhead on the client side.
-Communication.
 In the domain instruction augmentation stage, each client sends $\xi$ cluster centers to the server.
-Next, the server performs greedy client center selection and domain data retrieval, then sends retrieved public data $D_{k}^{g}$ to the client $c_{k}$, whose size is $N_{k}^{p}$.
-In addition, in the fine-tuning stage, FedDCA follows the standard FedIT procedure, which exchanges the LoRA parameters $\Delta\phi_{k}$ between clients and the server.
-Specifically, for the Llama3-8B model, the number of trainable LoRA parameters is just 13.6 million.
-Compared with the number of frozen pre-trained LLM parameters, the communication for LoRA tuning is quite efficient.
+The server returns $N_{k}^{p}$ retrieved public instructions to the client $c_{k}$.
+FedDCA follows standard FedIT procedures, efficiently exchanging LoRA parameters $\Delta\phi_{k}$ between clients and the server.
 Privacy.
-The proposed method, FedDCA, does not violate the client’s privacy.
-Compared with other FedIT methods (Zhang et al., 2024b; Ye et al., 2024b), the difference lies in the greedy client center selection.
+Comparing FedDCA with other FedIT methods Zhang et al.
+(2024b); Ye et al.
+(2024b), the difference lies in the greedy client center selection.
 In this stage, the client only uploads the cluster center to the server, which is the average of embeddings to its cluster.
-In addition, the potential privacy leakage can be further avoided through homomorphic encryption (Acar et al., 2018), which allows the server to directly compute on ciphertext for matrix multiplication for dense retrieval.
+In addition, the potential privacy leakage can be further avoided through homomorphic encryption Acar et al.
+(2018), which allows the server to directly compute on ciphertext for matrix multiplication for dense retrieval.
 6 Experiments
 To demonstrate the effectiveness of FedDCA and its variant FedDCA, we conduct extensive experiments across various domains and with several baselines.
-For additional details and results, please refer to Appendix A.
+For additional results and analysis, please refer to Appendix C.
 6.1 Experimental Setup
+Alpaca CodeAlpaca HumanEval MedAlpaca MMLU-Med
+$N_{sample}$ 52,002 20,022 164 33,955 1,089
+Metric - - Pass@1 - Acc
+FinGPT FPB FiQA TFNS MathInstruct GSM8K
+$N_{sample}$ 76,772 152 35 299 224,567 1,319
+Metric - Acc Acc Acc - Exact Match
+Table 1: Dataset information of each domain.
+The test set is marked by $\ddagger$.
+The public dataset is composed of these five train sets.
+We denote HumanEval and MMLU-Med as H-Eval and M-Med for short.
 Dataset and Evaluation Metrics.
-To evaluate the performance of FedDCA, we conduct experiments on four domains: code, medical, financial, and mathematical.
-The detail of constructing the public dataset is described in Appendix A.5.
-For evaluation, we select a range of datasets, including HumanEval (H-Eval for short) for coding (Chen et al., 2021), MMLU-Med (abbreviated as M-Med) for medical (Hendrycks et al., 2021), and GSM8K for mathematics (Cobbe et al., 2021).
-Additionally, financial datasets such as FPB, FiQA, and TFNS (Yang et al., 2023a) are utilized.
-HumanEval is evaluated using Pass@1.
-For MMLU-Med, FPB, FiQA, and TFNS, we use accuracy as the evaluation metric.
-While GSM8K is evaluated using exact match (EM).
-Please see Appendix A.5 for more dataset details.
+As shown in Table 1, to evaluate the performance of FedDCA, we conduct experiments on four domains: code, medical, financial, and mathematical.
+We set the number of clients to 10 by default, while each client has 100 local instructions and obtains 5k augmented public instructions from the server.
 Baselines.
-We compare FedDCA and FedDCA with the following baselines: 1) Unaugmented methods, including zero-shot inference and FedIT, which are composed of four widely used FL methods, including FedAvg (McMahan et al., 2017), FedProx (Li et al., 2020), SCAFFOLD (Karimireddy et al., 2020) and FedAvgM (Hsu et al., 2019).
-2) Augmented methods, which include random sampling, direct retrieval, LESS (Xia et al., 2024), and Self-Instruct (Wang et al., 2022).
-Additionally, we report FedDCA’s performance with different FL strategies in Table 2.
-Specifically, zero-shot inference shows the performance of the pre-trained LLM without FedDIT, which gives the lower performance bound.
-Direct retrieval is described in Appendix A.3.
-Self-Instruct augments instruction in a generative way through GPT-3.5-turbo (Sun et al., 2023).
-Based on the prompt provided by Self-Instruct, we define the prompts for generating instructions and responses as shown in Appendix A.7.
-5.2 Performance Analysis
+We compare FedDCA and FedDCA with the following baselines: 1) Unaugmented methods, including zero-shot inference and FedIT, which are composed of four widely used FL methods, including FedAvg McMahan et al.
+(2017), FedProx Li et al.
+(2020), SCAFFOLD Karimireddy et al.
+(2020) and FedAvgM Hsu et al.
+(2019).
+2) Augmented methods, including Random Sampling, Direct Retrieval, LESS Xia et al.
+(2024), and Self-Instruct Wang et al.
+(2022).
+Specifically, Random Sampling randomly samples $N_{k}^{p}$ public data for each client, while Direct Retrieval is a simplified version of FedDCA that directly retrieves public data based on each client’s cluster center set without greedy client center selection, not considering the cross-client domain coverage.
+6.2 Main Results
 Performance.
-We evaluate the performance of FedDCA and compare it with several baselines on four domains (see training details in Appendix A.6).
-As shown in Table 2, FedDCA outperforms the other nine baselines in all domains.
-In particular, FedDCA+FedProx achieves the best performance in most domains, with a substantial improvement from at least 5.61% to the maximum of 14.63% over other baselines on the average performance.
-In addition, to reduce the computation overhead of clients, FedDCA attempts to utilize heterogeneous encoders.
-We see that although FedDCA has a performance drop of 1.85% than FedDCA, it still outperforms other baselines in the average performance.
-Furthermore, we report FedDCA’s performance with different FL strategies, and we can see that no FL method can keep the leading position in all domains.
-In specific, FedProx and SCAFFOLD FL strategies perform better in the average performance.
-Overall, the result shows the effectiveness of FedDCA and FedDCA.
-Table 2: Performance(%) of FedDCA, FedDCA and other nine baselines on various domains.
-We report FedDCA with different FL strategies in the last four rows, and FedDCA+FedProx achieves the best average performance across four domains.
-Method H-Eval M-Med FPB FiQA TFNS GSM8K Avg.
-Zero-shot 29.88 70.60 55.94 18.54 59.21 23.27 42.91
-FedAvg 39.03 68.40 58.25 14.18 66.62 47.46 48.99
-FedProx 37.20 69.10 56.51 14.90 66.45 47.15 48.55
-SCAFFOLD 37.80 70.20 62.71 15.27 66.49 49.27 50.29
-FedAvgM 32.32 64.70 68.14 29.27 70.32 46.85 51.93
-Random Sampling 32.93 71.30 64.19 13.09 65.53 47.38 49.07
-Direct Retrieval 34.14 72.20 66.31 19.11 67.62 50.87 51.71
-LESS 28.04 71.00 60.56 16.00 61.14 43.13 46.65
-Self-Instruct 32.92 71.90 59.73 20.67 66.54 50.79 50.43
-FedDCA 34.75 73.30 67.10 30.54 71.01 51.55 54.71
-FedDCA+FedAvg 36.58 74.50 67.24 35.27 73.32 52.46 56.56
-FedDCA+FedProx 32.92 72.40 72.93 38.18 77.55 51.25 57.54
-FedDCA+SCAFFOLD 39.87 73.20 72.68 33.09 75.50 50.26 57.29
-FedDCA+FedAvgM 33.53 68.90 71.45 31.45 72.52 49.76 54.60
-Additionally, Table 2 can be divided into two parts: unaugmented methods and augmented methods.
+Table 2 shows the performance and the corresponding domain coverage of FedDCA and other baselines on four domains.
+FedDCA outperforms the other nine baselines in all domains, with a substantial improvement from at least 0.84% to the maximum of 29.19% over other baselines.
+In particular, FedDCA+FedAvg and FedDCA+FedProx perform better across four domains.
+We see that although FedDCA has a performance drop than FedDCA, it still outperforms other baselines.
 We can observe that FedIT methods perform well in the code domain, even better than most augmented methods.
-This could be attributed to two reasons: 1) The code domain is more about following a certain paradigm.
-2) As the local data is few, so compared with the same epoch and batch size, the unaugmented methods learn the same data more times, which is kind of not a fair setting.
-However, FedDCA+SCAFFOLD still surpasses the best baseline in the code domain, further demonstrating the effectiveness of FedDCA.
-Efficiency.
-We see that direct retrieval is the best baseline in the average performance.
-However, it does not consider cross-client domain coverage, resulting in an overlapping retrieved data distribution.
-Self-Instruct (Wang et al., 2022) represents the upper limit performance of the generation-based method, which is restricted to the high expense of cost, limited API call frequency, and heavy quality screening process.
-On the other hand, FewFedPIT (Zhang et al., 2024b) attempts to augment local data by leveraging the pre-trained Llama2 model to perform self-instructed generation during training.
-However, two concerns exist: 1) The pre-trained LLM cannot provide effective, stable, and high-quality instruction generation.
-2) Generating instructions locally is very costly regarding both time and computational resources.
-Overall, generating instructions self-instructively is not a satisfactory method for the FedDIT scenario.
-LESS (Xia et al., 2024) represents the augmentation methods through gradient feature retrieval.
-However, its performance is underwhelming.
-Compounding this issue, LESS presupposes access to the validation set, which is not always available, and necessitates an initial warmup on the public dataset before calculating gradients for both the public and validation data.
-This process becomes computationally burdensome with large public datasets.
-In conclusion, FedDCA achieves better performance while requiring lower computational resources and a more relaxed training condition, which shows its efficiency.
-Table 3: Domain coverage of FedDCA, FedDCA and other baselines on four domains, where FedIT methods only use local data without domain augmentation.
-Method Code Med. Fin. Math. Avg.
-FedIT 0.8126 0.6990 0.8529 0.7871 0.7879
-Random Sampling 0.8512 0.7940 0.9196 0.8651 0.8575
-Direct Retrieval 0.9396 0.8830 0.9293 0.8967 0.9122
-LESS 0.8509 0.7737 0.8917 0.8352 0.8379
-Self-Instruct 0.8966 0.8586 0.9015 0.8811 0.8845
-FedDCA 0.9532 0.8972 0.9538 0.9096 0.9285
-FedDCA 0.9766 0.9348 0.9815 0.9320 0.9562
+This could be attributed to two reasons: 1) The code domain follows a certain paradigm.
+2) With limited local data, unaugmented methods learn the same data more times given the same epoch and batch size, which may be an unfair comparison.
+Nonetheless, FedDCA still surpasses the best baseline in the code domain.
 Domain Coverage.
-Each method’s domain coverage is shown in Table 3.
-Additionally, visualization of domain coverage across different strategies can be found in Appendix A.10.
-For augmented methods, a correlation is evident between higher domain coverage in a specific domain and improved performance of the method within that domain.
-Given that FedDCA aims to maximize the domain coverage through greedy client center selection, therefore it achieves the highest domain coverage in all domains, which surpasses other baselines from 4.82% to 21.36% average relative improvement.
-Furthermore, while FedDCA employs a smaller encoder on the client side to enhance computational efficiency, this leads to a slight compromise in semantic precision, observed as a relative drop of 2.89% in average domain coverage.
-Nevertheless, FedDCA still outperforms the best baseline by an average of 2.78%.
-In brief, the result proves the effectiveness of FedDCA and FedDCA in domain coverage augmentation.
-5.3 Computation Analysis
-Table 4: Computational cost comparison between FedDCA and FedDCA. We report the encoding time on the public dataset and the model size of each encoder.
-Method Encoding Time Model Size
-FedDCA 15 min 46 s 335 M
-FedDCA 5 min 02 s 22.7 M
-Heterogenous encoder setting FedDCA allows the system to be more scalable and flexible based on different client capacities.
-To show the computational efficiency of FedDCA compared with FedDCA, we compare these two methods from the following aspects: 1) Encoder model size.
+FedDCA achieves the highest domain coverage across all domains, surpassing other baselines by 4.82% to 21.36% on average.
+Although FedDCA uses a smaller encoder on the client side for computational efficiency, it results in a slight 2.89% drop in average domain coverage but still outperforms the best baseline by 2.78% on average.
+These results demonstrate the effectiveness of both FedDCA and FedDCA in enhancing domain coverage.
+Method H-Eval M-Med FPB FiQA TFNS GSM8K
+Zero-shot 29.88/- 70.60/- 55.94/- 18.54/- 59.21/- 23.27/-
+FedAvg McMahan et al. (2017) 39.03/0.8126 68.40/0.6990 58.25/0.8529 14.18/0.8529 66.62/0.8529 47.46/0.7871
+FedProx Li et al. (2020) 37.20/- 69.10/- 56.51/- 14.90/- 66.45/- 47.15/-
+SCAFFOLD Karimireddy et al. (2020) 37.80/- 70.20/- 62.71/- 15.27/- 66.49/- 49.27/-
+FedAvgM Hsu et al. (2019) 32.32/- 64.70/- 68.14/- 29.27/- 70.32/- 46.85/-
+Random Sampling 32.93/0.8512 71.30/0.7940 64.19/0.9196 13.09/0.9196 65.53/0.9196 47.38/0.8651
+Direct Retrieval 34.14/0.9396 72.20/0.8830 66.31/0.9293 19.11/0.9293 67.62/0.9293 50.87/0.8967
+LESS Xia et al. (2024) 28.04/0.8509 71.00/0.7737 60.56/0.8917 16.00/0.8917 61.14/0.8917 43.13/0.8352
+Self-Instruct Wang et al. (2022) 32.92/0.8966 71.90/0.8586 59.73/0.9015 20.67/0.9015 66.54/0.9015 50.79/0.8811
+FedDCA 34.75/0.9532 73.30/0.8972 67.10/0.9538 30.54/0.9538 71.01/0.9538 51.55/0.9096
+FedDCA+FedAvg 36.58/0.9766 74.50/0.9348 67.24/0.9815 35.27/0.9815 73.32/0.9815 52.46/0.9320
+FedDCA+FedProx 32.92/- 72.40/- 72.93/- 38.18/- 77.55/- 51.25/-
+FedDCA+SCAFFOLD 39.87/- 73.20/- 72.68/- 33.09/- 75.50/- 50.26/-
+FedDCA+FedAvgM 33.53/- 68.90/- 71.45/- 31.45/- 72.52/- 49.76/-
+Table 2: Performance (%) and domain coverage of FedDCA, FedDCA and other nine baselines.
+Due to limited local data or inappropriate data augmentation strategy, occurs performance degradation in the Unaugmented and Random Sampling settings (underlined values).
+FedDCA outperforms all other baselines.
+Plug-and-Play.
+To demonstrate the plug-and-play capability of FedDCA with other augmentation methods, we first perform instruction sampling through different baselines to obtain augmented data (5k instructions per client) and then utilize FedDCA to conduct data selection (select 1k instructions per client) for FedDIT.
+As shown in Table 3, using FedDCA in combination with other augmentation methods still achieves comparable performance despite using only 20% of the sampled data.
+FedDCA efficiently selects the most representative samples based on domain coverage, enabling effective training with less data while maintaining performance.
+Computation Analysis.
+To show the computational efficiency of FedDCA, we compare these to FedDCA from two aspects: 1) Encoder model size.
 2) Encoding time overhead.
-We first give the evaluation setup of computation analysis.
-Evaluation setup.
-The encoders used for FedDCA and FedDCA on the client side are bge-large-en-v1.5 and all-MiniLM-L6-v2 respectively (detailed in Appendix A.6).
-Then, we show the model size and the time overhead of encoding the public instructions in Table 4
-Compared to FedDCA, FedDCA has significant computational and time advantages while maintaining acceptable performance (shown in Table 2).
-This is accomplished by employing heterogeneous encoders and configuring a projector on the server to achieve dimensional mapping.
-5.4 Privacy Analysis
-Next, we evaluate the privacy-preserving capability of different ratios of public data against memory extraction attacks (Carlini et al., 2021; Zhang et al., 2024a), which utilizes the autoregression nature of LLM to extract information from the memory (Xu et al., 2024).
-Table 5: Privacy-preserving capability against memory extraction attack with different amounts of augmented public data. We report the average ROUGE-L score of each setting in four domains.
-Retrieval Num. Code Med. Fin. Math.
-Base Data 0.2338 0.1579 0.1527 0.1718
-1000 0.2228 0.1505 0.1451 0.1635
-2000 0.2226 0.1531 0.1445 0.1627
-3000 0.2195 0.1513 0.1454 0.1555
-4000 0.2162 0.1434 0.1438 0.1631
-5000 0.2119 0.1501 0.1445 0.1647
-Evaluation setup.
-We focus on one client’s instruction tuning in FedDIT, using FedDCA for instruction augmentation with 1000 to 5000 public instructions.
+We test the encoders’ encoding time on the public dataset.
+As shown in Figure 6(a), FedDCA requires only about 30% encoded time overhead of FedDCA.
+Moreover, FedDCA achieves a substantial reduction in model size, which is less than 10% of FedDCA’s.
+Privacy Analysis.
+We evaluate the privacy-preserving capability of different ratios of public data against memory extraction attacks Carlini et al.
+(2021); Zhang et al.
+(2024a), which utilizes the autoregression nature of LLMXu et al.
+(2024).
+We analyze the domain inference attack in Section C.3.
+We focus on one client’s instruction tuning in FedDIT, using FedDCA for instruction augmentation with 1k to 5k public instructions.
 We set up 10 clients with full participation for 10 rounds.
-Specifically, we record the average ROUGE-L score (Lin, 2004) of client $c_{0}$ for each round.
-The designed prompt to extract instructions memorized by the LLM is detailed in Appendix A.8.
-For each setting, we repeat memory extraction 100 times and report the average ROUGE-L score based on each generated instruction and all local data instructions to evaluate the privacy-preserving capability of different public data amounts.
-Specifically, denoting the generated $\mathcal{N}$ instructions as $\mathcal{I}$ and the client’s local instructions $\mathcal{I}^{l}$.
-The calculation is defined as $\frac{1}{\mathcal{N}}\sum_{i=1}^{\mathcal{N}}\texttt{ROUGE-L}(\mathcal{I}_{i},\mathcal{I}^{l})$.
-Figure 3: Average ROUGE-L score of each pair of base-data-only and augmented fine-tuning in four domains per round, where different monochromatic represents each domain.
-Results.
-We report the average ROUGE-L scores for base-data-only and various public data fine-tuning in Table 5.
-The results show no significant correlation between the public data ratio and privacy-preserving capability in the same training round.
-After 10 rounds of FedDIT, base-data-only fine-tuning outperforms the instruction-augmented settings in all four domains by 6.69%, 5.55%, 5.55%, and 6.16% on average.
-Figure 3 shows the average ROUGE-L score trends per round for each domain, where augmented fine-tuning uses 5000 public data.
-Initially, the ROUGE-L scores for augmented settings increase, then decrease or converge, while the base-data-only scores continue to rise, especially in the code domain.
+Specifically, we record the average ROUGE-L score Lin (2004) denoted as $\rho$ for client $c_{0}$ in each round.
+The prompt to extract instructions is in Section C.1.
+For each setting, we repeat memory extraction 100 times and report $\rho$.
+Specifically, denote the generated $\mathcal{N}$ instructions as $\mathcal{I}$ and the client’s local instructions $\mathcal{I}^{l}$.
+The calculation is defined as $\rho=\frac{1}{\mathcal{N}}\sum_{i=1}^{\mathcal{N}}\texttt{ROUGE-L}(\mathcal{I}_{i},\mathcal{I}^{l})$.
+(a) Different amounts of augmented public data for FedDIT.
+(b) $\rho$ per round. Dotted lines represent the base-data-only setting.
+Figure 4: Privacy analysis of memory extraction attacks. Data augmentation can effectively mitigate the privacy leakage risk.
+Figure 4(a) shows no significant correlation between the public data ratio and privacy-preserving capability in the same training round.
+In addition, only using local data has a higher risk of privacy leakage than augmented methods.
+Additionally, Figure 4(b) shows $\rho$ trends per round.
+Initially, the $\rho$ for augmented settings increases, then decreases or converges, while the base-data-only scores continue to rise, especially in the code domain.
 This indicates that with more training rounds, base-data-only fine-tuning captures more privacy information, while the privacy leakage risk in augmented fine-tuning decreases or converges.
+Scalability.
+We scale the client number to 100 and 2 clients are randomly selected for FedDIT in each round.
+As the number of clients increases, the amount of local data on each client gradually grows.
+Table 4 shows that the model trained on only base data even outperforms Random Sampling in domains other than code and narrows the gap with FedDCA.
+However, because FedDCA aims to maximize the cross-client domain coverage, it achieves higher domain coverage and better performance.
+Data Selection Scenarios.
+We conduct the experiments on two settings: A) For each domain’s FedDIT, we randomly dispatch the in-domain data equally to each client.
+For example, in the code domain, as the size of the public dataset is 20,022, each client is assigned about 2,000 samples.
+Then we perform FedDIT on each client’s whole data.
+B) Following A’s setting, we first perform FedDCA to determine the client center set, and then we conduct the data selection based on the client center set through dense retrieval to select 200 samples for each client.
+Finally, we perform FedDIT on the selected data.
+Figure 6(b) demonstrates that FedDCA can achieve comparable performance to training with the full dataset by using only 10% data through data selection.
+Held-out Setting.
+If distributed clients aim to solve tasks based on existing knowledge, the public dataset will inevitably contain knowledge relevant to those domains.
+This could come from the original corpus (which can be converted into instruction-response pairs by GPT) or from pre-constructed instruction datasets on the website.
+So the distribution of the public dataset can be categorized as follows: containing held-in or held-out instructions.
+The held-in indicates that the public dataset contains instructions for the specific task that clients aim to solve, while the held-out indicates that the public dataset does not contain this task’s instructions.
+The paper’s default setting is the held-in setting.
+Considering the held-out setting in the financial domain, given that the training set FinGPT and the test sets FPB, FiQA, and TFNS are all related to sentiment analysis tasks.
+We keep the setting of test sets and replace the FinGPT’s instructions in public data with data from a financial QA dataset (Sujet).
+The clients’ local data are still randomly sampled from FinGPT.
+This approach yields held-out public data.
+Figures 6(d), 6(e) and 6(f) shows, FedDCA still achieves performance improvements compared to other baselines for the held-out setting.
+Additionally, using the Random Sampling data augmentation strategy resulted in performance degradation on the FiQA dataset.
+This further underscores the necessity of selecting an appropriate data augmentation strategy.
+6.3 Ablation Study
+Effect of Retrieval Number.
+We report the performance and the corresponding domain coverage of FedDCA with different retrieval amounts on the four domains in Figure 5(b).
+We can see that the domain coverage of FedDCA is increasing along with the retrieval amount in different trends, as well as the performance except for the code domain.
+Noted that although the code domain is more about following a certain paradigm, which could perform well with a few data and more fine-tuning rounds, it still could benefit from the instruction augmentation.
+Similarity Threshold.
+Figure 5(a) shows the performance and the domain coverage of FedDCA w and w/o similarity threshold $\alpha$, where FedDCA w/o $\alpha$ is marked as FedDCA.
+Results show that FedDCA is slightly better than FedDCA in code, financial, and mathematical domains.
+Impact of Different Cluster Number.
+The hyperparameter $\xi$ is the number of clusters in the k-means algorithm.
+The experiment is conducted on $\xi=[N,2N,4N,8N]$, where $N$ is the number of clients.
+We report the domain coverage of the augmented dataset via FedDCA with different $\xi$ on the four domains in Figure 6(c).
+Results show that there is no best $\xi$ for all domains and $\xi=N$ is usually an acceptable choice.
++ FedDCA H-Eval M-Med FPB GSM8K
+Random Sampling 32.51 (-0.42) 71.80 (+0.50) 63.92 (-0.27) 47.82 (+0.44)
+Direct Retrieval 33.89 (-0.25) 72.10 (-0.10) 67.78 (+1.47) 52.53 (+1.66)
+LESS 28.32 (+0.28) 70.80 (-0.20) 61.03 (+0.47) 43.92 (+0.79)
+Self-Instruct 33.15 (+0.23) 70.70 (-1.20) 60.21 (+0.48) 50.25 (-0.54)
+Table 3: Plug-and-Play. Performance (%) of FedDCA with different data augmentation methods. FedDCA can match full dataset performance using just 10% of the data through selection.
+Metric Zero-shot Base Data Random Sampling FedDCA
+H-Eval 29.88/- 34.14/0.8282 34.75/0.8685 35.92/0.9242
+M-Med 70.60/- 72.40/0.8377 69.90/0.8497 73.30/0.9090
+FPB 55.94/- 66.74/0.9339 61.05/0.9408 67.16/0.9800
+GSM8K 23.27/- 49.12/0.8709 47.23/0.8812 50.26/0.9118
+Table 4: Scalability. Performance (%) and domain coverage of FedDCA and other baselines. 100 clients with 2 clients per round. Results show the strong scalability of FedDCA.
+(a) W and w/o $\alpha$
+(b) Different Retrieval Amounts
+Figure 5: Solid bars and lines are performance.
+Performance (%) and domain coverage of FedDCA (a) w and w/o similarity threshold $\alpha$; (b) w different retrieval amounts.
+Experiments show the necessity of similarity threshold for amplifying diversity and the strong applicability of FedDCA with various augmentation scale.
+(a) Computational Cost
+(b) Data Selection Scenarios
+(c) Different $\xi$
+(d) Held-out. FPB
+(e) Held-out. FiQA
+(f) Held-out. TFNS
+Figure 6: (a) Computational cost comparison.
+(b) Domain coverage for different $\xi$.
+(c) Data selection scenarios.
+(d-f) Performance in the held-out setting.
+Results show the efficiency of FedDCA, the effectiveness of FedDCA in data selection scenarios, and the robustness of FedDCA in the held-out setting.
 7 Conclusion
-We reveal that in the context of FedDIT, the model performance is not linearly correlated with data heterogeneity but is significantly impacted by the cross-client domain coverage.
-In response, we propose a novel FedDIT method called FedDCA, which optimizes the domain coverage through greedy client center selection and retrieval-based instruction augmentation.
-Additionally, FedDCA leverages heterogeneous encoders to reduce the client-side computation overhead and improve system scalability.
-Experiments across four domains demonstrate the effectiveness of FedDCA and the efficiency of FedDCA.
-Further privacy analysis indicates that as fine-tuning advances, the risk of private data leakage diminishes or converges in FedDIT with instruction augmentation.
-Appendix A Appendix
-A.1 Related Work
-Federated Instruction Tuning.
-Instruction tuning has been widely applied across various application areas of large language models (LLM), serving as a key technique to enhance the capabilities and controllability of LLM (Zhang et al., 2023d; Wei et al., 2022).
-Recently, federated instruction tuning (FedIT) has emerged as an effective strategy for the distributed optimization of LLMs, leveraging federated learning (FL) protocols to improve the handling of privacy-sensitive tasks in real-world scenarios.
-So far, several FedIT frameworks (Ye et al., 2024b; a) have been established to evaluate the effectiveness of FedIT across multiple datasets, tasks, and FL methods.
-While these platforms provide a foundation for research, they have not yet introduced more complex federated algorithms and deeply investigate the challenging problems and factors affecting FedIT, which are crucial for advancing this field.
-PrivateLoRA (Wang et al., 2023b) addresses privacy and efficiency issues by exploiting the low-rank properties of residual activations to reduce communication costs, significantly lowering the communication overhead through collaborative computation between the server and clients while effectively maintaining the privacy of local data.
-While FewFedPIT (Zhang et al., 2024b) focuses on the few-shot learning setting in FedIT, using self-generated data by pre-trained LLMs locally to mitigate the paucity of data and first discussing memory extraction attacks within FedIT.
-Domain Instruction Augmentation.
-In the real world, there is an urgent need for training LLMs with specific functionalities (e.g., reasoning capabilities) or domain-specific LLMs (e.g., code (Nijkamp et al., 2023; Luo et al., 2024), medical (Zhang et al., 2023e), financial (Yang et al., 2023b; a; Zhang et al., 2023a; Wu et al., 2023), mathematical (Yue et al., 2024; Luo et al., 2023)).
-Existing works tend to use open-source domain-specific instruction tuning datasets for training.
-However, the target domain may not always have corresponding ready-made domain-specific instruction datasets.
-Even if they exist, these datasets are often limited in scale.
-Several studies investigate domain-specific instruction augmentation, which can be categorized into three aspects: 1) Reusing human-curated public datasets (Wang et al., 2023a; Zhang et al., 2023f; Jiao et al., 2023; Lee et al., 2024; Xia et al., 2024).
-For instance, Parrot (Jiao et al., 2023) enhances translation capabilities of LLMs in chat by converting bilingual sentences into instruction-following datasets.
-Furthermore, works like INSTA (Lee et al., 2024) and LESS (Xia et al., 2024) attempt efficient domain-specific instruction augmentation via dense retrieval.
-INSTA (Lee et al., 2024) uses instructions without responses for effective retrieval.
-LESS (Xia et al., 2024) assumes access to a validation set and uses the warmup LLM’s gradients of the train and validation sets for retrieval.
-2) Using seed tasks for self-instruct (Luo et al., 2024; Wan et al., 2023): Explore-Instruct (Wan et al., 2023), for example, employs activate exploration to tree-model domain tasks from both depth and breadth, increasing seed instructions’ coverage of the domain, and subsequently generating broader in-domain instructions through self-instruct.
-3) Scaling instructions from the web: Recent works (Yue et al., 2024; Zhou et al., 2024) highlight the immense potential of mining naturally occurring instructions from the internet.
-Compared to generated data, web-mined instructions exhibit less bias and greater diversity.
-MAmmoTH2 (Yue et al., 2024) firstly retrieves domain-relevant texts and employs a LLM to extract Q-A pairs, further refining them into instruction-response pairs.
-Jiuzhang3.0 (Zhou et al., 2024) distills GPT’s instruction generation capabilities into a smaller model and then uses it to generate instructions from the internet.
-In conclusion, models fine-tuned with augmented instructions have shown promising domain-specific capabilities.
-To obtain a well-performing LLM in the specific domain within the distributed environment, we utilize a multi-domain dataset as public data on the server side and perform domain-specific instruction augmentation based on the client’s local instructions.
-A.2 What Truly Counts in FedDIT
-This section will first analyze the correlation between different non-iid levels and the model’s performance with separate experiments for both single and multiple domains in Section A.2.1.
-Further, to demonstrate the impact of non-iid on FedDIT, we compare the performance of the global model trained on augmented data based on iid and non-iid cross-client data distribution.
-Additionally, we suggest that domain coverage is a key factor for FedDIT in Section A.2.2.
-A.2.1 Data Heterogeneity Is Not Matter In FedDIT
-Following the traditional approach of constructing different degrees of non-iid, which are widely used in federated learning (Wang et al., 2020; Yurochkin et al., 2019), we adopt Dirichlet distribution to construct various heterogeneity and use k-means with the cluster num $\xi=100$ to pseudo labeling instructions.
-Dirichlet distribution is affected by the hyperparameter $\alpha$, which enhances heterogeneity with a smaller $\alpha$ and decreases heterogeneity with a larger $\alpha$.
-We choose four widely used heterogeneity, which are $\alpha=[0.01,0.1,1,10]$ (Ye et al., 2023; Zhang et al., 2023b; Li et al., 2021).
-Figure LABEL:fig:data_heterogeneity shows a visualization of the data distribution with different heterogeneity on the code dataset.
-We perform instruction tuning both on single-domain and multi-domain with different heterogeneity.
-Specifically, for the single domain, we only use the in-domain data and then perform FedIT.
-For multi-domain, we use the mixed dataset, which consists of four domain-specific instruction datasets and a general instruction dataset, which are described in Appendix A.5.
-The training details are shown in Appendix A.6.
-Results for single domain and multi-domain are shown in Table 6.
-Both single-domain and multi-domain clearly show that the performance of LLM does not decrease due to the increase of data heterogeneity but shows a nonlinear correlation, which indicates that the performance of LLM does not directly depend on data heterogeneity and other factors that play a key role.
-Table 6: Performance(%) of different heterogeneity on single domain and multi-domain.
-$\alpha$ Single Domain Multi-Domain
-10 1 0.1 0.01 10 1 0.1 0.01
-H-Eval 32.92 34.14 31.09 34.75 34.14 31.70 34.75 34.75
-M-Med 70.6 69.9 71.0 71.3 71.7 71.3 69.9 71.3
-FPB 67.90 72.11 73.76 68.89 60.06 61.63 60.97 62.04
-FiQA 43.63 45.45 41.81 32.00 10.90 12.72 17.81 9.09
-TFNS 74.37 76.00 75.71 72.69 65.36 65.24 64.69 64.61
-GSM8K 53.75 51.40 51.93 52.76 46.5 43.5 41.5 45.5
-Avg. 57.20 58.17 57.55 55.40 48.11 47.68 48.27 47.88
-A.2.2 Domain Coverage: A Key Factor in FedDIT
-Explore-Instruct (Wan et al., 2023) enhances the coverage of domain-specific seed tasks through active exploration, then uses the self-instruct method for instruction data augmentation.
-This approach highlights the impact of domain coverage on domain-specific instruction tuning.
-Inspired by Explore-Instruct, we attempt to conduct more in-depth and extensive experiments to study the effect of domain coverage on FedDIT.
-Firstly, we define the domain coverage of cross-client data in the FL setting.
-Assume the dataset of in-domain data $D^{d}$ represents the latent data distribution of this domain and the cross-client data is defined as $D^{c}=\cup_{k=1}^{N}\left(D_{k}^{l}\cup D_{k}^{g}\right)$.
-Inspired by the facility location function (Cornuejols et al., 1983), we define the domain coverage of $D^{c}$ respect to $D^{d}$ as follows:
-$d(D^{d},D^{c})=\frac{1}{\left|D^{c}\right|}\sum_{d\in D^{d}}\text{max}_{v\in(D^{c}\cap D^{d})}sim(d,v),$ (5)
-where $sim(d,v)$ is the similarity between $d$ and $v$.
-Specifically, we use the cosine similarity function in FedDCA.
-Note that we only use the in-domain data in $D^{c}$ to calculate the domain coverage because the out-of-domain data would mislead the domain coverage evaluation, as shown in Figure 7.
-Figure 7: Misleadning of out-of-domain data on domain coverage calculation.
-To better align with the real-world scenarios, we explore instruction augmentation based on both iid and non-iid cross-client data distribution and adopt direct data retrieval as described in Appendix A.3 for FedDIT.
-We set the number of clients to 10, while each client has 100 local instructions and obtains 5000 augmented public instructions from the server.
-To construct iid data distribution, we randomly sample 1,000 from multi-domain datasets (code, medical, financial, mathematical, and general), which is detailed in Table 8 and divide them into 10 shards as each client’s local data.
-To construct non-iid data distribution, we perform k-means clustering with $\xi=100$.
-Each client randomly samples 100 instructions from different randomly selected clusters.
-Furthermore, for both iid and non-iid settings, direct data retrieval is performed based on the client’s local data.
-Table 7: Performance(%) and domain coverage of iid and non-iid settings on different domains. The higher domain coverage correlates with better performance.
-Test Set Performance (%) Domain Coverage
-iid non-iid iid non-iid
-H-Eval 35.36 33.53 0.8538 0.7994
-M-Med 70.20 71.00 0.7800 0.8027
-FPB 58.58 64.19 0.8523 0.9327
-FiQA 17.09 19.27
-TFNS 66.16 69.09
-GSM8K 40.50 38.50 0.9137 0.8448
-Table 7 presents the performance of FedDIT on different domains with iid and non-iid settings and shows the domain coverage in four domains.
-We can observe that both iid and non-iid settings outperform in some domains, but both collectively indicate that higher domain coverage correlates with better performance.
-A.3 Direct domain data retrieval
-In this section, we first show the detail of instruction-based dense retrieval in Appendix A.3.1, which is both used in direct retrieval and FedDCA.
-Then, we explain the direct retrieval algorithm in Appendix A.3.2.
-A.3.1 Instruction Based Dense Retrieval
+We reveal that in FedDIT, data heterogeneity has a non-monotonic relationship with model performance, while cross-client domain coverage significantly affects model effectiveness.
+To address this, we propose FedDCA, which optimizes domain coverage through greedy client center selection and retrieval-based instruction augmentation.
+The selection process ensures comprehensive cross-client domain coverage by greedily picking centers that represent maximal distinct regions of the instruction space.
+FedDCA further reduces client-side computation and enhances scalability using heterogeneous encoders.
+Extensive experiments comparing FedDCA with other nine baselines and across four domains demonstrate its effectiveness and plug-and-play capability.
+Appendix
+Appendix A Methodology
+A.1 Complementary Description
+Notation Table.
+For convenience, we provide a detailed notation table in Table 5.
+NOTATIONS REMARK
+$N,C$ Clients number, client set $C=\{c_{1},c_{2},\ldots,c_{N}\}$.
+$D^{p},D^{d},D^{c},D^{l}$ The public datasets, the in-domain data, the cross-client dataset, client’s local private data.
+$N_{k}^{l},N_{k}^{p}$ Number of local private data on the $k$-th client, number of the retrieved public data on the $k$-th client.
+$D_{k}^{l},D_{k}^{g},D_{k}$ The local private data on the $k$-th client, the retrieved public data on the $k$-th client, the augmented dataset on the $k$-th client.
+$\Lambda,\mathcal{P},\mathcal{C}$ A specific sampling strategy that performs instruction augmentation on the server side, selected client center set, the cluster centers obtained locally and sent to the server for the greedy client center selection.
+$\phi,\Delta\phi$ LLM’s pre-trained parameters, additional LoRA parameters.
+$w,w_{enc},w_{p}$ Merged model parameters from the frozen LLM’s parameters $\phi$ and the additional LoRA parameters $\Delta\phi$, encoder model, projector model.
+$F_{k}(w;\mathcal{D}),l(w;x,y)$ Loss of model $w$ over a specific dataset $\mathcal{D}$, the instructon tuning loss of model $w$ over a data sample $\left(x,y\right)$.
+Table 5: Frequently used notation.
+A.2 Direct Domain Data Retrieval
+In this section, we first show the detail of instruction-based dense retrieval, which is both used in direct retrieval and FedDCA. Then, we explain the algorithm of the method direct retrieval.
+A.2.1 Instruction Based Dense Retrieval
 […]
-$\textbf{Score}(\mathcal{E},I)=sim(\mathcal{E},w_{enc}(I)),$ (6)
-where $sim(\cdot,\cdot)$ is the cosine similarity function.
-Based on the computed similarity between $\mathcal{E}$ and each $I$ in the public data, we then select the top-$N_{k}^{p}$ instructions as the retrieved public data for domain data augmentation.
-A.3.2 Direct Retrieval
+A.2.2 Direct Retrieval
 The direct domain data retrieval only utilizes instructions of the client’s local data without responses to perform the retrieval-based domain data augmentation.
@@
 16:end for
-A.4 Projector
-To reduce the computational overhead of clients, we propose a heterogeneous encoder method called FedDCA.
-We utilize a projector to perform the dimension alignment between the small encoder on the client side and the large decoder on the server side.
-As shown in Figure 8, where $d_{l}$ is the dimension of the client-side encoder, $d_{g}$ is the dimension of the server-side encoder.
-Figure 8: The projector used on the server side for feature alignment, which consists of two fully connected layers and a ReLU activation layer in between.
-A.5 Train and Test Dataset Information
-Here we provide the train and test dataset details of code (CodeAlpaca [^2] https://huggingface.co/datasets/sahil2801/CodeAlpaca-20k), medical (MedAlpaca [^3] https://huggingface.co/datasets/medalpaca), financial (FinGPT (Yang et al., 2023a)) and mathematical (MathInstruct [^4] https://huggingface.co/datasets/TIGER-Lab/MathInstruct) domain, respectively.
-Table 8: Train set information of each domain.
-Dataset name Domain $N_{sample}$
-CodeAlpaca Code 20,022
-MedAlpaca Medical 33,955
-FinGPT Financial 76,772
-MathInstruct Mathematical 224,567
-Alpaca General 52,002
-As shown in Table 8, the public data consists of four domain-specific instruction datasets and a general instruction dataset, which are CodeAlpaca, MedAlpaca, FinGPT, MathInstruct, and Alpaca, respectively.
-For each domain’s FedDIT, we randomly select 1000 samples from the in-domain instruction and split them into 10 shards as each client’s local dataset.
-The rest of the instructions are used as the public dataset $D^{p}$.
-Table 9: Test set information of each domain.
-Dataset name Domain $N_{sample}$ Metric
-HumanEval Code 164 Pass@1
-MMLU-Med Medical 1,089 Acc
-FPB Financial 152 Acc
-FiQA Financial 35 Acc
-TFNS Financial 299 Acc
-GSM8K Mathematical 1,319 Exact Match
-Table 9 shows the detail of the test set used for evaluation: HumanEval is used for the code domain; MMLU-Med is used for the medical domain; FPB, FiQA and TFNS are used for the financial domain, and GSM8K is used for the mathematical domain.
-Specifically, MMLU-Med uses subjects anatomy, clinical_knowledge, college_biology, college_medicine, medical_genetics and professional_medicine in MMLU.
-A.6 Implementation Details
+Appendix B Related Works
+Federated Instruction Tuning.
+Instruction tuning has been widely applied across various application areas of large language models (LLM), serving as a key technique to enhance the capabilities and controllability of LLM Zhang et al.
+[2023c]; Wei et al.
+[2022].
+Recently, federated instruction tuning (FedIT) has emerged as an effective strategy for the distributed optimization of LLMs, leveraging federated learning (FL) protocols to improve the handling of privacy-sensitive tasks in real-world scenarios.
+So far, several FedIT frameworks Ye et al.
+[2024b, a]; Zhang et al.
+[2023b] have been established to evaluate the effectiveness of FedIT across multiple datasets, tasks, and FL methods.
+While these platforms provide a foundation for research, they have not yet introduced more complex federated algorithms and deeply investigate the challenging problems and factors affecting FedIT, which are crucial for advancing this field.
+PrivateLoRA Wang et al.
+[2023b] addresses privacy and efficiency issues by exploiting the low-rank properties of residual activations to reduce communication costs, significantly lowering the communication overhead through collaborative computation between the server and clients while effectively maintaining the privacy of local data.
+While FewFedPIT Zhang et al.
+[2024b] focuses on the few-shot learning setting in FedIT, using self-generated data by pre-trained LLMs locally to mitigate the paucity of data and first discussing memory extraction attacks within FedIT.
+Non-IID data distribution is a common challenge in FL Karimireddy et al.
+[2020]; Li et al.
+[2020]; Sattler et al.
+[2019]; Li et al.
+[2019, 2021].
+However, the impact of data heterogeneity on FedIT has not been fully explored.
+Shepherd Zhang et al.
+[2023b] constructing heterogeneity based on the topic of general instruction datasets.
+Results demonstrate that, unlike the consensus of traditional FL, for general FedIT, heterogeneity has a positive effect.
+By aggregating diverse instructions from clients, the diversity increases, thereby enhancing the model’s adaptability to various tasks.
+However, it just scratches the surface and does not explore issues in FedDIT.
+To fully understand the impact of data heterogeneity on FedDIT, we conduct a comprehensive analysis in this work and unveil a crucial finding: there is no monotonic correlation between the degree of non-independent and identically distributed (non-iid) and LLM’s performance in the context of FedDIT (Section 3).
+Domain Instruction Augmentation.
+In the real world, there is an urgent need for training LLMs with specific functionalities (e.g., reasoning capabilities) or domain-specific LLMs (e.g., code Nijkamp et al.
+[2023]; Luo et al.
+[2024], medical Zhang et al.
+[2023d], financial Yang et al.
+[2023b, a]; Zhang et al.
+[2023a]; Wu et al.
+[2023], mathematical Yue et al.
+[2024]; Luo et al.
+[2023]).
+Existing works tend to use open-source domain-specific instruction tuning datasets for training.
+However, the target domain may not always have corresponding ready-made domain-specific instruction datasets.
+Even if they exist, these datasets are often limited in scale.
+Several studies investigate domain-specific instruction augmentation, which can be categorized into three aspects: 1) Reusing human-curated public datasets Wang et al.
+[2023a]; Zhang et al.
+[2023e]; Jiao et al.
+[2023]; Lee et al.
+[2024]; Xia et al.
+[2024].
+For instance, Parrot Jiao et al.
+[2023] enhances translation capabilities of LLMs in chat by converting bilingual sentences into instruction-following datasets.
+Furthermore, works like INSTA Lee et al.
+[2024] and LESS Xia et al.
+[2024] attempt efficient domain-specific instruction augmentation via dense retrieval.
+INSTA Lee et al.
+[2024] uses instructions without responses for effective retrieval.
+LESS Xia et al.
+[2024] assumes access to a validation set and uses the warmup LLM’s gradients of the train and validation sets for retrieval.
+2) Using seed tasks for self-instruct Luo et al.
+[2024]; Wan et al.
+[2023]: Explore-Instruct Wan et al.
+[2023], for example, employs activate exploration to tree-model domain tasks from both depth and breadth, increasing seed instructions’ coverage of the domain, and subsequently generating broader in-domain instructions through self-instruct.
+3) Scaling instructions from the web: Recent works Yue et al.
+[2024]; Zhou et al.
+[2024] highlight the immense potential of mining naturally occurring instructions from the internet.
+Compared to generated data, web-mined instructions exhibit less bias and greater diversity.
+MAmmoTH2 Yue et al.
+[2024] firstly retrieves domain-relevant texts and employs a LLM to extract Q-A pairs, further refining them into instruction-response pairs.
+Jiuzhang3.0 Zhou et al.
+[2024] distills GPT’s instruction generation capabilities into a smaller model and then uses it to generate instructions from the internet.
+In conclusion, models fine-tuned with augmented instructions have shown promising domain-specific capabilities.
+To obtain a well-performing LLM in the specific domain within the distributed environment, we utilize a multi-domain dataset as public data on the server side and perform domain-specific instruction augmentation based on the client’s local instructions.
+Appendix C Experiments
+C.1 Implementation Details
 We consider FedDIT in the cross-device scenario, $N$ = 10 clients, $\mathcal{R}$ = 30 rounds, where we randomly sample 2 clients to be available for each round.
@@
 The rank of LoRA is 16, and the scalar alpha is 16.
-For k-means (Wu, 2012), we set cluster num $\xi$ = 10 and for FedDCA we set the similarity threshold $\alpha$ = 0.7.
-For FedDCA, we set the temperature parameter $\tau$ = 0.5 for contrastive learning (Khosla et al., 2020).
-We utilize bge-large-en-v1.5666https://huggingface.co/BAAI/bge-large-en-v1.5 as both the client and server’s encoder as default, which outputs embeddings of 1024 dimensions.
-While we utilize all-MiniLM-L6-v2777https://huggingface.co/sentence-transformers/all-MiniLM-L6-v2 as the client’s small encoder which outputs embeddings of 384 dimensions and bge-large-en-v1.5 as the server’s encoder for FedDCA.
-A.7 Prompts used in the Self-Instruct data generation
+For k-means Wu [2012], we set cluster num $\xi$ = 10 and for FedDCA we set the similarity threshold $\alpha$ = 0.7.
+For FedDCA, we set the temperature parameter $\tau$ = 0.5 for contrastive learning Khosla et al.
+[2020].
+We utilize bge-large-en-v1.5333https://huggingface.co/BAAI/bge-large-en-v1.5 as both the client and server’s encoder as default, which outputs embeddings of 1024 dimensions.
+While we utilize all-MiniLM-L6-v2444https://huggingface.co/sentence-transformers/all-MiniLM-L6-v2 as the client’s small encoder which outputs embeddings of 384 dimensions and bge-large-en-v1.5 as the server’s encoder for FedDCA.
+Projector.
+To reduce the computational overhead of clients, we propose a heterogeneous encoder method called FedDCA.
+We utilize a projector to perform the dimension alignment between the small encoder on the client side and the large decoder on the server side.
+As shown in Figure 7, where $d_{l}$ is the dimension of the client-side encoder, $d_{g}$ is the dimension of the server-side encoder.
+Figure 7: The projector used on the server side for feature alignment, which consists of two fully connected layers and a ReLU activation layer in between.
+Prompts Used in the Self-Instruct Data Generation.
 To generate the Self-Instruct data, we prompt GPT-3.5 to generate the instruction with the designed prompt in Figure 8.
@@
 We prompt GPT-3.5 to generate responses with a randomly selected example for one-shot in-context learning.
-A.8 Prompt used for memory extraction attack
+Prompt Used for Memory Extraction Attack.
 As we use Llama3-8B as our base model and format the instructions and responses into the Alpaca’s format, to utilize the auto-regression nature of LLM to extract the instruction, we prompt the model to generate the instruction using the prompt in Figure 9, which is exactly the prefix of the Alpaca’s template.
 Figure 9: Prompt used for memory extraction attack.
-A.9 Further Analysis
-To further study the effect of different hyperparameters of FedDCA, we undertake a thorough analysis including various retrieval amounts and different k-means cluster numbers $\xi$.
-In addition, we perform the ablation study on whether using the similarity threshold $\alpha$ in the greedy client center selection.
-Effect of Retrieval Number.
-We report the performance and the corresponding domain coverage of FedDCA with different retrieval amounts on the four domains in Figure 11, respectively.
-We can see that the domain coverage of FedDCA is increasing along with the retrieval amount in different trends, as well as the performance.
-Specifically, the domain coverage of each domain increases by 6.36%, 18.69%, 5.04%, and 8.06% in relative, respectively.
-Along with domain coverage increasing, the performance of FedDCA is increasing by 3.05%, 2.20%, 4.08%, and 6.95% for each domain.
-Noted that although the code domain is more about following a certain paradigm, which could perform well with a few data and more fine-tuning rounds, it still could benefit from the instruction augmentation.
-Figure 11: Effect of different retrieval amounts on the performance of FedDCA and its domain coverage.
-We show the results on four domains separately.
-Here, we use the FPB test set to evaluate the performance in the financial domain.
-Impact of Different Cluster Number.
-The hyperparameter $\xi$ is the number of clusters in the k-means algorithm.
-The experiment is conducted on $\xi=[N,2N,4N,8N]$, where $N$ is the number of clients.
-Following the setting in Appendix A.6, we set $N=10$.
-We report the domain coverage of the augmented dataset via FedDCA with different cluster numbers $\xi$ on the four domains in Figure 12, respectively.
-Results show that there is no best $\xi$ for all domains.
-Specifically, the best $\xi$ of code, medical, financial, and mathematical domains are 80, 80, 40, and 10, respectively.
-Figure 12: Impact of different cluster number $\xi$ on the cross-client domain coverage.
-We report the domain coverage of the augmented dataset via FedDCA with different cluster numbers $\xi$ on the four domains.
-Ablation Study.
-We conduct the experiment with FedDCA w/o similarity threshold $\alpha$ on the four domains based on the FedAvg FL strategy.
-The performance and the corresponding domain coverage are shown in Table 10, where FedDCA without the similarity threshold $\alpha$ is marked as FedDCA.
-The result shows that the performance of FedDCA with similarity threshold $\alpha$ is slightly better than FedDCA without using $\alpha$ in code, financial, and mathematical domains, as the similarity scores in the medical domain are relatively lower.
-We show the similarity score distribution of the four domains in Figure LABEL:fig:similarity_score.
-For each domain, we plot each similarity score’s distribution of 10 clients.
-The similarity score is computed between the selected client center and the public data.
-Then, we show the similarity score distribution using the histogram plot.
-Table 10: Ablation study on the performance and domain coverage of FedDCA w/o similarity threshold $\alpha$.
-Metric Performance(%) Domain Coverage
-FedDCA FedDCA FedDCA FedDCA
-H-Eval 35.97 36.58 0.8972 0.9348
-M-Med 73.40 73.40 0.9348 0.9348
-FPB 66.25 67.24 0.9353 0.9815
-FiQA 23.27 35.27
-TFNS 69.34 73.32
-GSM8K 51.78 52.46 0.9128 0.9320
-A.10 Augmentation strategy visualization
+C.2 Baselines
+Method Privacy Preserving API Cost Additional Information Performance Degradation Domain Coverage Oriented
+FedAvg McMahan et al. [2017]
+Random Sampling
+LESS Xia et al. [2024]
+Self-Instruct Wang et al. [2022]
+Direct Retrieval
+FedDCA (ours)
+Table 6: Key differences between FedDCA and other baselines.
+FedDCA demonstrates the ability of: 1) privacy-preserving, 2) no API cost, 3) no additional information required, 4) avoiding performance degradation, and 5) aiming at domain coverage optimization.
+LESS requires additional information, as it needs access to the validation set for gradient-based retrieval.
+In Table 6, we highlight five critical aspects of our approach compared with other methods Xia et al. [2024]; Wang et al. [2022]; Zhang et al. [2024b]. Our comparison focuses on five key dimensions:
+  • Privacy Preserving: Most methods except Self-Instruct maintain data privacy, we analyze the privacy-preserving ability of FedDCA in Sections 5 and C.3.
+  • API Cost: FedDCA requires no api calls. As Self-Instruct utilizes the GPT to generate instructions, it requires additional API calls, which incurs costs.
+  • Additional Information: LESS needs validation set access for gradient-based retrieval, whereas FedDCA works with only the basic client’s local training data.
+  • Performance Degradation: From Table 2 we see that there occurs performance degradation for FedAvg, Random Sampling, and Self-Instruct, while FedDCA avoids this issue.
+  • Domain Coverage: Uniquely, FedDCA is specifically designed to optimize domain coverage for distributed environments, a feature absent in other baseline methods.
+Specifically, Self-Instruct Wang et al.
+[2022] represents the upper limit performance of the generation-based method, which is restricted to the high expense of cost, limited API call frequency, and heavy quality screening process.
+On the other hand, FewFedPIT Zhang et al.
+[2024b] attempts to augment local data by leveraging the pre-trained Llama2 model to perform self-instructed generation during training.
+However, two concerns exist: 1) The pre-trained LLM cannot provide effective, stable, and high-quality instruction generation.
+2) Generating instructions locally is very costly regarding both time and computational resources.
+Overall, generating instructions self-instructively is not a satisfactory method for the FedDIT scenario.
+LESS Xia et al.
+[2024] represents the augmentation methods through gradient feature retrieval.
+However, its performance is underwhelming.
+Compounding this issue, LESS presupposes access to the validation set, which is not always available, and necessitates an initial warmup on the public dataset before calculating gradients for both the public and validation data.
+This process becomes computationally burdensome with large public datasets.
+In conclusion, these comparisons demonstrate FedDCA’s distinctive approach to balancing privacy preservation, computational overhead and domain coverage optimization, while maintaining minimal additional requirements.
+C.3 Domain Inference Attacks
+The server may inference the clients’ data domain when the domain data retrieval is performed on the server side.
+However, the proposed algorithm FedDCA is independent of the presence of a public dataset on the server.
+Even if the server does not have a public dataset, clients can upload their cluster centers to the server, which selects a set of client centers and sends them back to the clients.
+Each client can then retrieve data from the website based on the received client center by itself, thereby achieving data augmentation while maximizing the cross-client domain coverage.
+In that case, since the server does not know the encoder used by the client, it cannot infer the semantic meaning of the embedding.
+Thus, for the server, it becomes significantly more challenging to infer the client’s domain, let alone apply any privacy protection techniques to the embeddings.
+We provide two examples for illustration.
+Two different encoders are used as client’s and server’s respectively: BAAI/bge-large-en-v1.5 (denoted as $w_{1}$) and google-bert/bert-large-uncased (denoted as $w_{2}$).
+Both encoders output 1024-dimensional features.
+Example 1: Both $w_{1}$ and $w_{2}$ take "hello world" as input, and the cosine similarity between their embeddings is 0.1829.
+Example 2: Three instructions are used:
+  • Instruction 1: Create an array of length 5 which contains all even numbers between 1 and 10.
+  • Instruction 2: Write a replace method for a string class which replaces the given string with a given set of characters.
+  • Instruction 3: What is the sentiment of this news?
+Please choose an answer from {negative/neutral/positive}.
+Teollisuuden Voima Oyj, the Finnish utility known as TVO, said it shortlisted Mitsubishi Heavy’s EU-APWR model along with reactors from Areva, Toshiba Corp., GE Hitachi Nuclear Energy, and Korea Hydro & Nuclear Power Co.
+For these instructions, Instruction 1 is passed to $w_{1}$ and Instructions 2 and 3 are passed to $w_{2}$, which will result in three embeddings: $e_{1}$, $e_{2}$, and $e_{3}$.
+The cosine similarity between $e_{1}$ and $e_{2}$ is 0.1464, while the similarity between $e_{1}$ and $e_{3}$ is 0.1879.
+Instructions 1 and 2 are in the same domain, whereas they have a lower cosine similarity.
+In conclusion, as demonstrated above, when the clients do not perform domain-specific instruction retrieval on the server side, the server cannot infer the client’s domain based on the uploaded embeddings.
+C.4 How Far can FedDCA Go
+Assume an extreme setting, where exists a pre-trained model that has already fine-tuned on the whole in-domain data of the internet or even all open-source public instruction datasets (it is actually unrealistic or temporary, as new data is constantly being generated), is there any room for instruction augmentation or FedDCA?
+To answer this question, we conduct the following experiments on four settings: A) Fine-tuning on the whole public dataset.
+B) Following A, perform further federated fine-tuning with only private clients’ local dataset.
+C) Following A, each client randomly samples 100 public data for fine-tuning.
+D) Following A, each client samples 100 public data through FedDCA and then performs FedDIT.
+Setting H-Eval M-Med FPB FiQA TFNS GSM8K
+A 38.41 69.10 75.33 41.09 71.60 53.75
+B 43.90 64.30 77.97 66.54 78.26 56.40
+C 41.46 66.90 74.50 29.81 70.51 57.01
+D 44.12 70.90 81.43 72.00 78.81 58.90
+Table 7: Performance (%) of setting A, B, C, and D in each domain.
+As shown in Table 7, even with only 100 samples selected via FedDCA, the method not only prevents performance degradation but also helps the model achieve better generalization within the domain.
+This is attributed to exposure to more diverse in-domain data during training.
+Meanwhile, it can be observed that random sampling still leads to performance degradation compared to Setting A, and even performs worse than Setting B, where no data augmentation is applied.
+In the financial domain, performance degradation is also evident.
+In conclusion, this experiment further highlights the necessity and effectiveness of FedDCA.
+C.5 Augmentation Strategy Visualization
 To more intuitively compare the domain coverage of different instruction augmentation methods, we randomly sample 5,000 instructions obtained through these methods and 10,000 in-domain instructions as the background, representing the distribution of specific domains in the public dataset.
-We then visualized the results using t-SNE (van der Maaten & Hinton, 2008), as shown in Figure LABEL:fig:augmentation_strategy_visualization.
+We then visualized the results using t-SNE van der Maaten and Hinton [2008], as shown in Figure 10.
 The plot shows that FedDCA encompasses most of the in-domain data, which is consistent with FedDCA’s domain coverage of each domain shown in Table 2.
 Also, we can observe that the random sampling strategy selects a lot of out-of-domain data while does not have good coverage in specific domains.
-A.11 Frequently Used Notation
-Table 11: Frequently used notation.
-NOTATIONS REMARK
-$N,C$ Clients number, client set $C=\{c_{1},c_{2},\ldots,c_{N}\}$.
-$D^{p},D^{d},D^{c},D^{l}$ The public datasets, the in-domain data, the cross-client dataset, client’s local private data.
-$N_{k}^{l},N_{k}^{p}$ Number of local private data on the $k$-th client, number of the retrieved public data on the $k$-th client.
-$D_{k}^{l},D_{k}^{g},D_{k}$ The local private data on the $k$-th client, the retrieved public data on the $k$-th client, the augmented dataset on the $k$-th client.
-$\Lambda,\mathcal{P},\mathcal{C}$ A specific sampling strategy that performs instruction augmentation on the server side, selected client center set, the cluster centers obtained locally and sent to the server for the greedy client center selection.
-$\phi,\Delta\phi$ LLM’s pre-trained parameters, additional LoRA parameters.
-$w,w_{enc},w_{p}$ Merged model parameters from the frozen LLM’s parameters $\phi$ and the additional LoRA parameters $\Delta\phi$, encoder model, projector model.
-$F_{k}(w;\mathcal{D}),l(w;x,y)$ Loss of model $w$ over a specific dataset $\mathcal{D}$, the instructon tuning loss of model $w$ over a data sample $\left(x,y\right)$.
+(a) Code
+(b) Medical
+(c) Financial
+(d) Mathematical
+Figure 10: Visualization of cross-client data distribution in different domains, performing t-SNE dimensionality reduction on retrieved instructions through various augmentation strategies.
+We randomly sample 10,000 in-domain samples as background while randomly sampling 5,000 samples from the cross-client augmented dataset for different instruction augmentation methods for comparison.

### v6

@@
-Federated Instruction Tuning of LLMs with Domain Coverage Augmentation
+Optimizing Cross-Client Domain Coverage for Federated Instruction Tuning of Large Language Models
 Abstract
-Federated Domain-specific Instruction Tuning (FedDIT) utilizes limited cross-client private data together with various strategies of instruction augmentation, ultimately boosting model performance within specific domains.
-To date, the factors affecting FedDIT remain unclear, and existing instruction augmentation methods primarily focus on the centralized setting without considering distributed environments.
-Our experiments reveal that the cross-client domain coverage, rather than data heterogeneity, drives model performance in FedDIT.
-In response, we propose FedDCA, which optimizes domain coverage through greedy client center selection and retrieval-based augmentation.
-At its core, the greedy selection procedure iteratively picks client centers that maximize the diversity and coverage of the instruction space while avoiding redundancy with previously selected centers.
-This ensures broad yet efficient coverage of the domain distribution across clients.
-For client-side computational efficiency and system scalability, FedDCA, the variant of FedDCA, utilizes heterogeneous encoders with server-side feature alignment.
-Extensive experiments across code, medical, financial, and mathematical domains substantiate the effectiveness of both methods, as well as plug-and-play capability.
-We further analyze privacy preservation against memory extraction attacks, showing that while privacy leakage risk is independent of augmented public data ratio, it decreases or converges as training progresses.
+Federated domain-specific instruction tuning (FedDIT) for large language models (LLMs) aims to enhance performance in specialized domains using distributed private and limited data, yet identifying key performance drivers and optimal augmentation strategies remains challenging.
+We empirically establish that cross-client domain coverage, rather than data heterogeneity, is the pivotal factor.
+We then introduce FedDCA, an algorithm that explicitly maximizes this coverage through diversity-oriented client center selection and retrieval-based augmentation, constructing diverse, non-redundant cross-client instruction sets.
+Extensive experiments across multiple domains demonstrate FedDCA’s superiority over eleven baselines, achieving performance gains of up to 29.19% and domain coverage improvements of 4.82%-21.36%.
+FedDCA maintains its effectiveness in diverse and challenging scenarios, including data selection, held-out settings where task-specific public data is scarce and various data heterogeneity, with manageable privacy risks.
+This work clarifies critical FedDIT dynamics and presents FedDCA as an effective, privacy-preserving, and scalable solution for advancing domain-specific LLM tuning.
+Optimizing Cross-Client Domain Coverage for Federated Instruction Tuning of Large Language Models
+Zezhou Wang1, Yaxin Du2, Xingjun Ma3, Yugang Jiang3, Zhuzhong Qian1 *, Siheng Chen2 *, 1Nanjing University, 2Shanghai Jiao Tong University, 3Fudan University {zzw.cs@smail, qzz@}nju.edu.cn, {dorothydu, sihengc}@sjtu.edu.cn, {xingjunma, ygj}@fudan.edu.cn
+\$*\$\$*\$footnotetext: Corresponding author.
 1 Introduction
@@
 (2017); Ye et al.
-(2024b); Zhang et al.
+(2024c); Zhang et al.
 (2023b).
@@
 Despite the establishment of various FedIT frameworks Ye et al.
-(2024b); Kuang et al.
+(2024c); Kuang et al.
 (2023); Zhang et al.
@@
 For instance, FedIT generally necessitates a sufficient amount of instruction data for fine-tuning, which is often a shortage in domain-specific fine-tuning contexts Zhang et al.
-(2024b).
-Figure 1: Visualization of different instruction augmentation strategies in the mathematical domain.
-The Random Sampling strategy randomly samples instructions from the public dataset.
-Direct Retrieval performs dense retrieval from the public dataset based on the client’s local data distribution without considering the cross-client domain coverage.
-FedDCA strategically selects client centers to maximize cross-client domain coverage and then performs domain data retrieval.
-The t-SNE visualization shows FedDCA has a better domain coverage than other sampling strategies.
-In this study, we investigate FedDIT, a novel approach within the FL paradigm aimed at boosting the performance of LLMs in specific domains.
-Unlike general FedIT, which enhances model effectiveness across diverse tasks without accounting for local data shortages, FedDIT addresses the unique challenge of clients only having limited domain-specific data.
-To overcome this, FedDIT enriches local data through a specific instruction augmentation strategy, crucial for effective instruction tuning and avoiding performance degradation.
-Our results reveal that, except for the code domain, where a standardized paradigm is primarily adhered to, relying solely on high-quality in-domain local data can be insufficient due to limited scale, leading to a decline in performance (see underlined values in Table 2).
-FedDIT aims to develop a domain-specific LLM through collaborative training and instruction augmentation while safeguarding client privacy, ensuring proficiency in tasks pertinent to its designated domain.
-For simplification and better instruction quality Zhang et al.
-(2024b); Toshniwal et al.
-(2024), we focus on a specific scenario of FedDIT involving a server-hosted public dataset, abstracting open-source instruction datasets on the web across multiple domains.
-This dataset is used for various instruction augmentation strategies, enhancing model performance in specific domains.
+(2024c).
+We explore Federated Domain-specific Instruction Tuning (FedDIT), an innovative approach that harnesses federated learning (FL) to unlock the potential of Large Language Models (LLMs) in specialized domains.
+A significant hurdle in deploying such models arises when multiple entities hold valuable but limited and privacy-sensitive data—a common scenario in fields like medical diagnosis Guan and Liu (2023); Hu et al.
+(2023) and financial risk assessment Abadi et al.
+(2024).
+These entities require collaborative training of domain-specific LLMs without direct data sharing Xu et al.
+(2024a); Zhang et al.
+(2024c).
+Our findings underscore a key limitation: exclusive reliance on local, in-domain data, despite its quality, often results in subpar performance due to insufficient scale (see Table˜1).
+FedDIT tackles this by integrating carefully designed instruction augmentation strategies (detailed in Section˜2) that expand local datasets while upholding privacy.
+This augmentation is not merely a supplement but a crucial enabler for robust instruction tuning and for averting performance decline.
+Focusing on practical and high-quality augmentation Zhang et al.
+(2024c); Toshniwal et al.
+(2024), we investigate a FedDIT setup employing a server-hosted public dataset (a cross-domain instruction set, detailed in Section˜B.1) and utilize various sampling strategies to realize the data augmentation (discussed in Section˜3).
+This dataset, abstracting diverse open-source instructions, provides a versatile foundation for augmentation techniques designed to elevate model performance within specific target domains.
 Additionally, the factors affecting FedDIT are still unclear.
@@
 Our results show that domain coverage significantly influences model performance within the corresponding domain.
-To maximize the cross-client domain coverage without compromising client data privacy, we propose a novel FedDIT algorithm, Federated Instruction Tuning of LLMs with Domain Coverage Augmentation, termed FedDCA.
-This algorithm employs a greedy client center selection process and implements instruction augmentation through dense retrieval on the server side.
-The fundamental idea of FedDCA is to select client centers to expand the diversity and coverage of augmented instruction datasets within a specific domain (as shown in Figure 1).
-By strategically optimizing domain coverage at each step, FedDCA efficiently constructs the augmented train set that enhances both the learning and generalization capabilities of the model, leading to superior performance on domain-specific tasks.
-Furthermore, to mitigate computational overhead on the client side and enhance the system scalability, we propose FedDCA, which employs heterogeneous encoders of different sizes and capacities.
-To achieve feature alignment, we train a projector on the server side using public data and employ contrastive learning techniques.
-We demonstrate the effectiveness and plug-and-play capability of FedDCA through comprehensive experiments conducted across four domains: code, medical, financial, and mathematical.
-These are compared against a range of baselines, which can be categorized into unaugmented and augmented methods.
-Results show that FedDCA outperforms other methods from at least 0.84% to the maximum of 29.19% in terms of model performance, with a relative domain coverage improvement from 4.82% to 21.36%.
-Extensive experiments across privacy analysis, computational efficiency and other further investigations are also conducted.
+To maximize cross-client domain coverage without compromising client data privacy, we propose a novel FedDIT algorithm, FedDCA (Domain Coverage Augmentation).
+The inability of direct data sharing under privacy constraints (detailed in Section˜5.1) motivates our approach.
+Initially, each client computes a set of candidate cluster centers from its local data; these serve as privacy-preserving proxies for its semantic distribution.
+FedDCA then employs a server-side, domain-coverage-oriented selection algorithm (detailed in Section˜5.2) to choose a strategic subset (denoted as client centers) from all uploaded candidate centers.
+These client centers then guide server-side dense retrieval for instruction augmentation.
+This selection process serves two critical purposes: 1) it provides a privacy-preserving way to capture the semantic diversity of distributed data, and 2) it enables efficient retrieval of relevant public instructions that complement the collective needs of the clients.
+By strategically optimizing the cross-client domain coverage, FedDCA efficiently constructs an augmented train set that enhances the model’s generalization capability, leading to superior performance on domain-specific tasks.
+To substantiate our claims, we conduct a rigorous empirical study.
+This includes foundational experiments identifying the key performance drivers in FedDIT, the development and evaluation of our proposed FedDCA algorithm across multiple specialized domains (medical, financial, and mathematical) against eleven baselines, and in-depth analyses of its practical aspects such as privacy, scalability, and robustness in challenging settings (detailed in Sections˜6 and B).
+Our key contributions are as follows:
+  • We empirically reveal a critical finding: in Federated Domain-specific Instruction Tuning (FedDIT), cross-client domain coverage, rather than data heterogeneity, substantially impacts LLM effectiveness.
+We propose a novel representation-based metric to quantify this cross-client domain coverage.
+  • We propose FedDCA, an algorithm that maximizes cross-client domain coverage through diversity-oriented client center selection and retrieval-based instruction augmentation.
+  • Extensive experiments across diverse domains demonstrate FedDCA’s superior effectiveness and plug-and-play capability, outperforming baselines by at most 29.19% in model performance and improving relative domain coverage by 4.82% to 21.36%.
+We also show its privacy-preserving capability in Section˜B.5.
+  • We demonstrate FedDCA’s robustness and practical applicability by evaluating its performance in data selection scenarios (achieves comparable performance using only 10% of baseline’s data), held-out settings where the server lack of task-specific public data, and its consistent effectiveness across varying inter-client data heterogeneity (Sections˜6.2, B and C).
+2 Related Work
+Federated Instruction Tuning.
+Instruction tuning has been widely applied across various application areas of large language models (LLM), serving as a key technique to enhance the capabilities and controllability of LLM Zhang et al.
+(2023c); Wei et al.
+(2022).
+Recently, federated instruction tuning (FedIT) has emerged as an effective strategy for the distributed optimization of LLMs, leveraging federated learning (FL) protocols to improve the handling of privacy-sensitive tasks in real-world scenarios.
+So far, several FedIT frameworks Ye et al.
+(2024c, b); Zhang et al.
+(2023b) have been established to evaluate the effectiveness of FedIT across multiple datasets, tasks, and FL methods.
+While these platforms provide a foundation for research, they have not yet introduced more complex federated algorithms and deeply investigated the challenging problems and factors affecting FedIT, which are crucial for advancing this field.
+Some progress has been made in the study of FedDIT Zhang et al.
+(2024c); Wang et al.
+(2024); Ye et al.
+(2024a), such as FewFedPIT, which addresses data scarcity by locally generating data using pre-trained LLMs and is the first to explore memory extraction attacks within FedIT.
+Another relevant approach is FedIT-U2S Ye et al.
+(2024a), which focuses on enabling FedIT when clients only possess raw, unstructured documents.
+Similar to our conceptualization of a server-hosted public dataset for augmentation, FedIT-U2S also leverages server-side resources (the example database and the LLM for instruction generation) to enrich client data, albeit with a focus on structuring initially unstructured data rather than augmenting existing instructions with diverse new ones.
+Non-IID data distribution is a common challenge in FL Karimireddy et al.
+(2020); Li et al.
+(2020); Sattler et al.
+(2019); Li et al.
+(2019, 2021).
+However, the impact of data heterogeneity on FedDIT has not been fully explored.
+To fully understand the impact of data heterogeneity on FedDIT and what truly matters in FedDIT, we conduct a comprehensive analysis in Section˜4.
+Domain Instruction Augmentation.
+In the real world, there is an urgent need for training LLMs with specific functionalities (e.g., reasoning capabilities) or domain-specific LLMs (e.g., code Nijkamp et al.
+(2023); Luo et al.
+(2024), medical Zhang et al.
+(2023d), financial Yang et al.
+(2023b, a); Zhang et al.
+(2023a); Wu et al.
+(2023), mathematical Yue et al.
+(2024); Luo et al.
+(2023)).
+Existing works tend to use open-source domain-specific instruction tuning datasets for training.
+However, the target domain may not always have corresponding ready-made domain-specific instruction datasets.
+Even if they exist, these datasets are often limited in scale.
+Augmentation strategies broadly fall into two categories: those leveraging existing/mined data and those generating new instructions.
+Methods based on existing or mined data include reusing human-curated public datasets Wang et al.
+(2023); Zhang et al.
+(2023e), retrieving from large instruction pools Xia et al.
+(2024); Jiao et al.
+(2023), or scaling instruction acquisition from the web Yue et al.
+(2024); Zhou et al.
+(2024).
+These approaches offer access to diverse, potentially high-quality instructions with generally favorable efficiency and privacy, but depend on the availability and relevance of such external data for the target domain.
+Conversely, generative methods Wang et al.
+(2022); Zhang et al.
+(2024c); Wang et al.
+(2024) create new instructions using LLMs (locally or via APIs).
+While potentially highly tailored, they often incur significant computational/API costs, raise privacy concerns, and face challenges in ensuring consistent instruction quality and diversity, as detailed in Appendix˜B.
+In summary, adapting existing instruction augmentation techniques to federated learning presents significant challenges in balancing data quality, diversity, efficiency, privacy, and scalability.
+Current methods often involve trade-offs between costly, quality-variable generative approaches and data-dependent retrieval strategies.
+This necessitates robust augmentation frameworks tailored for FedDIT, motivating our work in this area.
 3 Problem Formulation
 Federated Domain-specific Instruction Tuning (FedDIT) is a federated learning approach designed to improve the performance of LLMs in specific domains by utilizing limited cross-client private data in combination with domain-specific instruction augmentation strategies Zhang et al.
-(2024b); Xia et al.
+(2024c); Xia et al.
 (2024); Wang et al.
 (2022).
-We abstract the public data as a server-hosted multi-domain dataset (the composition is detailed in Table 1).
-The presence of a public dataset on the server is only one possible scenario, and the proposed algorithm is independent of the presence of a public dataset on the server.
-Even if the server does not have a public dataset, clients can retrieve public instructions by themselves based on the received client centers from the server or perform self-instruct, thereby achieving data augmentation.
+For practical and scalable augmentation within FedDIT, we conceptualize the primary source of augmentation instructions as a server-hosted, multi-domain public dataset (detailed in Section˜B.1).
+This strategic choice offers several key advantages in a federated context.
+Firstly, it provides a unified and manageable resource pool, abstracting the complex origins of public instructions (whether curated, mined, or pre-generated).
+Secondly, server management of this dataset decouples data sourcing from the core federated augmentation logic (such as retrieval strategies), allowing the latter to focus on optimizing instruction selection and distribution.
+Finally, a centralized public dataset enables efficient server-side pre-processing and indexing, which ensures consistent data quality and supports sophisticated retrieval mechanisms beneficial to all participating clients.
 Consider $N$ distributed clients, each with local private data $D_{k}^{l}$ of size $N_{k}^{l}$, and augmented data $D_{k}^{g}$ from the server’s public dataset $D^{p}$.
@@
 (2022) as the fine-tuning method, tuning additional parameters $\Delta\phi$ while keeping pre-trained LLM parameters $\phi$ frozen.
-The objective of FedDIT is to enhance the domain-specific performance of LLMs through FL without sharing private data Ye et al. (2024b); Zhang et al. (2024b, 2023b), which is defined as:
-$\small\operatorname*{arg\,min}_{\Delta\phi}\left\{F(\phi,\Delta\phi)\triangleq\sum_{k=1}^{N}p_{k}F_{k}\left(\phi,\Delta\phi_{k};D_{k}^{l},D_{k}^{g};\alpha_{k}\right)\right\},$ (1)
-where $F_{k}(\cdot)$ is the accumulated instruction fine-tuning loss of model $P_{\phi+\Delta\phi_{k}}$ computed on the client $c_{k}$’s augmented dataset $D_{k}^{l}\cup D_{k}^{g}$.
-$p_{k}$ is client $k$’s weight based on data ratio, and $\alpha_{k}=\frac{N_{k}^{p}}{N_{k}^{l}+N_{k}^{p}}$ is the ratio of public data.
-Eq.1 minimizes the summed empirical loss across clients’ augmented instructions to pursue the in-domain utility of the obtained global model.
-The client $c_{k}$’s empirical loss $F_{k}(\phi,\Delta\phi_{k};\mathcal{D})$ is calculated as:
-$F_{k}(\phi,\Delta\phi_{k};\mathcal{D})\triangleq\frac{1}{|\mathcal{D}|}\sum_{j=1}^{|\mathcal{D}|}l(P_{\phi+\Delta\phi_{k}};x_{j}),$ (2)
-where $x_{j}\in D,\forall j\in\{1,2,\dots,|D|\}$.
-The instruction tuning loss $l(\cdot;\cdot)$ on a sample $(x,y)$ is defined as $-\sum_{t=1}^{\left|y\right|}\log\left(w\left(y_{t}|x,y_{<t}\right)\right)$, where $x$ is the formated instruction and $y$ is the corressponding response.
-In summary, we formalize FedDIT as a novel FL framework that focuses specifically on domain-specific instruction tuning.
-In contrast, FedIT typically demands a substantial quantity of instructions for fine-tuning.
-Domain-specific instruction augmentation can be divided into two categories: 1) Generative methods, which generate instructions through the pre-trained LLM locally or external APIs.
-2) Retrieval-based methods, which retrieve instructions from the web.
-The former methods either compromise privacy or have high computational overhead.
-Thus, in this work, we focus on retrieval-based methods to utilize diverse and high-quality public data.
+The objective of FedDIT is to enhance the domain-specific performance of LLMs through FL without sharing private data Ye et al. (2024c); Zhang et al. (2024c, 2023b), which is defined as:
+$\operatorname*{arg\,min}_{\Delta\phi}\{F(\phi,\Delta\phi)\triangleq\sum_{k=1}^{N}p_{k}F_{k}\left(\phi,\Delta\phi_{k};D_{k};\alpha_{k}\right)\},$ (1)
+where $F_{k}(\cdot)$ represents the accumulated instruction fine-tuning loss of model $w_{\phi+\Delta\phi_{k}}$ evaluated on client $c_{k}$’s augmented dataset $D_{k}^{l}\cup D_{k}^{g}$.
+Here, $p_{k}$ denotes client $k$’s weight based on data ratio, and $\alpha_{k}=\frac{N_{k}^{p}}{N_{k}^{l}+N_{k}^{p}}$ represents the proportion of public data.
+Equation 1 optimizes the global model by minimizing the weighted sum of empirical losses across clients’ augmented instructions, thereby enhancing in-domain utility.
+The empirical loss for client $c_{k}$, $F_{k}(\phi,\Delta\phi_{k};\mathcal{D})$, is computed as $\frac{1}{|\mathcal{D}|}\sum_{j=1}^{|\mathcal{D}|}l(w_{\phi+\Delta\phi_{k}};x_{j})$, where $x_{j}\in D,\forall j\in\{1,2,\dots,|D|\}$ and $l(\cdot)$ is the instruction tuning loss function.
+For a client $c_{k}$ with private instruction set $\mathcal{D}_{k}$, we first encode instructions into embeddings using encoder $w_{\mathrm{enc}}$.
+These $d$-dimensional vectors are then clustered into $\xi$ groups via $k$-means, yielding centroid set $\mathcal{C}_{k}=\{\mathbf{c}_{k,1},\mathbf{c}_{k,2},\ldots,\mathbf{c}_{k,\xi}\}$, which is subsequently transmitted to the server.
+We first introduce the Direct Retrieval baseline, which performs independent dense retrieval Zhao et al.
+(2022) for each client without explicit consideration of cross-client domain coverage.
+For each centroid $\mathbf{c}_{k,j}$, the server retrieves the top-$\frac{N_{k}^{p}}{\xi}$ most similar public instructions using cosine similarity.
+The retrieved instructions are aggregated as client $c_{k}$’s augmented data $D_{k}^{p}=\bigcup_{j=1}^{\xi}\mathcal{R}_{k,j}$.
+Finally, client $c_{k}$ fine-tunes its model $w_{k}$ on the combined dataset $D^{k}=D_{k}^{l}\cup D_{k}^{g}$ of private and retrieved public instructions.
 4 What Truly Counts in FedDIT
-(a) Exp1. Code
-(b) Exp1. Med
-(c) Exp1. Fin
-(d) Exp1. Math
-(e) Exp2. Performance (%)
-(f) Exp2. Domain Coverage
-Figure 2: (a)-(d) Performance (%) of different heterogeneity in each domain with 10 and 100 clients.
-(e)-(f) Performance and domain coverage of iid and non-iid settings on different domains.
+(a) Exp1. Med
+(b) Exp1. Fin
+(c) Exp1. Math
+(d) Exp2. Performance (%)
+(e) Exp2. Domain Coverage
+Figure 1: (a)-(c) Performance (%) of different heterogeneity in each domain with 10 and 100 clients.
+(d)-(e) Performance and domain coverage of iid and non-iid settings on different domains.
 Experiments show that cross-client domain coverage significantly impacts model effectiveness, while data heterogeneity shows no monotonic correlation with performance.
-Data Heterogeneity Is Not Matter in FedDIT.
+4.1 Data Heterogeneity: A Minor Factor
 We adopt the Dirichlet distribution Wang et al.
 (2020); Yurochkin et al.
-(2019) to construct various heterogeneity ($\alpha=[0.01,0.1,1,10]$) and use k-means with $\xi=100$ to pseudo labeling instructions.
-We perform instruction tuning on both 10 and 100 clients with 2 randomly selected clients participating in each round.
+(2019) to construct various heterogeneity ($\beta=[0.01,0.1,1,10]$) and use k-means with $\xi=100$ to pseudo labeling instructions.
+We then perform instruction tuning on both 10 and 100 clients with 2 randomly selected clients participating in each round.
 For each domain, we only use the in-domain data and then perform FedIT.
-We perform the experiments 3 times with different random seeds (42, 43 and 44) and report the average performance and the standard deviation in each domain.
-As shown in Figures 2(a), 2(b), 2(c) and 2(d), the performance of LLM does not decrease due to the increase of data heterogeneity but shows a non-monotonic correlation, which indicates that the performance of LLM does not directly depend on data heterogeneity and other factors that play a key role.
-Domain Coverage: A Key Factor in FedDIT.
+We repeat the experiments for 3 times with different random seeds (42, 43 and 44) and report the average performance and the standard deviation in each domain.
+As shown in Figures˜1(a), 1(b) and 1(c), the performance of LLM shows a non-monotonic correlation with the data heterogeneity, which indicates that the performance of LLM does not directly depend on data heterogeneity and exist other factors playing a key role.
+4.2 Domain Coverage: A Key Factor
 Different from Explore-Instruct, which defines domain coverage through the distribution of verb-noun pairs, we attempt to conduct more in-depth and extensive experiments to study the effect of domain coverage on FedDIT.
@@
 Assume the dataset of in-domain data $D^{d}$ represents the latent data distribution of this domain and the cross-client data is defined as $D^{c}=\cup_{k=1}^{N}\left(D_{k}^{l}\cup D_{k}^{g}\right)$.
-Inspired by the facility location function Cornuejols et al. (1983), we define the domain coverage of $D^{c}$ respect to $D^{d}$ as follows:
-$\small d(D^{d},D^{c})=\frac{1}{\left|D^{d}\right|}\sum_{d\in D^{d}}\text{max}_{v\in(D^{c}\cap D^{d})}sim(d,v),$ (3)
-where $sim(d,v)$ is the cosine similarity between $d$ and $v$. Note that we only use the in-domain data in $D^{c}$ for the domain coverage calculation to avoid misleading.
-To better align with the real-world scenarios, we explore instruction augmentation based on both iid and non-iid cross-client data distribution and then perform dense retrieval for FedDIT.
+Inspired by the submodular function Krause and Golovin (2014), we first define a general coverage metric.
+For any two sets of item embeddings, $S_{1}$ (the reference set) and $S_{2}$ (the covering set), their coverage, denoted $d(S_{1},S_{2})$, is given by:
+$d(S_{1},S_{2})=\frac{1}{|S_{1}|}\sum_{s_{1}\in S_{1}}\max_{s_{2}\in S_{2}}\operatorname{sim}(s_{1},s_{2}),$ (2)
+where $\operatorname{sim}(\cdot,\cdot)$ is the cosine similarity between embeddings.
+Then the domain coverage of $D^{c}$ respect to $D^{d}$ is $d(D^{d},D^{c}\cap D^{d})$.
+The use of $D^{c}\cap D^{d}$ ensures that we only consider the in-domain portion of the aggregated client data for this particular coverage calculation, preventing misleading scores from out-of-domain client data, consistent with our original formulation.
+To better align with our setup, we explore instruction augmentation based on both iid and non-iid cross-client local data distribution and then perform dense retrieval Zhao et al.
+(2022) for FedDIT (the Direct Retrieval method).
 We set the number of clients to 10, while each client has 100 local instructions and obtains 1000 augmented public instructions from the server.
-For the iid setting, we randomly sample 1k from the public dataset and divide them into 10 shards as each client’s local data.
+For the iid setting, we randomly sample 1000 from the public dataset and divide them into 10 shards as each client’s local data.
 For the non-iid distribution, we perform k-means clustering with $\xi=100$.
-Each client randomly samples 100 instructions from randomly selected distinct clusters.
-Figures 2(e) and 2(f) presents the performance and the according domain coverage of FedDIT in different domains with iid and non-iid settings.
+Each client randomly samples 100 instructions from randomly selected distinct clusters and then performs retrieval for both settings.
+Figures˜1(d) and 1(e) presents the performance and the according domain coverage of FedDIT in different domains with iid and non-iid settings.
 We can observe that both iid and non-iid settings outperform in some domains, but both collectively indicate that higher domain coverage correlates with better performance.
 5 Method
-Figure 3: Overview of FedDCA, which consists of three stages: 1) The client $c_{k}$ performs local instructions clustering and sends cluster centers $\mathcal{C}_{k}$ to the server.
-2) The server first does greedy client center selection to maximize domain coverage (obtains $\mathcal{P}$) and performs client-center-based domain data retrieval, then sends the augmented instructions $D_{k}^{g}$ to the client $c_{k}$.
-3) Clients fine-tune the LLM collaboratively, and the LoRA parameters $\Delta\phi$ are exchanged between clients and the server.
-Based on the above empirical observations, we propose FedDCA, which enhances domain coverage to obtain a LLM that performs well on domain-specific tasks (shown in Figure 3).
-We first formulate and analyze the cross-client domain coverage optimization problem.
-Then, we introduce the FedDCA algorithm, which consists of two main stages: greedy client center selection and client-center-based domain data retrieval.
-For computational efficiency and system scalability, we introduce a variant of FedDCA with heterogeneous encoder settings, named FedDCA.
+Figure 2: Overview, a three-stage process.
+1) Client-side: Clients $c_{k}$ perform local instructions clustering and send cluster centers $\mathcal{C}_{k}$ to the server.
+2) Server-side Augmentation: The server executes diversity-oriented client center selection to maximize domain coverage (obtaining $\mathcal{P}$), retrieves augmented data $D_{k}^{g}$ based on these centers, and distributes it to clients.
+3) Collaborative Fine-tuning: Clients collaboratively fine-tune the LLM, exchanging LoRA parameters $\Delta\phi$ with the server.
+Based on the above empirical observations, we propose FedDCA, which enhances domain coverage to obtain a LLM that performs well on domain-specific tasks (shown in Figure˜2).
+We formulate the cross-client domain coverage optimization problem and then introduce the FedDCA algorithm.
 5.1 Optimization Problem
 As domain coverage directly affects the in-domain performance of the LLM, FedDCA aims to maximize the domain coverage of the cross-client augmented data $D^{c}$ with respect to the in-domain data distribution $D^{d}$.
-However, as clients can not send the local data to the server for privacy, directly finding a cross-client dataset that maximizes the domain coverage is unrealistic.
-Instead, finding a proper client center set $\mathcal{P}$ with optimized domain coverage and using it to perform instruction retrieval on the public data is an approximation problem.
-Suppose there exists a metric space $\mathcal{X}$, a set of cross-client local instruction embeddings $\mathcal{E}\in\mathcal{X}$, and a set of cluster centers $\mathcal{P}\in\mathcal{X}$.
-Each client $c_{k}$ has maximum $\xi$ clusters obtained by k-means algorithm Wu (2012).
-Additionally, in the federated setting, communication cost is always a critical factor.
-Thus, we formulate the optimization problem as follows:
-$\footnotesize\operatorname*{arg\,min}_{\mathcal{P}}$$\left\{\sum_{i=1}^{N}|\mathcal{P}_{i}|-\frac{1}{|D^{d}|}\sum_{d\in D^{d}}\left(\max_{p\in\mathcal{P}}\operatorname{sim}(d,p)\right)\right\},$ (4)
-s.t. $\quad|\mathcal{P}_{i}|\leq\xi,\quad\forall i\in\{1,2,\dots,N\}$ (5)
-The second term is the domain coverage of the selected client center set $\mathcal{P}$, and $sim(\cdot,\cdot)$ is the cosine similarity function. However, this optimization problem is NP-hard.
-For simplicity, we let $\xi$ be the constant $N$ (further discussed in Section 6.3).
-To solve this problem, the time complexity is $\mathcal{O}\left(C_{\xi N}^{N}\left(\mathcal{N}\left(N+N\log N\right)\right)\right)$, where $\mathcal{N}$ is the size of the public data $D^{p}$.
-As it is a factorial equation, the computational cost increases exponentially with $N$.
-Furthermore, the enormous amount of public data results in a large $\mathcal{N}$.
-To enhance computational efficiency, we propose a greedy algorithm as described below.
+However, privacy constraints prevent clients from sharing local data, meaning the server lacks direct knowledge of clients’ domain information.
+Consequently, directly identifying a cross-client dataset that maximizes domain coverage is infeasible.
+Instead, we leverage client-provided cluster centers ($\xi$ clusters per client, obtained by k-means algorithm Wu (2012)) as privacy-preserving proxies for local data.
+These centers, typically vector averages, effectively represent local distributions with manageable privacy risks (elaborated in Section˜B.5).
+The core challenge thus shifts to selecting an optimal set of these client centers, $\mathcal{P}$ (one per client), such that retrieving public instructions based on $\mathcal{P}$ enhances cross-client domain coverage.
+This formulation provides a tractable approximation to the original optimization problem under inherent privacy constraints.
+Additionally, in the FL setting, communication cost is always a critical factor. Thus, we formulate the optimization problem as follows:
+$\operatorname*{arg\,min}_{\mathcal{P}}\left\{\eta(\mathcal{P})-d(D^{d},\mathcal{P})\right\},$ (3)
+where the first term $\eta(\mathcal{P})=\sum_{i=1}^{N}|\mathcal{P}_{i}|$ represents the communication overhead of transmitting client centers.
+And the second term is the domain coverage of the selected client center set $\mathcal{P}$, defined in Eq.2.
+For simplicity, we let $\xi$ be the constant $N$ (further discussed in Section˜6.3). To better fit the FL environment and enhance the computational efficiency, we propose FedDCA as follows.
 5.2 FedDCA
-As the goal of FedDCA is to find a proper client center set $\mathcal{P}$ to optimize the objective defined in Eq.
-4, we propose a greedy algorithm as shown in Algorithm 1 to solve this problem in polynomial time and obtain a sub-optimal solution.
-Given the cluster centers $\{\mathcal{C}_{i}\}_{i=1}^{N}$ received from each client, which are obtained by clustering local instructions.
-FedDCA on the server consists of two main steps: greedy client center selection and client-center-based domain data retrieval.
-Naive Greedy Client Center Selection.
-In this paper’s FedDIT setting, a natural attempt to achieve greedy domain coverage optimization is using the public dataset $D^{p}$ on the server and maximizing $d(D^{d},\mathcal{P})$.
-Here, each client provides a set of candidate embeddings, and the server greedily picks (or updates) the cluster center to maximize domain coverage on $D^{p}$.
-While this approach is intuitive, it has three main limitations: 1) High computation complexity when $D^{p}$ is large.
-Each step of finding the best center $q^{*}$ (to add or update) must re-evaluate coverage w.r.t.
-$|D^{d}|$ data points, often incurring complexity on the order of $\mathcal{O}\bigl{(}N\xi|D^{d}|(N+N\log N)\bigr{)}$ per selection step.
-This quickly becomes prohibitively expensive for large $|D^{d}|$.
-2) Difficulty in domain coverage computation.
-The server may have limited knowledge about which domain each public instruction belongs to, or only know it partially.
-Additionally, the definition of domain is flexible without unified annotation standards - clients’ perceptions of domains may differ from the server’s understanding.
-This makes it challenging for the server to accurately compute domain coverage.
-3) No guaranteed public data.
-In many realistic FedDIT deployments, the server may not actually have an in-domain public dataset.
-Clients might rely on external retrieval or self-instruct generation for data augmentation, leaving no $D^{p}$ accessible on the server.
-The coverage-based greedy approach is then infeasible.
-Cluster-center-based Client Center Selection.
-To address these issues, we propose an alternative method that does not rely on public dataset.
-We denote the full set of these uploaded cluster centers $\{\mathcal{C}_{i}\}_{i=1}^{N}$ as $\mathcal{C}_{\text{all}}$, which is the proxy for domain coverage space.
-Since each client must pick one cluster center from $\mathcal{C}_{\text{all}}$.
-We then define the coverage score for client center selection:
-$\quad C(\mathcal{P})$ $=\sum_{z\,\in\,\mathcal{C}_{\text{all}}}\max_{p\in\mathcal{P}}\mathrm{sim}(z,p),$ (6)
-$where\quad\mathcal{P}$ =$=\{\,p_{1},\dots,p_{N}\,\},\quad p_{i}\in\mathcal{C}_{\text{all}},$ (7)
-and seek to optimize it. This avoids relying on $D^{d}$, greatly reducing coverage-computation overhead and allowing operation without a public dataset.
+Optimizing Eq.
+3 to find the ideal client center set $\mathcal{P}$ presents significant practical hurdles.
+First, an exhaustive search over all $C_{\xi N}^{N}$ candidate center combinations is computationally infeasible due to its exponential complexity.
+Second, directly leveraging a server-hosted public dataset $D^{p}$ to guide client center selection for optimal domain coverage is often impractical.
+This stems from several factors: 1) privacy constraints limit the server’s knowledge of individual client data distributions and their specific domain focuses; 2) the inherent ambiguity and hierarchical nature of ‘domains’ (e.g., ‘medical’ versus specific medical sub-tasks) complicate precise server-side domain coverage calculation, which is already costly over a potentially massive $D^{p}$; 3) the assumption of a relevant $D^{p}$ being available on the server may not hold, particularly if clients utilize external retrieval or local self-instruct for augmentation.
+These challenges underscore the need for a more pragmatic selection strategy.
+Therefore, we propose FedDCA.
+As detailed in Alg.
+1, FedDCA seeks a sub-optimal solution in polynomial time (discussed in Section˜A.1) and operates on the server through two main stages: 1) a coverage-oriented client center selection mechanism, and 2) client-center-based dense retrieval for data augmentation.
+Coverage-oriented Client Center Selection.
+To address these challenges, we propose a novel client center selection strategy that, crucially, does not rely on a server-side public dataset $D^{p}$.
+Let $\mathcal{C}_{\text{all}}=\bigcup_{i=1}^{N}\mathcal{C}_{i}$ be the set of all candidate cluster centers uploaded by the $N$ clients (where each client $c_{i}$ provides $\xi$ centers in $\mathcal{C}_{i}$).
+Our approach leverages $\mathcal{C}_{\text{all}}$ in two ways: 1) We use $\mathcal{C}_{\text{all}}$ as a proxy for the target domain distribution.
+Domain coverage is then approximated by evaluating how well a selected subset of centers covers $\mathcal{C}_{\text{all}}$ itself, making the computation tractable and independent of $D^{p}$.
+2) Instead of an exhaustive search, we employ an iterative, coordinate ascent-style algorithm (detailed in Alg.
+1) to select a final set $\mathcal{P}$ of $N$ client centers.
+This efficiently yields a high-quality, albeit potentially suboptimal $\mathcal{P}$.
+Based on the $\mathcal{C}_{\text{all}}$, the domain coverage could be calculated as $d(\mathcal{C}_{\text{all}},\mathcal{P})$.
 Specifically, each client $c_{i}$ locally runs $k$-means to produce $\xi$ candidate centers $\mathcal{C}_{i}$, then uploads these centers to the server.
 The server initializes a selection $\mathcal{P}$ by picking one cluster center from each client (thus $|\mathcal{P}|=N$) and computes the overall coverage solely based on these uploaded centers.
-In each iteration, the algorithm employs a coordinate-ascent style update: it removes the currently chosen center from one client’s portion of $\mathcal{P}$, tries replacing it with other candidate cluster centers, and keeps the replacement if it improves the coverage; this process is repeated until no further improvement is found (as shown in Algorithm 1).
-Algorithm 1 FedDCA: Greedy Client Center Selection
+In each iteration, FedDCA performs the following update (as shown in Alg.
+1): it removes the currently chosen center from $\mathcal{P}$, tries replacing it with other candidate cluster centers in $\mathcal{C}_{\text{all}}$, and keeps the replacement if it improves the cross-client domain coverage; this process is repeated until no further improvement is found.
+Algorithm 1 FedDCA: Client Center Selection
 1:Initialize:
@@
 3: Let $\mathcal{P}=\{p_{1},\dots,p_{N}\}$ and compute its coverage $C(\mathcal{P})$.
-4:$\Omega^{*}\leftarrow C(\mathcal{P})$
-5:for $i=1,\dots,N$ do
-6:    $\mathcal{P}_{-i}\leftarrow\mathcal{P}\setminus\{p_{i}\}$triangleright Remove $p_{i}$ from the current center set
-7:    $\nu^{*}\leftarrow p_{i}$
+[…]
+5:while True do
+6:  $\mathcal{P}_{-i}\leftarrow\mathcal{P}\setminus\{p_{i}\}$ $\triangleright$ Remove $p_{i}$ from $\mathcal{P}$
+7:  $\nu^{*}\leftarrow p_{i}$; $\Omega_{\text{old}}\leftarrow\Omega^{*}$
 8:  for $q\in(\mathcal{C}_{\text{all}}-\mathcal{P}_{-i})$ do
-9:         $C_{i}=C(\mathcal{P}_{-i}\cup\{q\})$
-10:         if $C_{i}>\Omega^{*}$ then
-11:             $\Omega^{*}\leftarrow C_{i}$; $\nu^{*}\leftarrow q$
+9:   $\mathrm{C}=d(\mathcal{C}_{\text{all}},\mathcal{P}_{-i}\cup\{q\})$ $\triangleright$ Defined in Eq. 2
+10:   if $\mathrm{C}>\Omega^{*}$ then
+11:     $\Omega^{*}\leftarrow\mathrm{C};\nu^{*}\leftarrow q$
 12:   end if
 13:  end for
-14:    $p_{i}\leftarrow\nu^{*}$; $\mathcal{P}\leftarrow\mathcal{P}_{-i}\cup\{p_{i}\}$; $C(\mathcal{P})\leftarrow\Omega^{*}$
-15:end for
-16:Output: The final center set $\mathcal{P}$ and its coverage $C(\mathcal{P})$.
-Since coverage is computed over just $N^{2}$ local centers (i.e., $\mathcal{C}_{\text{all}}$) rather than a large $D^{p}$, the computational overhead is much more manageable.
-Moreover, by not relying on a public dataset, this approach is applicable for various scenarios, including both data augmentation and data selection.
+14:  if $\Omega^{*}=\Omega_{\text{old}}$ then
+15:   break
+16:  end if
+17:  $p_{i}\leftarrow\nu^{*}$; $\mathcal{P}\leftarrow\mathcal{P}_{-i}\cup\{p_{i}\}$; $C(\mathcal{P})\leftarrow\Omega^{*}$; $i\leftarrow(i+1)\bmod N$
+18:end while
+19:Output: The selected client center set $\mathcal{P}$.
+This $D^{p}$-agnostic selection core is computationally tractable and versatile, effective for both data augmentation and targeted data selection through the selected client centers $\mathcal{P}$.
 Domain Data Retrieval.
-For each client center $p_{k}$, the server performs dense retrieval (detailed in Appendix A.2) on public dataset $D^{p}$ to get the top-$N_{k}^{p}$ similar public instructions, then sends retrieved public datasets $\{D_{1}^{g},\dots,D_{N}^{g}\}$ to each clients.
+For each client center $p_{k}$, the server performs dense retrieval Zhao et al.
+(2022) on public dataset $D^{p}$ to get the top-$N_{k}^{p}$ similar public instructions, then sends retrieved public datasets $\{D_{1}^{g},\dots,D_{N}^{g}\}$ to each clients.
 Specifically, to avoid the overlap between public data and local private data, we set a threshold $\alpha$ to filter the public instructions that have a similarity score larger than $\alpha$ with the client center.
-FedDCA: Heterogeneous Encoders.
-To enhance client-side computational efficiency and system scalability, we introduce FedDCA.
-This method employs a small encoder $\omega$ on the client side to reduce computational overhead and a larger encoder $\omega^{\prime}$ on the server side.
-However, the output dimension of heterogeneous encoders may not be consistent.
-Therefore, we introduce a projector $w_{p}$ on the server for feature alignment, which consists of two fully connected layers and a ReLU activation layer in between.
-We use contrastive learning Khosla et al.
-(2020) and train $w_{p}$ on the public instructions.
-For $\forall{x_{i}}\in D^{p}$, $\omega$ outputs embedding $h_{i}$, $\omega^{\prime}$ outputs embedding $h_{i}^{\prime}$ and $w_{p}$ outputs embedding $\varphi_{i}$.
-For input $x_{i}$, the positive sample pair is $(h_{i}^{\prime},\varphi_{i})$ and negative sample pairs are $\{\varphi_{j},j\neq i\}$.
-Let the batch size be $B$ and $\tau$ be the temperature parameter, then the training objective is defined as follows:
-$\footnotesize\mathcal{L}=-\sum_{i=1}^{B}\log\frac{\exp\left(\text{sim}(\varphi_{i},h^{\prime}_{i})/\tau\right)}{\sum_{j=1}^{B}\mathbb{I}_{[j\neq i]}\exp\left(\text{sim}(\varphi_{i},h^{\prime}_{j})/\tau\right)},$ (8)
-In summary, FedDCA leverages greedy client center selection and domain data retrieval to maximize cross-client domain coverage, leading to better model performance on domain-specific tasks.
-Additionally, FedDCA employs heterogeneous encoders with server-side feature alignment to reduce client-side computational overhead while maintaining effectiveness.
-5 Discussions
-Computation & Communication.
-Through greedy client center selection, FedDCA solves the optimization problem defined in Eq.
-4 in the polynomial time.
-As mentioned in Section 4.1, the time complexity of greedy client center selection is $\mathcal{O}\left(N^{4}\xi^{2}\left(1+\log N\right)\right)$.
-For domain data retrieval, its time complexity is $\mathcal{O}\left(N(\mathcal{N}\log\mathcal{N})\right)$.
-In addition, the heterogeneous encoder setting FedDCA further reduces the computation overhead on the client side.
-In the domain instruction augmentation stage, each client sends $\xi$ cluster centers to the server.
-The server returns $N_{k}^{p}$ retrieved public instructions to the client $c_{k}$.
-FedDCA follows standard FedIT procedures, efficiently exchanging LoRA parameters $\Delta\phi_{k}$ between clients and the server.
+In summary, through the coverage-oriented client center selection and domain data retrieval, FedDCA obtains a better cross-client domain coverage of the augmented data $D_{k}^{g}$, which leads to a better model performance on the target domain.
+For the Discussions part please refer to Section˜A.1.
+6 Experiments
+To demonstrate the effectiveness of FedDCA, we conduct extensive experiments across various domains and with several baselines. For additional results and analysis, please refer to Appendix˜B.
+Method MMLU-Med FPB FiQA TFNS GSM8K
+Zero-shot 70.60/- 55.94/- 18.54/- 59.21/- 23.27/-
+FedAvg McMahan et al. (2017) 68.40/0.6990 58.25/0.8529 14.18/0.8529 66.62/0.8529 47.46/0.7871
+FedProx Li et al. (2020) 69.10/- 56.51/- 14.90/- 66.45/- 47.15/-
+SCAFFOLD Karimireddy et al. (2020) 70.20/- 62.71/- 15.27/- 66.49/- 49.27/-
+FedAvgM Hsu et al. (2019) 64.70/- 68.14/- 29.27/- 70.32/- 46.85/-
+Random Sampling 71.30/0.7940 64.19/0.9196 13.09/0.9196 65.53/0.9196 47.38/0.8651
+Direct Retrieval 72.20/0.8830 66.31/0.9293 19.11/0.9293 67.62/0.9293 50.87/0.8967
+LESS Xia et al. (2024) 71.00/0.7737 60.56/0.8917 16.00/0.8917 61.14/0.8917 43.13/0.8352
+FewFedPIT Zhang et al. (2024c) 68.50/0.8250 56.30/0.8780 14.20/0.8780 59.10/0.8780 42.30/0.8380
+Self-Instruct Wang et al. (2022) 71.90/0.8586 59.73/0.9015 20.67/0.9015 66.54/0.9015 50.79/0.8811
+KnowledgeSG Wang et al. (2024) 73.13/0.9171 66.00/0.9490 32.72/0.9490 71.10/0.9490 51.20/0.9088
+FedDCA+FedAvg 74.50/0.9348 67.24/0.9815 35.27/0.9815 73.32/0.9815 52.46/0.9320
+FedDCA+FedProx 72.40/- 72.93/- 38.18/- 77.55/- 51.25/-
+FedDCA+SCAFFOLD 73.20/- 72.68/- 33.09/- 75.50/- 50.26/-
+FedDCA+FedAvgM 68.90/- 71.45/- 31.45/- 72.52/- 49.76/-
+Table 1: Performance (%) and domain coverage of FedDCA and other eleven baselines.
+Due to limited local data or inappropriate data augmentation strategy, occurs performance degradation in the Unaugmented and Random Sampling settings (underlined values).
+Different FL strategy does not affect domain coverage.
+We can see that FedDCA outperforms all other baselines.
+6.1 Experimental Setup
+Dataset and Evaluation Metrics.
+To evaluate the performance of FedDCA, we conduct experiments utilizing a server-hosted public dataset consisting of multiple domains’ data (detailed in Section˜B.1): Alpaca, MedAlpaca Zhang et al.
+(2023d), FinGPT Yang et al.
+(2023a), and MathInstruct Toshniwal et al.
+(2024).
+Performance is evaluated on MMLU-Med (medical); FPB, FiQA, and TFNS (financial); and GSM8K (math).
+The former two domains’ metric is Acc.
+and math’s is Exact Match.
+We set the number of clients to 10 by default, while each client has 100 local instructions and obtains 1k augmented public instructions by default from the server.
+Specifically, following the construction of the client’s local data heterogeneity in Section˜4, we set the $\beta$ to 0.1 by default.
+We also exhibits FedDCA’s strong robustness to different levels of data heterogeneity in Appendix˜C.
+Baselines.
+We compare FedDCA against a range of baseline methods, categorized into unaugmented and augmented approaches (further details are available in Appendix˜B).
+Unaugmented Methods.
+These include Zero-shot inference, where the LLM directly predicts without any fine-tuning, and FedIT.
+FedIT represents the application of standard federated learning to clients’ local data, for which we employ four widely recognized FL algorithms: FedAvg McMahan et al.
+(2017), FedProx Li et al.
+(2020), SCAFFOLD Karimireddy et al.
+(2020), and FedAvgM Hsu et al.
+(2019).
+Augmented Methods.
+This category encompasses several strategies.
+Random Sampling serves as a simple baseline, randomly selecting $N_{k}^{p}$ public data instances for each client.
+Direct Retrieval, a precursor to our method, performs dense retrieval from public data based on each client’s local instructions independently, without optimizing for cross-client domain coverage.
+LESS Xia et al.
+(2024) utilizes gradients from a warmed-up LLM on both training and a required validation set for its similarity-based retrieval.
+For generation-based approaches, Self-Instruct Wang et al.
+(2022) generates new instructions by prompting a powerful external LLM (e.g., GPT-3.5).
+FewFedPIT Zhang et al.
+(2024c) is a federated few-shot method where the global model is used for local in-context learning to synthesize data, aiming to improve both performance and privacy.
+Finally, KnowledgeSG Wang et al.
+(2024) employs a server-hosted expert model to generate and refine synthetic data, which clients then use to enhance their local models.
+6.2 Main Results
+Performance & Domain Coverage.
+Table˜1 showcases the performance and corresponding domain coverage of FedDCA against various baselines across three distinct domains.
+FedDCA consistently outperforms all eleven baselines, demonstrating substantial performance improvements of up to 29.19%.
+This superior performance is strongly correlated with its ability to achieve the highest domain coverage, surpassing other methods by an average of 4.82% to 21.36%.
+While unaugmented methods naturally struggle due to limited local data and thus restricted domain coverage, many existing augmentation strategies also exhibit significant drawbacks.
+1) Generation-based methods present a mixed picture: KnowledgeSG, leveraging a powerful server-hosted expert model, achieves strong results while FewFedPIT underperform even unaugmented baselines, potentially due to inconsistent synthetic data quality and diversity.
+Moreover, these approaches often entail significant resource demands: KnowledgeSG and FewFedPIT involve high computational costs (server-side and client-side), while Self-Instruct incurs monetary expenses for API calls with the privacy leakage.
+The potential lack of quality and diversity in synthetic instructions further constrains their effectiveness.
+2) Random Sampling and Direct Retrieval, while less costly, lack a strategic mechanism to ensure comprehensive cross-client domain coverage.
+3) LESS, while performs domain specific retrieval, requiring access to a validation set, and additional computation (warmup training).
+In conclusion, FedDCA’s strategic data augmentation substantially enhances domain coverage and model performance, efficiently circumventing the computational burden and inherent quality/redundancy challenges of data generation.
+Plug-and-Play.
+We demonstrate FedDCA’s plug-and-play capability by combining it with other augmentation methods.
+After sampling 1k instructions per client through different baselines, FedDCA performs retrieval by the selected client center (200 per client).
+As shown in Table˜2, despite using only 20% of the sampled data, FedDCA achieves comparable performance with baselines.
+This highlights FedDCA’s efficient data selection capability based on the cross-client domain coverage, which is further discussed in Section˜B.7.
++ FedDCA MMLU-Med FPB GSM8K
+Random Sampling 71.80 (+0.50) 63.92 (-0.27) 47.82 (+0.44)
+Direct Retrieval 72.10 (-0.10) 67.78 (+1.47) 52.53 (+1.66)
+LESS 70.80 (-0.20) 61.03 (+0.47) 43.92 (+0.79)
+Self-Instruct 70.70 (-1.20) 60.21 (+0.48) 50.25 (-0.54)
+Table 2: Plug-and-Play. Performance (%) of FedDCA with different data augmentation methods. FedDCA matches baseline’s performance by just 10% of its data through selection.
+6.3 Ablation Study
+Impact of Different Cluster Number.
+The hyperparameter $\xi$ is the number of clusters in the k-means algorithm.
+The experiment is conducted on $\xi=[N,2N,4N,8N]$, where $N$ is the number of clients.
+We report the domain coverage of the augmented dataset via FedDCA with different $\xi$ on the three domains in Table˜3.
+Results show that there is no best $\xi$ for all domains and $\xi=N$ is usually an acceptable choice.
+(a) Performance
+(b) Domain Coverage
+Figure 3: Impact of similarity threshold $\alpha$ on FedDCA’s (a) performance (%) and (b) domain coverage.
+Results highlight the role of the similarity threshold in enhancing diversity.
+Lighter bars indicate FedDCA.
+Similarity Threshold.
+Figure˜3 presents an ablation study comparing FedDCA (w similarity threshold $\alpha$) against FedDCA (w/o $\alpha$).
+FedDCA achieves superior performance and domain coverage, highlighting the crucial role of the similarity threshold in effectively amplifying the diversity of augmented instructions and subsequently boosting model efficacy.
+Note that the medical domain remains unchanged, as the similarity of all pairwise embeddings is below 0.7 (the default value for the similarity threshold $\alpha$).
+Domain $\xi=10$ $\xi=20$ $\xi=40$ $\xi=80$
+Med. 0.9348 0.9478 0.9466 0.9618
+Fin. 0.9815 0.9814 0.9819 0.9813
+Math. 0.9320 0.9348 0.9344 0.9337
+Table 3: Domain Coverage for Different $\xi$ Values. $\xi=N$ is usually an acceptable choice.
+We also investigate the impact of different retrieval amounts on FedDCA’s performance, detailed in Section˜B.4.
+7 Conclusion
+This work advances our understanding of Federated Domain-specific Instruction Tuning (FedDIT), establishing cross-client domain coverage as paramount over data heterogeneity for high performance.
+FedDCA provides a practical pathway to optimize this coverage, demonstrating significant gains through strategic data augmentation.
+FedDCA’s plug-and-play capability and validated robustness in challenging scenarios (e.g., scarcity of task-specific public data, large-scale deployments) further positions it as a reliable building block for real-world, privacy-preserving specialized LLM applications.
+These findings offer both an applicable solution and crucial guidance for future federated system design where domain coverage is the key.
+Limitations
+While FedDCA provides a robust and scalable framework for federated domain-specific instruction tuning, several avenues remain for further exploration.
+Although we have empirically demonstrated FedDCA’s resilience to the similar cross-client data distributions (see Appendix˜C), its performance may still be influenced by the quality and diversity of both the server-hosted public data and the client-provided candidate centers.
+Future work could investigate the adaptive client-side clustering to further enhance coverage and robustness.
+Additionally, automating hyperparameter selection (e.g., the number of clusters per client) and extending FedDCA to handle evolving data distributions or continual learning scenarios are promising directions to further improve its practicality and generalization in real-world deployments.
+Appendix A Methodology
+A.1 Discussions
+Computation.
+The client center selection process, as detailed in Alg.
+1, operates on $\mathcal{C}_{\text{all}}$, the set of all $N\cdot\xi$ candidate centers uploaded by $N$ clients (each providing $\xi$ centers).
+Since FedDCA is an iterative optimization algorithm, let $T$ be the number of passes until convergence.
+In each pass, the algorithm tries to swap the currently selected center $p_{i}$ with one of the other candidate centers from $\mathcal{C}_{\text{all}}$ ($N\xi-N$ choices).
+Each potential swap requires re-evaluating the domain coverage $\mathrm{C}$, which is $N\xi(N+N\log N)$.
+Therefore, the total complexity for one pass is approximately $O(N\xi\cdot(N\xi(N+N\log N)))=O(N^{3}\xi^{2}(1+\log N))$.
+If the algorithm converges in $T$ passes, the total complexity for client center selection is $O(T\cdot N^{3}\xi^{2}(1+\log N))$.
+Overall, this complexity is polynomial in the number of clients $N$ ($\xi=N$ by default), rendering the selection process tractable for typical federated learning scenarios where $N$ and $\xi$ are manageable.
+To empirically validate the convergence speed ($T$) of our client center selection algorithm, we conducted the client center selection for each domain under the default setup for 100 times.
+The results show that FedDCA converges remarkably quickly, with an average of 1.1N passes across all experiments.
+This rapid convergence indicates that the algorithm typically requires only one full iteration through all clients, with occasional minor adjustments in a second round of iteration, making it highly efficient in practice.
+Communication.
+The communication overhead of FedDCA is mainly incurred in two phases: (1) In the domain instruction augmentation phase, each client uploads $\xi$ cluster centers (typically low-dimensional vectors, e.g., 1024-d) to the server, and the server returns $N_{k}^{p}$ retrieved public instructions to each client.
+(2) During model parameter synchronization, FedDCA follows standard federated learning (FedIT) procedures, requiring only the exchange of LoRA parameters $\Delta\phi_{k}$, which significantly reduces communication cost.
 Privacy.
 Comparing FedDCA with other FedIT methods Zhang et al.
-(2024b); Ye et al.
-(2024b), the difference lies in the greedy client center selection.
+(2024c); Ye et al.
+(2024c), the difference lies in the client center selection stage.
 In this stage, the client only uploads the cluster center to the server, which is the average of embeddings to its cluster.
@@
 (2018), which allows the server to directly compute on ciphertext for matrix multiplication for dense retrieval.
-6 Experiments
-To demonstrate the effectiveness of FedDCA and its variant FedDCA, we conduct extensive experiments across various domains and with several baselines.
-For additional results and analysis, please refer to Appendix C.
-6.1 Experimental Setup
-Alpaca CodeAlpaca HumanEval MedAlpaca MMLU-Med
-$N_{sample}$ 52,002 20,022 164 33,955 1,089
-Metric - - Pass@1 - Acc
-FinGPT FPB FiQA TFNS MathInstruct GSM8K
-$N_{sample}$ 76,772 152 35 299 224,567 1,319
-Metric - Acc Acc Acc - Exact Match
+Robustness to Outliers.
+The k-means clustering, a component of our client-side processing, can be sensitive to outliers in the local data.
+If a client’s local dataset contains significant outliers, these could potentially skew the resulting cluster centers, and subsequently affect the server-side selection and retrieval process.
+To mitigate this, two main approaches could be considered.
+Firstly, more robust clustering algorithms from existing literature could be adopted by clients.
+For instance, Deshpande and Pratap, 2023 enhances k-means++’s robustness by capping the selection probability for new centers, based on the overall clustering cost and an allowed number of outliers.
+This prevents distant, potentially outlier, points from dominating the initialization process and improves clustering in noisy data.
+Secondly, clients could perform preliminary data filtering or cleaning steps before initiating the FedDCA process.
+This might involve simple heuristic-based filtering (e.g., keyword-based removal of irrelevant instructions) or leveraging lightweight LLM-based tagging to identify and exclude potential outliers Bernsohn et al.
+(2024); Biester et al.
+(2024); Zhang et al.
+(2024a).
+A comprehensive investigation into the impact of outliers and the optimal strategies for enhancing FedDCA’s robustness in such scenarios is a promising direction for future work.
+A.2 Theoretical Foundation of FedDCA
+FedDCA’s optimization objective is equivalent to the classic submodular facility location problem, ensuring our greedy algorithm achieves at least a 11/e1-1/e approximation ratio.
+The Optimization Problem in FedDCA.
+The overall optimization problem in our paper is formulated as Eq.
+3.
+In our specific problem setting, the number of cluster centers ($\xi$) per client is considered fixed.
+Therefore, minimizing the objective function is equivalent to maximizing the domain coverage term $d(D^{d},\mathcal{P})$ through the selected client centers $\mathcal{P}$.
+The Submodular Facility Location Problem.
+In its abstract form, we are given: a set of “locations” or “customers” $\mathcal{U}$ that need to be serviced; a set of potential “facilities” $\mathcal{F}$ that can be opened; a benefit function $w(i,j)$ that quantifies the value of servicing customer $j\in\mathcal{U}$ with facility $i\in\mathcal{F}$.
+The objective is to select a subset of facilities $\mathcal{S}\subseteq\mathcal{F}$ of a given size $k$ to maximize the total benefit provided to all customers:
+$g(\mathcal{S})=\sum_{j\in\mathcal{U}}\max_{i\in\mathcal{S}}w(i,j)$ (4)
+Formal Equivalence.
+The connection between our domain coverage problem and the facility location problem becomes clear when we map the components: 1) Customers ($\mathcal{U}$) $\rightarrow$ The set of in-domain data, $D^{d}$; 2) Facilities ($\mathcal{F}$) $\rightarrow$ The set of all candidate centers, $\mathcal{C}_{\text{all}}$; 3) Benefit ($w(i,j)$) $\rightarrow$ The cosine similarity between two centers, $\operatorname{sim}(p,c)$.
+Candidate centers are obtained by clustering each client’s local data, whereas $D^{d}$ is the union of all public in-domain data points.
+This “non-diagonal” mapping still matches the standard facility-location structure because every data point is evaluated against the chosen center set, regardless of origin.
+With this mapping, our domain coverage objective $d(D^{d},\mathcal{P})$ has the form of the classic facility location function $g(\mathcal{P})$ normalized by the size of the in-domain data, $|D^{d}|$.
+That is, $d(D^{d},\mathcal{P})=\frac{1}{|D^{d}|}g(\mathcal{P})$.
+Method
+Privacy
+Preserving
+API
+Cost
+Additional
+Information
+Performance
+Degradation
+Domain Coverage
+Oriented
+Computational Overhead of
+Data Augmentation
+FedAvg McMahan et al. (2017) -
+Random Sampling -
+LESS Xia et al. (2024) high
+FewFedPIT Zhang et al. (2024c) high
+KnowledgeSG Wang et al. (2024) high
+Self-Instruct Wang et al. (2022) low
+Direct Retrieval low
+FedDCA (ours) low
+Table 4: Key differences between FedDCA and other baselines.
+FedDCA demonstrates the ability of: 1) privacy-preserving, 2) no API cost, 3) no additional information required, 4) avoiding performance degradation, and 5) aiming at domain coverage optimization.
+LESS requires additional information, as it needs access to the validation set for gradient-based retrieval.
+Dataset Size Metric
+Alpaca 52,002 -
+MedAlpaca 33,955 -
+MMLU-Med 1,089 Acc.
+FinGPT 76,772 -
+FPB 152 Acc.
+FiQA 35 Acc.
+TFNS 299 Acc.
+MathInstruct 224,567 -
+GSM8K 1,319 Exact Match
 Table 5: Dataset information of each domain.
-The test set is marked by $\ddagger$.
-The public dataset is composed of these five train sets.
-We denote HumanEval and MMLU-Med as H-Eval and M-Med for short.
-Dataset and Evaluation Metrics.
-As shown in Table 1, to evaluate the performance of FedDCA, we conduct experiments on four domains: code, medical, financial, and mathematical.
-We set the number of clients to 10 by default, while each client has 100 local instructions and obtains 5k augmented public instructions from the server.
-Baselines.
-We compare FedDCA and FedDCA with the following baselines: 1) Unaugmented methods, including zero-shot inference and FedIT, which are composed of four widely used FL methods, including FedAvg McMahan et al.
-(2017), FedProx Li et al.
-(2020), SCAFFOLD Karimireddy et al.
-(2020) and FedAvgM Hsu et al.
-(2019).
-2) Augmented methods, including Random Sampling, Direct Retrieval, LESS Xia et al.
-(2024), and Self-Instruct Wang et al.
-(2022).
-Specifically, Random Sampling randomly samples $N_{k}^{p}$ public data for each client, while Direct Retrieval is a simplified version of FedDCA that directly retrieves public data based on each client’s cluster center set without greedy client center selection, not considering the cross-client domain coverage.
-6.2 Main Results
-Performance.
-Table 2 shows the performance and the corresponding domain coverage of FedDCA and other baselines on four domains.
-FedDCA outperforms the other nine baselines in all domains, with a substantial improvement from at least 0.84% to the maximum of 29.19% over other baselines.
-In particular, FedDCA+FedAvg and FedDCA+FedProx perform better across four domains.
-We see that although FedDCA has a performance drop than FedDCA, it still outperforms other baselines.
-We can observe that FedIT methods perform well in the code domain, even better than most augmented methods.
-This could be attributed to two reasons: 1) The code domain follows a certain paradigm.
-2) With limited local data, unaugmented methods learn the same data more times given the same epoch and batch size, which may be an unfair comparison.
-Nonetheless, FedDCA still surpasses the best baseline in the code domain.
-Domain Coverage.
-FedDCA achieves the highest domain coverage across all domains, surpassing other baselines by 4.82% to 21.36% on average.
-Although FedDCA uses a smaller encoder on the client side for computational efficiency, it results in a slight 2.89% drop in average domain coverage but still outperforms the best baseline by 2.78% on average.
-These results demonstrate the effectiveness of both FedDCA and FedDCA in enhancing domain coverage.
-Method H-Eval M-Med FPB FiQA TFNS GSM8K
-Zero-shot 29.88/- 70.60/- 55.94/- 18.54/- 59.21/- 23.27/-
-FedAvg McMahan et al. (2017) 39.03/0.8126 68.40/0.6990 58.25/0.8529 14.18/0.8529 66.62/0.8529 47.46/0.7871
-FedProx Li et al. (2020) 37.20/- 69.10/- 56.51/- 14.90/- 66.45/- 47.15/-
-SCAFFOLD Karimireddy et al. (2020) 37.80/- 70.20/- 62.71/- 15.27/- 66.49/- 49.27/-
-FedAvgM Hsu et al. (2019) 32.32/- 64.70/- 68.14/- 29.27/- 70.32/- 46.85/-
-Random Sampling 32.93/0.8512 71.30/0.7940 64.19/0.9196 13.09/0.9196 65.53/0.9196 47.38/0.8651
-Direct Retrieval 34.14/0.9396 72.20/0.8830 66.31/0.9293 19.11/0.9293 67.62/0.9293 50.87/0.8967
-LESS Xia et al. (2024) 28.04/0.8509 71.00/0.7737 60.56/0.8917 16.00/0.8917 61.14/0.8917 43.13/0.8352
-Self-Instruct Wang et al. (2022) 32.92/0.8966 71.90/0.8586 59.73/0.9015 20.67/0.9015 66.54/0.9015 50.79/0.8811
-FedDCA 34.75/0.9532 73.30/0.8972 67.10/0.9538 30.54/0.9538 71.01/0.9538 51.55/0.9096
-FedDCA+FedAvg 36.58/0.9766 74.50/0.9348 67.24/0.9815 35.27/0.9815 73.32/0.9815 52.46/0.9320
-FedDCA+FedProx 32.92/- 72.40/- 72.93/- 38.18/- 77.55/- 51.25/-
-FedDCA+SCAFFOLD 39.87/- 73.20/- 72.68/- 33.09/- 75.50/- 50.26/-
-FedDCA+FedAvgM 33.53/- 68.90/- 71.45/- 31.45/- 72.52/- 49.76/-
-Table 2: Performance (%) and domain coverage of FedDCA, FedDCA and other nine baselines.
-Due to limited local data or inappropriate data augmentation strategy, occurs performance degradation in the Unaugmented and Random Sampling settings (underlined values).
-FedDCA outperforms all other baselines.
-Plug-and-Play.
-To demonstrate the plug-and-play capability of FedDCA with other augmentation methods, we first perform instruction sampling through different baselines to obtain augmented data (5k instructions per client) and then utilize FedDCA to conduct data selection (select 1k instructions per client) for FedDIT.
-As shown in Table 3, using FedDCA in combination with other augmentation methods still achieves comparable performance despite using only 20% of the sampled data.
-FedDCA efficiently selects the most representative samples based on domain coverage, enabling effective training with less data while maintaining performance.
-Computation Analysis.
-To show the computational efficiency of FedDCA, we compare these to FedDCA from two aspects: 1) Encoder model size.
-2) Encoding time overhead.
-We test the encoders’ encoding time on the public dataset.
-As shown in Figure 6(a), FedDCA requires only about 30% encoded time overhead of FedDCA.
-Moreover, FedDCA achieves a substantial reduction in model size, which is less than 10% of FedDCA’s.
-Privacy Analysis.
-We evaluate the privacy-preserving capability of different ratios of public data against memory extraction attacks Carlini et al.
-(2021); Zhang et al.
-(2024a), which utilizes the autoregression nature of LLMXu et al.
-(2024).
-We analyze the domain inference attack in Section C.3.
-We focus on one client’s instruction tuning in FedDIT, using FedDCA for instruction augmentation with 1k to 5k public instructions.
-We set up 10 clients with full participation for 10 rounds.
-Specifically, we record the average ROUGE-L score Lin (2004) denoted as $\rho$ for client $c_{0}$ in each round.
-The prompt to extract instructions is in Section C.1.
-For each setting, we repeat memory extraction 100 times and report $\rho$.
-Specifically, denote the generated $\mathcal{N}$ instructions as $\mathcal{I}$ and the client’s local instructions $\mathcal{I}^{l}$.
-The calculation is defined as $\rho=\frac{1}{\mathcal{N}}\sum_{i=1}^{\mathcal{N}}\texttt{ROUGE-L}(\mathcal{I}_{i},\mathcal{I}^{l})$.
-(a) Different amounts of augmented public data for FedDIT.
-(b) $\rho$ per round. Dotted lines represent the base-data-only setting.
-Figure 4: Privacy analysis of memory extraction attacks. Data augmentation can effectively mitigate the privacy leakage risk.
-Figure 4(a) shows no significant correlation between the public data ratio and privacy-preserving capability in the same training round.
-In addition, only using local data has a higher risk of privacy leakage than augmented methods.
-Additionally, Figure 4(b) shows $\rho$ trends per round.
-Initially, the $\rho$ for augmented settings increases, then decreases or converges, while the base-data-only scores continue to rise, especially in the code domain.
-This indicates that with more training rounds, base-data-only fine-tuning captures more privacy information, while the privacy leakage risk in augmented fine-tuning decreases or converges.
-Scalability.
-We scale the client number to 100 and 2 clients are randomly selected for FedDIT in each round.
-As the number of clients increases, the amount of local data on each client gradually grows.
-Table 4 shows that the model trained on only base data even outperforms Random Sampling in domains other than code and narrows the gap with FedDCA.
-However, because FedDCA aims to maximize the cross-client domain coverage, it achieves higher domain coverage and better performance.
-Data Selection Scenarios.
-We conduct the experiments on two settings: A) For each domain’s FedDIT, we randomly dispatch the in-domain data equally to each client.
-For example, in the code domain, as the size of the public dataset is 20,022, each client is assigned about 2,000 samples.
-Then we perform FedDIT on each client’s whole data.
-B) Following A’s setting, we first perform FedDCA to determine the client center set, and then we conduct the data selection based on the client center set through dense retrieval to select 200 samples for each client.
-Finally, we perform FedDIT on the selected data.
-Figure 6(b) demonstrates that FedDCA can achieve comparable performance to training with the full dataset by using only 10% data through data selection.
-Held-out Setting.
-If distributed clients aim to solve tasks based on existing knowledge, the public dataset will inevitably contain knowledge relevant to those domains.
-This could come from the original corpus (which can be converted into instruction-response pairs by GPT) or from pre-constructed instruction datasets on the website.
-So the distribution of the public dataset can be categorized as follows: containing held-in or held-out instructions.
-The held-in indicates that the public dataset contains instructions for the specific task that clients aim to solve, while the held-out indicates that the public dataset does not contain this task’s instructions.
-The paper’s default setting is the held-in setting.
-Considering the held-out setting in the financial domain, given that the training set FinGPT and the test sets FPB, FiQA, and TFNS are all related to sentiment analysis tasks.
-We keep the setting of test sets and replace the FinGPT’s instructions in public data with data from a financial QA dataset (Sujet).
-The clients’ local data are still randomly sampled from FinGPT.
-This approach yields held-out public data.
-Figures 6(d), 6(e) and 6(f) shows, FedDCA still achieves performance improvements compared to other baselines for the held-out setting.
-Additionally, using the Random Sampling data augmentation strategy resulted in performance degradation on the FiQA dataset.
-This further underscores the necessity of selecting an appropriate data augmentation strategy.
-6.3 Ablation Study
-Effect of Retrieval Number.
-We report the performance and the corresponding domain coverage of FedDCA with different retrieval amounts on the four domains in Figure 5(b).
-We can see that the domain coverage of FedDCA is increasing along with the retrieval amount in different trends, as well as the performance except for the code domain.
-Noted that although the code domain is more about following a certain paradigm, which could perform well with a few data and more fine-tuning rounds, it still could benefit from the instruction augmentation.
-Similarity Threshold.
-Figure 5(a) shows the performance and the domain coverage of FedDCA w and w/o similarity threshold $\alpha$, where FedDCA w/o $\alpha$ is marked as FedDCA.
-Results show that FedDCA is slightly better than FedDCA in code, financial, and mathematical domains.
-Impact of Different Cluster Number.
-The hyperparameter $\xi$ is the number of clusters in the k-means algorithm.
-The experiment is conducted on $\xi=[N,2N,4N,8N]$, where $N$ is the number of clients.
-We report the domain coverage of the augmented dataset via FedDCA with different $\xi$ on the four domains in Figure 6(c).
-Results show that there is no best $\xi$ for all domains and $\xi=N$ is usually an acceptable choice.
-+ FedDCA H-Eval M-Med FPB GSM8K
-Random Sampling 32.51 (-0.42) 71.80 (+0.50) 63.92 (-0.27) 47.82 (+0.44)
-Direct Retrieval 33.89 (-0.25) 72.10 (-0.10) 67.78 (+1.47) 52.53 (+1.66)
-LESS 28.32 (+0.28) 70.80 (-0.20) 61.03 (+0.47) 43.92 (+0.79)
-Self-Instruct 33.15 (+0.23) 70.70 (-1.20) 60.21 (+0.48) 50.25 (-0.54)
-Table 3: Plug-and-Play. Performance (%) of FedDCA with different data augmentation methods. FedDCA can match full dataset performance using just 10% of the data through selection.
-Metric Zero-shot Base Data Random Sampling FedDCA
-H-Eval 29.88/- 34.14/0.8282 34.75/0.8685 35.92/0.9242
-M-Med 70.60/- 72.40/0.8377 69.90/0.8497 73.30/0.9090
-FPB 55.94/- 66.74/0.9339 61.05/0.9408 67.16/0.9800
-GSM8K 23.27/- 49.12/0.8709 47.23/0.8812 50.26/0.9118
-Table 4: Scalability. Performance (%) and domain coverage of FedDCA and other baselines. 100 clients with 2 clients per round. Results show the strong scalability of FedDCA.
-(a) W and w/o $\alpha$
-(b) Different Retrieval Amounts
-Figure 5: Solid bars and lines are performance.
-Performance (%) and domain coverage of FedDCA (a) w and w/o similarity threshold $\alpha$; (b) w different retrieval amounts.
-Experiments show the necessity of similarity threshold for amplifying diversity and the strong applicability of FedDCA with various augmentation scale.
-(a) Computational Cost
-(b) Data Selection Scenarios
-(c) Different $\xi$
-(d) Held-out. FPB
-(e) Held-out. FiQA
-(f) Held-out. TFNS
-Figure 6: (a) Computational cost comparison.
-(b) Domain coverage for different $\xi$.
-(c) Data selection scenarios.
-(d-f) Performance in the held-out setting.
-Results show the efficiency of FedDCA, the effectiveness of FedDCA in data selection scenarios, and the robustness of FedDCA in the held-out setting.
-7 Conclusion
-We reveal that in FedDIT, data heterogeneity has a non-monotonic relationship with model performance, while cross-client domain coverage significantly affects model effectiveness.
-To address this, we propose FedDCA, which optimizes domain coverage through greedy client center selection and retrieval-based instruction augmentation.
-The selection process ensures comprehensive cross-client domain coverage by greedily picking centers that represent maximal distinct regions of the instruction space.
-FedDCA further reduces client-side computation and enhances scalability using heterogeneous encoders.
-Extensive experiments comparing FedDCA with other nine baselines and across four domains demonstrate its effectiveness and plug-and-play capability.
-Appendix
-Appendix A Methodology
-A.1 Complementary Description
-Notation Table.
-For convenience, we provide a detailed notation table in Table 5.
-NOTATIONS REMARK
-$N,C$ Clients number, client set $C=\{c_{1},c_{2},\ldots,c_{N}\}$.
-$D^{p},D^{d},D^{c},D^{l}$ The public datasets, the in-domain data, the cross-client dataset, client’s local private data.
-$N_{k}^{l},N_{k}^{p}$ Number of local private data on the $k$-th client, number of the retrieved public data on the $k$-th client.
-$D_{k}^{l},D_{k}^{g},D_{k}$ The local private data on the $k$-th client, the retrieved public data on the $k$-th client, the augmented dataset on the $k$-th client.
-$\Lambda,\mathcal{P},\mathcal{C}$ A specific sampling strategy that performs instruction augmentation on the server side, selected client center set, the cluster centers obtained locally and sent to the server for the greedy client center selection.
-$\phi,\Delta\phi$ LLM’s pre-trained parameters, additional LoRA parameters.
-$w,w_{enc},w_{p}$ Merged model parameters from the frozen LLM’s parameters $\phi$ and the additional LoRA parameters $\Delta\phi$, encoder model, projector model.
-$F_{k}(w;\mathcal{D}),l(w;x,y)$ Loss of model $w$ over a specific dataset $\mathcal{D}$, the instructon tuning loss of model $w$ over a data sample $\left(x,y\right)$.
-Table 5: Frequently used notation.
-A.2 Direct Domain Data Retrieval
-In this section, we first show the detail of instruction-based dense retrieval, which is both used in direct retrieval and FedDCA. Then, we explain the algorithm of the method direct retrieval.
-A.2.1 Instruction Based Dense Retrieval
-[…]
-A.2.2 Direct Retrieval
-The direct domain data retrieval only utilizes instructions of the client’s local data without responses to perform the retrieval-based domain data augmentation.
-The detailed algorithm is described in Algorithm 2.
-For each client, we start by encoding the local instructions using the encoder model $w_{enc}$.
-Next, we apply the k-means algorithm to cluster the embeddings into $\xi$ clusters.
-Then $\xi$ cluster centers are sent to the server for retrieval-based domain data augmentation.
-Subsequently, the retrieved public data is sent to the client for instruction tuning.
-Algorithm 2 Direct domain data retrieval
-1:Parameters:
-2:Clients’ local datasets $D=\{D_{1},D_{2},\ldots,D_{N}\}$; Public dataset $D^{p}$; Local datasets $D^{l}=\{D^{l}_{1},D^{l}_{2},\ldots,D^{l}_{N}\}$; Number of clusters $\xi$; Encoder model $w_{enc}$; Pre-encoded public instruction embeddings $\mathcal{E}$.
-3:$D^{g}\leftarrow\emptyset$
-4:for $i\in\{1,2,\ldots,N\}$ do
-5:     $\mathcal{E^{\prime}}\leftarrow\{w_{enc}(x)\mid x\in D_{i,instruction}^{l}\}$triangleright Encode the local instructions
-6:     $\mathcal{C}\leftarrow\text{k-means}(\xi).fit(\mathcal{E^{\prime}})$triangleright Cluster local instructions, return cluster centers $\mathcal{C}$
-7:     $\mathcal{S}\leftarrow\mathcal{C}\cdot\mathcal{E}^{T}$triangleright Compute the similarity score between $\mathcal{C}$ and $\mathcal{E}$
-8:     $\mathcal{I}\leftarrow\emptyset$
-9:     for $j\in\{1,2,\ldots,\xi\}$ do
-10:         $\mathcal{S}^{\prime}\leftarrow\{s\mid s\in S_{j},\text{index}(s)\notin\mathcal{I}\}$triangleright Filter the selected indices
-11:         $\mathcal{I^{\prime}}\leftarrow\text{indices of the top-}\frac{N_{k}^{p}}{\xi}\text{ elements in }\mathcal{S}^{\prime}$triangleright Retrieve the top $\frac{N_{k}^{p}}{\xi}$ indices
-12:         $\mathcal{I}\leftarrow\mathcal{I}\cup\mathcal{I^{\prime}}$triangleright Update the selected indices
-13:     end for
-14:     $D^{g}_{i}\leftarrow\{D^{p}_{j}\mid j\in\mathcal{I}\}$triangleright Selected public instructions
-15:     $D_{i}\leftarrow D^{g}_{i}\cup D_{i}^{l}$triangleright Obtain the augmented instruction dataset $D_{i}$
-16:end for
-Appendix B Related Works
-Federated Instruction Tuning.
-Instruction tuning has been widely applied across various application areas of large language models (LLM), serving as a key technique to enhance the capabilities and controllability of LLM Zhang et al.
-[2023c]; Wei et al.
-[2022].
-Recently, federated instruction tuning (FedIT) has emerged as an effective strategy for the distributed optimization of LLMs, leveraging federated learning (FL) protocols to improve the handling of privacy-sensitive tasks in real-world scenarios.
-So far, several FedIT frameworks Ye et al.
-[2024b, a]; Zhang et al.
-[2023b] have been established to evaluate the effectiveness of FedIT across multiple datasets, tasks, and FL methods.
-While these platforms provide a foundation for research, they have not yet introduced more complex federated algorithms and deeply investigate the challenging problems and factors affecting FedIT, which are crucial for advancing this field.
-PrivateLoRA Wang et al.
-[2023b] addresses privacy and efficiency issues by exploiting the low-rank properties of residual activations to reduce communication costs, significantly lowering the communication overhead through collaborative computation between the server and clients while effectively maintaining the privacy of local data.
-While FewFedPIT Zhang et al.
-[2024b] focuses on the few-shot learning setting in FedIT, using self-generated data by pre-trained LLMs locally to mitigate the paucity of data and first discussing memory extraction attacks within FedIT.
-Non-IID data distribution is a common challenge in FL Karimireddy et al.
-[2020]; Li et al.
-[2020]; Sattler et al.
-[2019]; Li et al.
-[2019, 2021].
-However, the impact of data heterogeneity on FedIT has not been fully explored.
-Shepherd Zhang et al.
-[2023b] constructing heterogeneity based on the topic of general instruction datasets.
-Results demonstrate that, unlike the consensus of traditional FL, for general FedIT, heterogeneity has a positive effect.
-By aggregating diverse instructions from clients, the diversity increases, thereby enhancing the model’s adaptability to various tasks.
-However, it just scratches the surface and does not explore issues in FedDIT.
-To fully understand the impact of data heterogeneity on FedDIT, we conduct a comprehensive analysis in this work and unveil a crucial finding: there is no monotonic correlation between the degree of non-independent and identically distributed (non-iid) and LLM’s performance in the context of FedDIT (Section 3).
-Domain Instruction Augmentation.
-In the real world, there is an urgent need for training LLMs with specific functionalities (e.g., reasoning capabilities) or domain-specific LLMs (e.g., code Nijkamp et al.
-[2023]; Luo et al.
-[2024], medical Zhang et al.
-[2023d], financial Yang et al.
-[2023b, a]; Zhang et al.
-[2023a]; Wu et al.
-[2023], mathematical Yue et al.
-[2024]; Luo et al.
-[2023]).
-Existing works tend to use open-source domain-specific instruction tuning datasets for training.
-However, the target domain may not always have corresponding ready-made domain-specific instruction datasets.
-Even if they exist, these datasets are often limited in scale.
-Several studies investigate domain-specific instruction augmentation, which can be categorized into three aspects: 1) Reusing human-curated public datasets Wang et al.
-[2023a]; Zhang et al.
-[2023e]; Jiao et al.
-[2023]; Lee et al.
-[2024]; Xia et al.
-[2024].
-For instance, Parrot Jiao et al.
-[2023] enhances translation capabilities of LLMs in chat by converting bilingual sentences into instruction-following datasets.
-Furthermore, works like INSTA Lee et al.
-[2024] and LESS Xia et al.
-[2024] attempt efficient domain-specific instruction augmentation via dense retrieval.
-INSTA Lee et al.
-[2024] uses instructions without responses for effective retrieval.
-LESS Xia et al.
-[2024] assumes access to a validation set and uses the warmup LLM’s gradients of the train and validation sets for retrieval.
-2) Using seed tasks for self-instruct Luo et al.
-[2024]; Wan et al.
-[2023]: Explore-Instruct Wan et al.
-[2023], for example, employs activate exploration to tree-model domain tasks from both depth and breadth, increasing seed instructions’ coverage of the domain, and subsequently generating broader in-domain instructions through self-instruct.
-3) Scaling instructions from the web: Recent works Yue et al.
-[2024]; Zhou et al.
-[2024] highlight the immense potential of mining naturally occurring instructions from the internet.
-Compared to generated data, web-mined instructions exhibit less bias and greater diversity.
-MAmmoTH2 Yue et al.
-[2024] firstly retrieves domain-relevant texts and employs a LLM to extract Q-A pairs, further refining them into instruction-response pairs.
-Jiuzhang3.0 Zhou et al.
-[2024] distills GPT’s instruction generation capabilities into a smaller model and then uses it to generate instructions from the internet.
-In conclusion, models fine-tuned with augmented instructions have shown promising domain-specific capabilities.
-To obtain a well-performing LLM in the specific domain within the distributed environment, we utilize a multi-domain dataset as public data on the server side and perform domain-specific instruction augmentation based on the client’s local instructions.
-Appendix C Experiments
-C.1 Implementation Details
-We consider FedDIT in the cross-device scenario, $N$ = 10 clients, $\mathcal{R}$ = 30 rounds, where we randomly sample 2 clients to be available for each round.
-Then, each available client performs FedDIT for 10 steps with AdamW optimizer, and the batch size is $B=32$ in a round.
-The initial learning rate is $55e-$ 5 with a cosine learning rate scheduler.
-Our experiment utilizes the widely used LLM, Llama3-8B [^2] https://huggingface.co/meta-llama/Meta-Llama-3-8B as the base model with 2048 max sequence length and adopts LoRA tuning method.
-The rank of LoRA is 16, and the scalar alpha is 16.
-For k-means Wu [2012], we set cluster num $\xi$ = 10 and for FedDCA we set the similarity threshold $\alpha$ = 0.7.
-For FedDCA, we set the temperature parameter $\tau$ = 0.5 for contrastive learning Khosla et al.
-[2020].
-We utilize bge-large-en-v1.5333https://huggingface.co/BAAI/bge-large-en-v1.5 as both the client and server’s encoder as default, which outputs embeddings of 1024 dimensions.
-While we utilize all-MiniLM-L6-v2444https://huggingface.co/sentence-transformers/all-MiniLM-L6-v2 as the client’s small encoder which outputs embeddings of 384 dimensions and bge-large-en-v1.5 as the server’s encoder for FedDCA.
-Projector.
-To reduce the computational overhead of clients, we propose a heterogeneous encoder method called FedDCA.
-We utilize a projector to perform the dimension alignment between the small encoder on the client side and the large decoder on the server side.
-As shown in Figure 7, where $d_{l}$ is the dimension of the client-side encoder, $d_{g}$ is the dimension of the server-side encoder.
-Figure 7: The projector used on the server side for feature alignment, which consists of two fully connected layers and a ReLU activation layer in between.
-Prompts Used in the Self-Instruct Data Generation.
-To generate the Self-Instruct data, we prompt GPT-3.5 to generate the instruction with the designed prompt in Figure 8.
-Specifically, we randomly sample two examples from the client’s local data to guide GPT-3.5 generating the in-domain instruction and one example from the client’s local data for one-shot in-context learning to guide GPT-3.5 generating responses into the example’s format.
+Public dataset is composed of these five train sets (Alpaca, MedAlpaca, FinGPT, MathInstruct, and GSM8K).
+Test sets are marked with $\dagger$ and used for evaluation.
+Monotonicity and Submodularity.
+The facility location function $g(\mathcal{P})$ is known to be both monotone and submodular. Our objective $d(D^{d},\mathcal{P})$ inherits these properties.
+Monotonicity: This property is guaranteed because our benefit function, $\operatorname{sim}(p,c)$, is the cosine similarity from a sentence encoder, which is non-negative [^1] https://huggingface.co/BAAI/bge-large-en-v1.5 (we have plotted histograms of the similarity scores for each domain, which fall within the range about [0.2, 0.85]).
+When adding a new center to $\mathcal{P}$, the maximum similarity for any data point $c$ can only increase or stay the same.
+If a negative value ever arises, we can rectify the score via applying an affine shift, which preserves the order of similarities and therefore the monotonicity.
+Consequently, the total sum of maximum similarities, $g(\mathcal{P})$, cannot decrease.
+Submodularity (Diminishing Returns): To formally show this, let’s define the marginal gain for a single data point $c\in D^{d}$ when adding a new center $x$ to a set of centers $S$ as:
+$\Delta_{c}(S,x)=\max\bigl{(}0,\operatorname{sim}(x,c)-\max_{p\in S}\operatorname{sim}(p,c)\bigr{)}$ (5)
+This formula calculates the new coverage a center $x$ provides to a point $c$, beyond what is already provided by the existing set $S$.
+Now, consider two sets of centers $A$ and $B$ such that $A\subseteq B$.
+For any data point $c$, the maximum similarity from set $B$ must be at least as large as from set $A$, so:
+$\max_{p\in A}\operatorname{sim}(p,c)\leq\max_{p\in B}\operatorname{sim}(p,c)$ (6)
+Consequently, the marginal gain from adding $x$ must be smaller for the larger set $B$. That is, for any $c$:
+$\Delta_{c}(A,x)\geq\Delta_{c}(B,x)$ (7)
+The total marginal gain of adding center $x$ to a set $S$ is:
+$g(S\cup\{x\})-g(S)=\sum_{c\in D^{d}}\Delta_{c}(S,x)$ (8)
+Since the inequality holds for each individual term in the sum, it must hold for the sum itself. This demonstrates the diminishing returns property:
+$g(A\cup\{x\})-g(A)\geq g(B\cup\{x\})-g(B)\text{ for }A\subseteq B$ (9)
+Overall, our problem of maximizing the domain coverage $d(D^{d},\mathcal{P})$ is equivalent to solving the facility location problem of maximizing $g(\mathcal{P})$, ensuring theoretical guarantees for FedDCA.
+In addition, we have provided the time complexity analysis in Section˜A.1 and our empirical results show that the FedDCA algorithm converges quickly (about 1.1N passes).
+Furthermore, Section˜B.9 demonstrates that FedDCA achieves near-optimal performance in practice.
+Appendix B Experiments
+B.1 Train and Test Dataset Information
+We evaluate our method on multiple domains including medical (MedAlpaca, MMLU-Med), financial (FinGPT, FPB, FiQA, TFNS), and mathematical reasoning (MathInstruct, GSM8K).
+Table 5 shows the dataset statistics for each domain.
+The public dataset used for retrieval consists of the training sets from these domains and the general instruction dataset (Alpaca) through concatenation and random shuffling, while the test sets (marked with $\dagger$) are used for evaluation.
+B.2 Baselines
+To highlight FedDCA’s advantages, Table 4 provides a comparative overview with existing baseline methods along six critical dimensions: Privacy Preserving, API Cost, Additional Information requirements, potential for Performance Degradation, orientation towards Domain Coverage, and the Computational Overhead of Data Augmentation.
+These dimensions are crucial for assessing the practical viability of federated instruction tuning strategies.
+We analyze each dimension in detail below:
 Figure 4: Prompts used in the Self-Instruct data generation.
@@
 We prompt GPT-3.5 to generate responses with a randomly selected example for one-shot in-context learning.
-Prompt Used for Memory Extraction Attack.
+Figure 5: Prompt used for memory extraction attack.
+  • Privacy Preserving: Most evaluated methods, including FedDCA, inherently preserve client data privacy by design, as they do not share raw local data.
+Self-Instruct is a notable exception as it involves sending prompts containing potentially sensitive information to an external API.
+Our detailed privacy analysis for FedDCA can be found in Sections˜A.1 and B.5.
+  • API Cost: FedDCA operates without requiring external API calls, thus incurring no associated costs.
+In contrast, Self-Instruct, by leveraging large models like GPT for instruction generation, entails API expenses.
+  • Additional Information: FedDCA primarily relies on clients’ local training data for its operations.
+Some methods, such as LESS Xia et al.
+(2024), necessitate additional resources like a validation set for its gradient-based retrieval mechanism.
+  • Performance Degradation Avoidance: A key objective is to enhance model performance or, at minimum, avoid degradation.
+As indicated in our main results (Table 1), FedDCA is designed to prevent performance drops.
+In contrast, methods like FedAvg (when unaugmented), Random Sampling, and even some generation-based approaches like FewFedPIT and KnowledgeSG (if generation quality is not optimal or well-aligned) can risk performance degradation.
+  • Domain Coverage Oriented: FedDCA is uniquely engineered to explicitly optimize cross-client domain coverage in distributed environments.
+This strategic focus is a distinguishing feature largely absent in the other compared baselines, which typically do not have a direct mechanism for maximizing instruction diversity across the federation.
+  • Computational Overhead of Data Augmentation: FedDCA maintains a low computational overhead for its data augmentation phase, comparable to methods like LESS (due to gradient computation and warm-up), FewFedPIT (local LLM generation), and KnowledgeSG (expert model generation on the server) inherently involve higher computational demands for data augmentation.
+In conclusion, these comparisons underscore FedDCA’s well-rounded design.
+It effectively addresses key practical challenges by preserving privacy, avoiding API costs, minimizing reliance on additional information, preventing performance degradation, strategically optimizing domain coverage, and maintaining a low computational overhead for augmentation.
+B.3 Implementation Details
+We consider FedDIT in the cross-device scenario, $N$ = 10 clients, $\mathcal{R}$ = 30 rounds, where we randomly sample 2 clients to be available for each round.
+Then, each available client performs FedDIT for 10 steps with AdamW optimizer, and the batch size is $B=32$ in a round.
+The initial learning rate is 5e55e-5 with a cosine learning rate scheduler.
+Our experiment utilizes the widely used LLM, Llama3-8B [^2] https://huggingface.co/meta-llama/Meta-Llama-3-8B as the base model with 2048 max sequence length and adopts LoRA tuning method.
+The rank of LoRA is 16, and the scalar alpha is 16.
+For k-means Wu (2012), we set cluster num $\xi$ = 10 and for FedDCA we set the similarity threshold $\alpha$ = 0.7.
+We utilize bge-large-en-v1.5333https://huggingface.co/BAAI/bge-large-en-v1.5 as both the client and server’s encoder as default, which outputs embeddings of 1024 dimensions.
+Figure 6: Impact of different retrieval amounts on FedDCA’s performance (%) and domain coverage.
+The dotted lines represent the domain coverge.
+Results demonstrate FedDCA’s robust applicability across different augmentation scales.
+Specifically, for LESS, we uses LoRA with rank 128 and dropout 0.1 for warmup training on a random 5% subset of the public dataset for 4 epochs.
+We apply a learning rate peak of $2\times 10^{-5}$ with linear warm-up and cosine decay.
+Gradients are extracted from each epoch checkpoint and projected to a 1024-dimensional space using Johnson-Lindenstrauss random projections.
+Data selection is performed by computing cosine similarity between training and validation gradient features aggregated across epochs.
+The top $N_{k}^{p}$ scoring examples are selected for final client $c_{k}$’s instruction tuning.
+Additionally, for KnowledgeSG, we utilize AlpacaCare Zhang et al.
+(2023d), FinGPT Yang et al.
+(2023a) and WizardMath Luo et al.
+(2023) as the server-hosted expert model for each domain respectively.
+To generate the Self-Instruct data, we prompt GPT-3.5 to generate the instruction with the designed prompt in Figure 4.
+Specifically, we randomly sample two examples from the client’s local data to guide GPT-3.5 generating the in-domain instruction and one example from the client’s local data for one-shot in-context learning to guide GPT-3.5 generating responses into the example’s format.
+B.4 Effect of Retrieval Number
+(a) Different amounts of augmented public data for FedDIT.
+(b) The average ROUGE-L score per round. Dotted lines represent the base-data-only setting.
+Figure 7: Privacy analysis of memory extraction attacks. Data augmentation can effectively mitigate the privacy leakage risk.
+Experimental Setup.
+To assess the impact of the volume of augmented data, we vary the number of public instructions retrieved per client by FedDCA.
+Specifically, we test retrieval amounts of [100, 200, 400, 800, 1000] instructions per client, while keeping other experimental parameters consistent with the main setup described in Section˜6.1.
+The experiments are conducted across the three primary domains (medical, financial, and mathematical).
+Results.
+Figure˜6 shows FedDCA’s performance and domain coverage with varying retrieval amounts.
+Both generally increase with more retrieved instructions, indicating richer augmentation enhances learning.
+The improvement rates slow after about 400 instructions per client, likely due to increasing overlap with already covered semantic space or the model reaching its capacity to benefit from further diverse examples under the given training regime.
+Overall, performance and coverage continue to rise with retrieval volume.
+Notably, domain coverage’s impact on performance differs by domain.
+For example, medical domain gains are modest, while financial domain gains are more aligned with coverage increases.
+This variance may stem from: 1) The base model’s initial domain proficiency, which influences its baseline performance and the marginal benefits of additional data.
+2) The degree of similarity between the public in-domain data distribution and the test set distribution, where higher relevance yields more proportional performance benefits.
+B.5 Privacy Analysis
+Memory Extraction Attack.
+We evaluate the privacy-preserving capability of different ratios of public data against memory extraction attacks Carlini et al.
+(2021); Zhang et al.
+(2024b), which utilizes the autoregression nature of LLMXu et al.
+(2024b).
+We focus on one client’s instruction tuning in FedDIT, using FedDCA for instruction augmentation with 100 to 1k public instructions.
+We set up 10 clients with full participation for 10 rounds.
+Specifically, we record the average ROUGE-L score Lin (2004) for client $c_{0}$ in each round (following the setup in Zhang et al.
+(2024c)).
 As we use Llama3-8B as our base model and format the instructions and responses into the Alpaca’s format, to utilize the auto-regression nature of LLM to extract the instruction, we prompt the model to generate the instruction using the prompt in Figure 5, which is exactly the prefix of the Alpaca’s template.
-Figure 9: Prompt used for memory extraction attack.
-C.2 Baselines
-Method Privacy Preserving API Cost Additional Information Performance Degradation Domain Coverage Oriented
-FedAvg McMahan et al. [2017]
-Random Sampling
-LESS Xia et al. [2024]
-Self-Instruct Wang et al. [2022]
-Direct Retrieval
-FedDCA (ours)
-Table 6: Key differences between FedDCA and other baselines.
-FedDCA demonstrates the ability of: 1) privacy-preserving, 2) no API cost, 3) no additional information required, 4) avoiding performance degradation, and 5) aiming at domain coverage optimization.
-LESS requires additional information, as it needs access to the validation set for gradient-based retrieval.
-In Table 6, we highlight five critical aspects of our approach compared with other methods Xia et al. [2024]; Wang et al. [2022]; Zhang et al. [2024b]. Our comparison focuses on five key dimensions:
-  • Privacy Preserving: Most methods except Self-Instruct maintain data privacy, we analyze the privacy-preserving ability of FedDCA in Sections 5 and C.3.
-  • API Cost: FedDCA requires no api calls. As Self-Instruct utilizes the GPT to generate instructions, it requires additional API calls, which incurs costs.
-  • Additional Information: LESS needs validation set access for gradient-based retrieval, whereas FedDCA works with only the basic client’s local training data.
-  • Performance Degradation: From Table 2 we see that there occurs performance degradation for FedAvg, Random Sampling, and Self-Instruct, while FedDCA avoids this issue.
-  • Domain Coverage: Uniquely, FedDCA is specifically designed to optimize domain coverage for distributed environments, a feature absent in other baseline methods.
-Specifically, Self-Instruct Wang et al.
-[2022] represents the upper limit performance of the generation-based method, which is restricted to the high expense of cost, limited API call frequency, and heavy quality screening process.
-On the other hand, FewFedPIT Zhang et al.
-[2024b] attempts to augment local data by leveraging the pre-trained Llama2 model to perform self-instructed generation during training.
-However, two concerns exist: 1) The pre-trained LLM cannot provide effective, stable, and high-quality instruction generation.
-2) Generating instructions locally is very costly regarding both time and computational resources.
-Overall, generating instructions self-instructively is not a satisfactory method for the FedDIT scenario.
-LESS Xia et al.
-[2024] represents the augmentation methods through gradient feature retrieval.
-However, its performance is underwhelming.
-Compounding this issue, LESS presupposes access to the validation set, which is not always available, and necessitates an initial warmup on the public dataset before calculating gradients for both the public and validation data.
-This process becomes computationally burdensome with large public datasets.
-In conclusion, these comparisons demonstrate FedDCA’s distinctive approach to balancing privacy preservation, computational overhead and domain coverage optimization, while maintaining minimal additional requirements.
-C.3 Domain Inference Attacks
+For each setting, we repeat memory extraction 100 times and report the average ROUGE-L score.
+Specifically, denote the generated $\mathcal{N}$ instructions as $\mathcal{I}$ and the client’s local instructions $\mathcal{I}^{l}$.
+The calculation of the average ROUGE-L score is defined as $\frac{1}{\mathcal{N}}\sum_{i=1}^{\mathcal{N}}\texttt{ROUGE-L}(\mathcal{I}_{i},\mathcal{I}^{l})$.
+Figure˜7(a) shows no significant correlation between the public data ratio and privacy-preserving capability in the same training round.
+In addition, only using local data has a higher risk of privacy leakage than augmented methods.
+Additionally, Figure˜7(b) shows the trends of the average ROUGE-L score per round.
+Initially, the average ROUGE-L score for augmented settings increases, then decreases or converges, while the base-data-only scores continue to rise, especially in the code domain.
+This indicates that with more training rounds, base-data-only fine-tuning captures more privacy information, while the privacy leakage risk in augmented fine-tuning decreases or converges.
+Domain Inference Attack.
 The server may inference the clients’ data domain when the domain data retrieval is performed on the server side.
@@
 Both encoders output 1024-dimensional features.
-Example 1: Both $w_{1}$ and $w_{2}$ take "hello world" as input, and the cosine similarity between their embeddings is 0.1829.
+Example 1: Both $w_{1}$ and $w_{2}$ take ‘‘hello world’’ as input, and the cosine similarity between their embeddings is 0.1829.
 Example 2: Three instructions are used:
@@
 In conclusion, as demonstrated above, when the clients do not perform domain-specific instruction retrieval on the server side, the server cannot infer the client’s domain based on the uploaded embeddings.
-C.4 How Far can FedDCA Go
-Assume an extreme setting, where exists a pre-trained model that has already fine-tuned on the whole in-domain data of the internet or even all open-source public instruction datasets (it is actually unrealistic or temporary, as new data is constantly being generated), is there any room for instruction augmentation or FedDCA?
-To answer this question, we conduct the following experiments on four settings: A) Fine-tuning on the whole public dataset.
-B) Following A, perform further federated fine-tuning with only private clients’ local dataset.
-C) Following A, each client randomly samples 100 public data for fine-tuning.
-D) Following A, each client samples 100 public data through FedDCA and then performs FedDIT.
-Setting H-Eval M-Med FPB FiQA TFNS GSM8K
-A 38.41 69.10 75.33 41.09 71.60 53.75
-B 43.90 64.30 77.97 66.54 78.26 56.40
-C 41.46 66.90 74.50 29.81 70.51 57.01
-D 44.12 70.90 81.43 72.00 78.81 58.90
-Table 7: Performance (%) of setting A, B, C, and D in each domain.
-As shown in Table 7, even with only 100 samples selected via FedDCA, the method not only prevents performance degradation but also helps the model achieve better generalization within the domain.
-This is attributed to exposure to more diverse in-domain data during training.
-Meanwhile, it can be observed that random sampling still leads to performance degradation compared to Setting A, and even performs worse than Setting B, where no data augmentation is applied.
-In the financial domain, performance degradation is also evident.
-In conclusion, this experiment further highlights the necessity and effectiveness of FedDCA.
-C.5 Augmentation Strategy Visualization
-To more intuitively compare the domain coverage of different instruction augmentation methods, we randomly sample 5,000 instructions obtained through these methods and 10,000 in-domain instructions as the background, representing the distribution of specific domains in the public dataset.
-We then visualized the results using t-SNE van der Maaten and Hinton [2008], as shown in Figure 10.
+MMLU-Med FPB GSM8K
+Zero-shot 70.60/- 55.94/- 23.27/-
+Base Data 72.40/0.8377 66.74/0.9339 49.12/0.8709
+Random Sampling 69.90/0.8497 61.05/0.9408 47.23/0.8812
+FedDCA 73.30/0.9090 67.16/0.9800 50.26/0.9118
+Table 6: Scalability. Performance (%) and domain coverage of FedDCA and other baselines. 100 clients with 2 clients per round. Results show the strong scalability of FedDCA.
+B.6 Scalability
+(a) FPB
+(b) FiQA
+(c) TFNS
+Figure 8: Held-out setting. We report the performance of FedDCA and other baselines for (a) FPB, (b) FiQA, and (c) TFNS datasets. Results show the robustness of FedDCA in the held-out setting.
+Figure 9: Data selection scenarios.
+We compare the performance of using the full dataset (Setting A) and using only 10% of data selected through FedDCA (Setting B), demonstrating FedDCA’s efficiency in data utilization.
+Experimental Setup.
+To evaluate the scalability of FedDCA, we conduct experiments with 100 clients, where 2 clients are randomly selected for FedDIT in each round.
+As the number of clients increases, the amount of local data on each client gradually grows, allowing us to assess how FedDCA performs in larger-scale distributed settings.
+Results.
+Table˜6 presents the performance and domain coverage across different domains.
+Interestingly, we observe that the model trained on only base data even outperforms Random Sampling in domains other than code and narrows the gap with FedDCA.
+This suggests that with sufficient local data, the benefits of data augmentation become more nuanced.
+However, FedDCA maintains its advantage by consistently achieving higher domain coverage and better performance across all domains, demonstrating its effectiveness in large-scale federated settings.
+This superior performance can be attributed to FedDCA’s strategic approach to maximizing cross-client domain coverage, which remains crucial even when dealing with a larger overall volume of local data distributed across a greater number of clients.
+B.7 Data Selection Scenarios
+Experimental Setup.
+We conduct the experiments on two settings: A) For each domain’s FedDIT, we randomly dispatch the in-domain data equally to each client.
+For example, in the code domain, as the size of the public dataset is 20,022, each client is assigned about 2,000 samples.
+Then we perform FedDIT on each client’s whole data.
+B) Following A’s setting, we first perform FedDCA to determine the client center set, and then we conduct the data selection based on the client center set through dense retrieval to select 200 samples for each client.
+Finally, we perform FedDIT on the selected data.
+Results.
+Figure˜9 demonstrates that FedDCA can achieve comparable performance to training with the full dataset by using only 10% data through data selection.
+This efficiency in data utilization highlights the effectiveness of FedDCA’s data selection strategy, which maintains model performance while significantly reducing the required training data.
+B.8 Held-out Setting
+If distributed clients aim to solve tasks based on existing knowledge, the public dataset will inevitably contain knowledge relevant to those domains.
+This could come from the original corpus (which can be converted into instruction-response pairs by GPT) or from pre-constructed instruction datasets on the website.
+So the distribution of the public dataset can be categorized as follows: containing held-in or held-out instructions.
+The held-in indicates that the public dataset contains instructions for the specific task that clients aim to solve, while the held-out indicates that the public dataset does not contain this task’s instructions.
+The paper’s default setting is the held-in setting.
+Considering the held-out setting in the financial domain, given that the training set FinGPT and the test sets FPB, FiQA, and TFNS are all related to sentiment analysis tasks.
+We keep the setting of test sets and replace the FinGPT’s instructions in public data with data from a financial QA dataset (Sujet-Finance-Instruct-177k [^4] https://huggingface.co/datasets/sujet/Sujet-Finance-Instruct-177k).
+The clients’ local data are still randomly sampled from FinGPT.
+This approach yields held-out public data.
+Figures˜8(a), 8(b) and 8(c) shows, FedDCA still achieves performance improvements compared to other baselines for the held-out setting.
+Additionally, using the Random Sampling data augmentation strategy resulted in performance degradation on the FiQA dataset.
+This further underscores the necessity of selecting an appropriate data augmentation strategy.
+B.9 Empirical Tightness of FedDCA’s Greedy Approximation
+FedDCA’s optimization objective is equivalent to the classic sub-modular facility-location problem, which guarantees that the simple greedy algorithm attains at least a 11/e1-1/e fraction of the optimal value in the worst case.
+To gauge how tight this bound is in practice, we benchmark the FedDCA algorithm against increasingly stronger (but more expensive) baselines on three domains: Medical, Financial and Mathematical.
+In particular we position FedDCA between the beam search (with various beam widths) and an infeasible brute-force solver that attempts to enumerate all candidate subsets.
+The beam search, which explores multiple promising branches simultaneously, serves as a realistic upper bound that is strictly better than greedy yet still tractable.
+Experimental Setup.
+  1. 1.
+    Metric: Domain coverage $d(D^{d},\mathcal{P})$ where $\mathcal{P}$ is the selected $N$ client centers, defined as Eq. 2.
+  2. 2.
+    Baselines: $\mathcal{C}_{\text{all}}$ (default FedDCA) selects client centers using the union of each client’s uploaded cluster center set $\mathcal{C}_{\text{all}}$ as the selection reference to compute marginal gains; $\mathcal{P}^{*}_{C}$ is the unreachable optimum under the same selection reference $\mathcal{C}_{\text{all}}$—marked timeout, requiring evaluation of $\binom{100}{10}\approx 1.73\times 10^{13}$ distinct candidate sets; $D^{d}$ uses $D^{d}$ as the selection reference to compute marginal gains, which is often impractical in real-world federated settings (please refer to Section˜5.2); $\mathcal{P}^{*}_{D}$ is the optimum under $D^{d}$—also timeout; Beam search with widths $256,512,1024,2048$ and, for Medical, $9182$ (does not exceed the baseline $D^{d}$ until width $9182$).
+  3. 3.
+    Approximation ratio: We report
+    $\frac{d(D^{d},\mathcal{P}_{\mathcal{C}_{\text{all}}})}{\max_{w}d(D^{d},\mathcal{P}_{\text{beam-}w})}\times 100\%,$ (10)
+    where the denominator is the best coverage attained by the beam search.
+$\mathcal{P}_{\mathcal{C}_{\text{all}}}$ is the selected client center set by FedDCA using $\mathcal{C}_{\text{all}}$ as the selection reference.
+Domain $\mathcal{C}_{\text{all}}$ $\mathcal{P}^{*}_{C}$ $D^{d}$ $\mathcal{P}^{*}_{D}$ Beam-256 Beam-512 Beam-1024 Beam-2048 Beam-9182 Approx. Ratio
+Med. 0.6772 timeout 0.6782 timeout 0.6776 0.6776 0.6776 0.6778 0.6782 99.85%
+Fin. 0.8379 timeout 0.8418 timeout 0.8422 0.8427 0.8427 0.8427 99.43%
+Math. 0.6391 timeout 0.6699 timeout 0.7401 0.7401 0.7404 0.7405 86.30%
+Table 7: Empirical tightness analysis of FedDCA’s greedy approximation.
+Both theoretical optima ($\mathcal{P}^{*}_{C}$ and $\mathcal{P}^{*}_{D}$) are computationally unattainable (timeout).
+The approximation ratio is computed against the best beam search result.
+Results.
+As shown in Table˜7, FedDCA using $\mathcal{C}_{\text{all}}$ already achieves nearly optimal coverage in Medical and Financial domains, within 99.85% and 99.43% of an extensive beam search respectively, and still attains a respectable 86.30% in the more challenging Mathematical domain.
+Note that both theoretical optima ($\mathcal{C}_{\text{all}}$ and $D^{d}$ as reference) are computationally unattainable (timeout), so the beam search constitutes the strongest practical competitor.
+In conclusion, these real-world ratios vastly exceed the theoretical lower bound of $1-1/e\approx 63\%$, further corroborating that the simple greedy procedure is both efficient and effectively near-optimal for real-world federated deployments.
+Appendix C Robustness to Data Heterogeneity
+The efficacy of retrieval-based augmentation strategies in FedDIT, such as Direct Retrieval and our proposed FedDCA, is inherently linked to the semantic characteristics of clients’ local private data, as these inform the selection of public instructions.
+A critical consideration is how varying degrees of inter-client data heterogeneity impact the ability to construct diverse and comprehensively covering augmented datasets.
+Intuitively, if clients possess highly similar local data, methods relying on these local signals might struggle to achieve broad cross-client domain coverage and instruction diversity in the augmented sets, which are crucial for a well-generalized global model.
+This section, therefore, investigates the robustness of FedDCA, Direct Retrieval, and Random Sampling to a spectrum of inter-client data heterogeneity.
+We focus on these methods as their augmentation processes are either directly influenced by local data distributions (FedDCA, Direct Retrieval) or serve as a data-agnostic baseline (Random Sampling); other generative or complex retrieval methods often have different operational dependencies.
+We control heterogeneity using a Dirichlet distribution ($\text{Dir}(\beta)$) over semantic clusters of clients’ private instructions, where smaller $\beta$ values yield higher inter-client heterogeneity (more distinct local data) and larger $\beta$ values lead to lower heterogeneity (more similar local data).
+We evaluate performance across $\beta\in\{0.01,0.1,1.0,10\}$.
+C.1 Experimental Setup
+Inter-Client Data Heterogeneity Control.
+We focus on the Financial domain, using FinGPT as the source for private instructions.
+First, we perform k-means clustering ($\xi=100$ clusters, consistent with Section˜4.1) on FinGPT instruction embeddings to create semantic pseudo-labels.
+To generate local data for 10 clients, each holding 100 instructions, we employ a Dirichlet distribution $\text{Dir}(\beta)$ over these clusters to define each client’s data composition, simulating varying degrees of inter-client data heterogeneity for $\beta\in\{0.01,0.1,1.0,10\}$.
+Augmentation and Training.
+Following the general setup in Section˜6.1, Random Sampling, Direct Retrieval (introduced in Section˜3), and FedDCA are used to obtain 1k augmented instructions per client before commencing FedDIT.
+Evaluation Metrics.
+Model performance is evaluated using FPB, FiQA, and TFNS accuracy. Augmented data characteristics are assessed using three key metrics:
+1) Domain Coverage ($\uparrow$): As defined in Eq. 2, this measures how well the aggregated client data (local + augmented) $D^{c}$ covers the target financial domain $D^{d}$.
+2) ICACS (Inter-Client Augmented Centroids Similarity $\downarrow$): For each client $k$, its augmented instruction embeddings are clustered (k-means, $\xi=10$) into centers $\mathcal{C}_{k}$.
+ICACS is the average pairwise cosine similarity between all such cluster centers across different clients.
+Lower values indicate more distinct augmented sets at a granular level.
+3) RUAI (Ratio of Unique Augmented Instructions $\uparrow$): The ratio of unique instructions within the total cross-client augmented data pool $D^{c}$. Higher values signify less redundancy.
+C.2 Results and Analysis
+Dirichlet $\beta$ Augmentation Model Performance (%) Augmented Data Characteristics
+Strategy FPB FiQA TFNS Domain Coverage ($\uparrow$) ICACS ($\downarrow$) RUAI ($\uparrow$)
+0.01 Random Sampling 64.20 13.10 65.55 0.920 0.472 0.867
+Direct Retrieval 67.20 21.50 69.00 0.938 0.434 0.909
+FedDCA 67.90 36.30 74.40 0.985 0.405 0.943
+0.1 Random Sampling 64.19 13.09 65.53 0.920 0.472 0.867
+Direct Retrieval 66.31 19.11 67.62 0.929 0.441 0.882
+FedDCA 67.24 35.27 73.32 0.982 0.411 0.938
+1.0 Random Sampling 64.19 13.09 65.53 0.919 0.472 0.867
+Direct Retrieval 65.50 17.80 66.79 0.924 0.448 0.879
+FedDCA 66.94 34.82 72.66 0.977 0.414 0.927
+10 Random Sampling 64.20 13.05 65.45 0.918 0.472 0.867
+Direct Retrieval 64.37 15.20 66.10 0.922 0.452 0.871
+FedDCA 66.73 34.51 72.23 0.972 0.423 0.921
+Table 8: Performance (%), domain coverage, ICACS and RUAI across different levels of client data heterogeneity (controlled by Dirichlet parameter $\beta$) in the Financial Domain.
+Best model performance, domain coverage, ICACS and RUAI per $\beta$ setting are in bold.
+Table˜8 systematically quantifies the impact of inter-client data heterogeneity on augmentation strategies.
+For small $\beta$ (i.e., high heterogeneity), the cross-client data exhibits substantial diversity, enabling Direct Retrieval to access a broader range of semantic regions in the public pool.
+This diversity translates into clear gains over Random Sampling in both performance (e.g., +6% FiQA at $\beta=0.01$) and coverage metrics.
+However, as $\beta$ increases and client data distributions become more homogeneous, the advantage of Direct Retrieval diminishes, with its performance and diversity metrics converging toward those of Random Sampling.
+This trend underscores that the effectiveness of Direct Retrieval is fundamentally driven by the diversity present in cross-client data.
+Conversely, as inter-client heterogeneity decreases (i.e., local data distribution becomes more similar), Direct Retrieval’s effectiveness continuously degrades.
+Its model performance and domain coverage drop, showing a more sensitive effectiveness to the heterogeneity.
+Results indicate that with more homogeneous local data distributions, Direct Retrieval tends to fetch more similar public instructions due to the similarity between each client’s cluster centers.
+Random Sampling exhibits stable, albeit consistently sub-optimal, performance and diversity metrics across all heterogeneity levels.
+This stability stems from its sampling strategy (oblivious to local data distributions) of the public dataset, which results in similar augmented data.
+Additionally, FedDCA showcases strong robustness.
+It consistently achieves the best model performance across all tested $\beta$ values.
+While its diversity metrics (domain coverage, ICACS, RUAI) show minor sensitivity to the cross-client data distribution (optimally diverse at highest heterogeneity).
+This resilience stems from its coverage-oriented client center selection (Algorithm˜1), which effectively prevents redundant augmentation by selecting public instructions that add novel semantic aspects to the collective, even with similar cross-client data distributions.
+In conclusion, these findings underscore FedDCA’s robustness and adaptability irrespective of the specific degree of inter-client heterogeneity, makes it a more reliable and effective solution for real-world FedDIT deployments.
+Appendix D Augmentation Strategy Visualization
+To more intuitively compare the domain coverage of different instruction augmentation methods, we randomly sample 5k cross-client instructions obtained through these methods and 10k in-domain instructions as the background, representing the distribution of specific domains in the public dataset.
+We then visualized the results using t-SNE van der Maaten and Hinton (2008), as shown in Figure 10.
 The plot shows that FedDCA encompasses most of the in-domain data, which is consistent with FedDCA’s domain coverage of each domain shown in Table 1.
 Also, we can observe that the random sampling strategy selects a lot of out-of-domain data while does not have good coverage in specific domains.
-(a) Code
-(b) Medical
-(c) Financial
-(d) Mathematical
+(a) Medical
+(b) Financial
+(c) Mathematical
 Figure 10: Visualization of cross-client data distribution in different domains, performing t-SNE dimensionality reduction on retrieved instructions through various augmentation strategies.
-We randomly sample 10,000 in-domain samples as background while randomly sampling 5,000 samples from the cross-client augmented dataset for different instruction augmentation methods for comparison.
+We randomly sample 10k in-domain samples as background while randomly sampling 5k samples from the cross-client augmented dataset for different instruction augmentation methods for comparison.